### v1

# Intelligent-Based Neural Networks and Optimal Control of Fractional Order Ebola Virus Dynamics

Abstract
Ebola virus disease is a severe, often deadly form of hemorrhagic fever, and is rapidly transmissible. It’s spread through human contact is increased through the direct exchange of infected blood, body fluids, and organs, as well as contact with surfaces on which these infected fluids have been dropped. Ebola’s complicated dynamics spread beyond integer order models, representing a central intersection for the application of machine learning to control theory. In the present study, we construct an innovative prime framework of fractional order calculus which integrates data-driven machine learning with optimal control in order to model and analyze the Ebola virus disease. We develop an eight-compartment model based on the Caputo fractional derivative to incorporate the non-Markovian memory effects inherent in the disease’s progression and regression. The system explicitly structures the distinct transmission pathways from symptomatic and asymptomatic infections, as well as the critical transmission dynamics associated with deceased individuals. We analytically prove the global well-posedness of the model, compute the basic reproduction number $\mathcal{R}_{0}$ to characterize the epidemic threshold, and prove global stability theorems. Sensitivity analysis indicates that the basic reproduction number is most critically sensitive to transmission rate, incubation period, and the relative infectivity of the deceased. The treatment-safe burial synergy optimization control provides a simultaneous 86.5% of the target morbidity and mortality rate control of the population. The safe burial host intervention is the most effective control. Our study categorizes the best control strategies and orders them in terms of effectiveness. The proposed disease-informed neural network achieves practically perfect predictive accuracy and describe the test set $R^{2}$ values of 0.991 - 0.999, and correspondingly 99.1% - 99.9% prediction accuracy as well as align with the real-world epidemic behavior. The first successful combination of process-based modeling with artificial intelligence for dealing an outbreak of an infectious disease has been established in this work. This work presents an innovative mechanistic modeling framework for infectious diseases, which has the potential to advance epidemic response analysis.
Keywords: Fractional Calculus, Ebola Modeling, Epidemic Threshold ($\mathcal{R}_{0}$), Sensitivity Analysis, Optimal Control, Disease-Informed Neural Networks.
1 Introduction
Ebola Virus Disease (EVD) continues to be one of the most difficult problems for the world health organization to tackle, with estimated case fatality rates varying from concerning 25% to 90% across various outbreaks [1]. The Ebola Virus Disease (EVD) is classified as a tropical zoonosis which involves cross transmission from animal reservoirs to human populations [2, 3]. The earliest recorded case of the disease occurred in Zaire (now known as Democratic Republic of Congo) in 1976, named after the Ebola River where the first outbreak took place [4]. With data around confirmed cases and deaths standing at 28,646 and 11,323, respectively, and spanning through three countries; Liberia, Guinea, and Sierra Leone, the ‘West African Ebola Virus Disease epidemic’ (2014-16) is considered the biggest ‘Ebola’ outbreak in history. The surprising number of cases and deaths during the outbreak and the new ways of fighting the complexity both demonstrated the failure of prediction and outbreak control traditional prediction and control methods used [5]. The outbreak still emphasized the complexity controlled in the actual mathematics and logic of the base prediction and control frameworks prediction, controlling, and modeling control frameworks outlined around the multi and complex dimensions in the spread of the EVD. Mathematical techniques in epidemiology and their application span new prospects and new dimensions in the field [6, 7]. While these techniques were developed in pure science, their fundamental calculations and distributions resource divisions and practical works approximations [8]. The epidemiology of the western African case outbreak hence has triggered the much-needed efficiency flow in prediction case numbers and evaluating the efficiency of preventive measures and their real-time value [9]. The greatest puzzling challenge suffering in the environmental and acoustic scales spread while modalities of EVD transmissions is the combination of factors that various phoneme extents growing paradigm [10]. The epidemiology of Ebola Virus Disease (EVD) has many distinguishing features which imply that estimating the disease outbreaks remaining traditional techniques will not be sufficient. For instance, the disease’s incubation period is highly variable and spans anywhere from 2 to 21 days, with the 8-10 days interval more commonly renowned [11]. There is a further complication within the disease’s clinical presentation and development as initial fever, weakness, and diarrhea unaccompanied by bleeding and other clinical symptoms do not signal an active infection and contagiousness. Additionally, contact with infected animals and the consumption of bush-meat are the first chance for the disease’s human transmission from its natural reservoirs of fruit bats [12, 13]. Subsequently, the disease transmission post- infection is unlinked and random. Furthermore, symptomatic ($I_{s}$), asymptomatic ($I_{a}$), and post-mortem ($D$) transmission routes for the human population. [14]. The high level of infectiousness of dead body, coupled with the perseverance of customarily traditional burial practices in the region, requires mathematically sophisticated treatment beyond typical integer-order differential equations, as outlined in earlier works [15]. Evidence shows that washing the deceased as part of ethnically prescribed performs is a relevant vector for transmission. Besides, the prospect of reinfection with EVD remains very real, as survivors own only quickly protective, low-potency antibodies, with research showing the complete absence of Ebola antibodies in some survivor’s years post-recovery [16, 17]. Such an immunological situation is certainly an additional complication from the modeling perspective whereby the absence of an unified framework for archiving immune waning over a long period rightly remains a gap. The natural, spatial spread of EVD transmission improves so far, another level of complexity to the analytical modelling of the disease dynamics. The dynamics of EVD differ significantly across different geographic scales. Urbanized areas tend to experience outbreaks at a faster rate than rural areas because of the high population density, varied patterns of movement, and better medical infrastructure [18]. Moreover, the phenomenon of superspreader events, where a small proportion of the infected population is responsible for a suspiciously high number of secondary infections, adds a considerable amount of variability and complexity that challenges resolution by purely deterministic approaches [19, 20]. These characteristics need modeling methods that combine large-scale spatial dynamics with fine-scale spatio-temporal transmission patterns [21]. Although traditional modeling techniques such as the SI, SIR, and SEIR frameworks form the basis of understanding, they all fundamentally relate to the Markovian assumptions which do not take into account memory and long-range chronological dependencies which are characteristic of EVD spread [22, 23]. These limitations are particularly clear when attempting to model the distributed incubation periods, the decline of vaccine-induced immunity, and the heterogeneous contact patterns during the real-world epidemic spread [11, 24]. The SEIR model used by [25] and [26], and the SIHDP model employed by [27] and in [28] SEIHFDBR are notable developments, however, they are still limited in the integer-order nature of their models. Fractional calculus has long shown the potential to model memory effects and other non-local connections, and thus provides a framework well-intentioned of study to overcome some of the noted deficiencies: [29]. Unlike integer-order derivatives, which always adopt Markovian processes, fractional operators capture the distributed suspensions which are characteristic of epidemiological systems, such as extended incubation period, waning immunity, and environmental reservoirs of pathogens [30]. Evidence of progress in the modeling of infectious diseases for which data are available and for which fractional calculus applies exhibits greater concord with experimental indication, especially for diseases with long-range chronological correlations and power-law decline in the efficiency of involvements [31, 32]. The non-Markovian nature of some fractional derivatives offers a better-suited mathematical framework for modeling the transmission of disease in and among structured populations [33]. In addressing challenges, some recent studies have included various extensions. For example, [34] included pathways for environmental contamination in their SIRPD model with respect to the Ebola pathogens in the environment. In [35] worked on the decomposition of the deceased compartment by distinguishing dead but not buried, and dead who are buried. Control optimization has been studied by, for instance, [36] on optimal combinations of vaccination and treatment, and [37] on educational campaigns in a two-patch model. Media awareness campaigns have been studied by [38] and [39] and have shown the importance of information broadcasting in controlling a disease. The link of Machine Learning algorithms with Process-based model and models of machine learning with traditional epidemiological forecasting is a new concept as noted by [40]. Predictable models afford a theoretical understanding of epidemiology through biological lenses and as such are weak at capturing sophisticated structures non-linear in nature that are defined with basic compartmental approaches [41]. Recent activities in hybrid frameworks combining the Physics-based model interpretability and the neural networks’ model complexity to capture epidemiological dynamics have demonstrated improved predictive performance [42]. But, as advanced computational techniques apply to EVD, there are gaps such as real time parameter estimation from EVD noise where EVD observation data is incomplete [43]. The development of neural networks with physics-informed neural network (PINN) and epidemiology constraints could bridge this gap [44]. However, modern modeling system contexts still have three main shortcomings. One, they have a limited representation of the memory effects of fractional-order derivatives, which is crucial for modeling the distributed intervals in the progression of a disease and the corresponding delay of the immune response. Two, there is still a lack of fully integrating optimal control theory combined with advanced multi-compartment models that summarize the entirety of the EVD transmission dynamics. Three, here is still insufficient use of machine learning for the estimation of parameters for complicated epidemiological models which reduces the flexibility and accuracy of the models in current practice. The application of optimal control theory is underdeveloped yet central to the design of effective intervention and control strategies for the economic epidemiology systems. Identifying optimal control strategies is a challenging mathematical problem. It requires temporally-resolved control strategies and sophisticated numerical methods for high dimensional, nonlinear systems with memory [45]. Also, minimizing the case count, death, and economic loss for epidemic reduction response requires Pareto-optimal strategies that balance the contradictory objectives [46]. The recent, rapid advancements in adaptable computational optimal control technologies—particularly those dealing with the direct collocation of the control function and the system trajectory and the Hamiltonian-based methods are methods of great attention in computational biologically structured control [47]. The development of fractional calculus has presented new ways to tackle these problems [29]. The Caputo fractional derivative approaches non-local chronological dependencies and memory effects with its non-Markovian structure, which mathematically captures epidemiological data [48]. Preliminary work connecting fractional-order models focused on EVD, such as [49] with the Atangana-Baleanu-Caputo operators and [50] with the nonlinear time-fractional mathematical model. Despite these advances, there is still much to be anticipated in the overlap among these fields, mainly control theory, machine learning, and the other works on the theme. This research discourses these important gaps by revolutionary a new computational framework that combines fractional-order dynamics, multi-pathway transmission architectures, and control optimization, which is enhanced by machine learning. Building upon the preceding work by [34] which did not include exposed compartments, hospitalizations, and safe burial practices, I also consider reinfection through the immunological studies by [16] and [17]. This research directs numerous contributions and novelties:
[…]
  • Comprehensive mathematical analysis signifying the well-posedness of the system through Banach’s fixed-point theorem, the basic reproduction number of the model $\mathcal{R}_{0}$ exogenously resulting through the next generation matrix, and the proof of global stability of the system through fractional Lyapunov methods [48].
  • Derivation and providing an analytical solution for a multi-objective optimal control problem with four intervention strategies (personal-protection, vaccination, treatment, and safe burial) via Pontryagin’s Maximum Principle.
  • Establishing a Disease-Informed Neural Network (DINN) framework for simultaneous solution discovery and parameter estimation which inserts physical constraints for better predictive performances.
  • Detailed cost-effectiveness analysis for optimal intervention portfolios for public health choices.
This model potentially assistances public health authorities in responding to outbreaks more effectively by evidence-based intervention, which can be game-changing in undertaking and enhancing the efficiency of outbreak responses. Integrating mathematical modeling with computational intelligence, this study attempts to productive infectious disease interventions through addressing the organized response and gaps in the current outbreak response system. This paper proceeds as follows. In Section 2, I discuss the elementary ideas of fractional calculus which is key to developing the model. In Section 3, I develop an eight-compartment, fractional-order Ebola transmission model. Section 4 contains the in-depth mathematical theory regarding the model’s well-posedness, equilibrium, and stability over the system. Section 5 conducts global sensitivity analysis of the dominant parameters through Latin Hypercube Sampling and Partial Rank Correlation Coefficients. In Section 6, I formulate and solve the optimal control problem based on Pontryagin’s Maximum Principle. Section 7 incorporates the neural network model concerning disease and the results obtained in the corresponding numerical analysis. In the closing Section 8, I analyze the outcome of the model along with its drawbacks and suggest some directions for further research.
2 Preliminaries
This part sets out the basic principles of fractional calculus which is essential to the construction and analysis of our model.
Definition 2.1 (Caputo Fractional Derivative)
Let $\alpha\in(0,1]$ and $u\in C^{1}([0,T])$. For any function $u$ which is continuously differentiable in the closure $[0,T]$ the Caputo fractional derivative of order $\alpha$ is defined in [29] as follows:
$^{C}D_{t}^{\alpha}u(t)=\frac{1}{\Gamma(1-\alpha)}\int_{0}^{t}(t-\tau)^{-\alpha}u^{\prime}(\tau)d\tau.$
This operator is preferred for modeling dynamical systems as it accommodates standard initial conditions, and $^{C}D_{t}^{\alpha}u(t)\to u^{\prime}(t)$ as $\alpha\to 1^{-}$, recovering the classical integer-order derivative.
Definition 2.2 (Mittag-Leffler Function)
The two-parameter Mittag-Leffler function is defined by the series [29]:
$E_{\alpha,\beta}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{\Gamma(\alpha k+\beta)},\quad\alpha,\beta>0,\,z\in\mathbb{C}.$
This function generalizes the exponential function and is crucial for constructing solutions to fractional differential equations, playing a role analogous to the exponential in integer-order systems.
Definition 2.3 (Fractional Lyapunov Stability)
The equilibrium point $y^{*}$ of the fractional-order system $^{C}D_{t}^{\alpha}y(t)=f(t,y(t))$ is said to be Mittag-Leffler stable if
$\|y(t)-y^{*}\|\leq\left[m(y(0)-y^{*})E_{\alpha,1}(-\lambda t^{\alpha})\right]^{b},$
where $\alpha\in(0,1)$, $\lambda\geq 0$, $b>0$, $m(0)=0$, and $m(y)\geq 0$. Mittag-Leffler stability implies asymptotic stability [48].
Definition 2.4 (Physics-Informed Neural Network Loss)
For a system governed by $^{C}D_{t}^{\alpha}\mathbf{Y}(t)-\mathbf{F}(t,\mathbf{Y}(t))=0$, the loss function for a Physics-Informed Neural Network (PINN) $NN(t;\theta)$ approximating $\mathbf{Y}(t)$ is defined as:
$\mathcal{L}(\theta)=\mathcal{L}_{data}+\mathcal{L}_{physics},$
where $\mathcal{L}_{data}=\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}|NN(t_{i};\theta)-\mathbf{Y}_{i}|^{2}$ enforces data fidelity, and $\mathcal{L}_{physics}=\frac{1}{N_{c}}\sum_{j=1}^{N_{c}}|{}^{C}D_{t}^{\alpha}NN(t_{j};\theta)-\mathbf{F}(t_{j},NN(t_{j};\theta))|^{2}$ enforces the physical law described by the fractional differential equation.
Theorem 2.1 (Stability Condition for Fractional Systems)
Let $\alpha\in(0,1]$. An equilibrium point $y^{*}$ of the autonomous system $^{C}D_{t}^{\alpha}y(t)=f(y(t))$ is asymptotically stable if all eigenvalues $\lambda_{i}$ of the Jacobian matrix $J=\partial f/\partial y|_{y^{*}}$ satisfy the condition [48]:
$|\arg(\lambda_{i})|>\frac{\alpha\pi}{2},\quad\forall i.$
This condition generalizes the classical Routh-Hurwitz criterion to fractional-order systems.
3 Fractional-Order Ebola Transmission Dynamics
We formulate a Caputo fractional-order model ($\alpha\in(0,1]$) which incorporates the multi-pathway transmission, memory effects, and time-varying interventions of Ebola. The governing system is given as follows.
${}^{C}D_{t}^{\alpha}S(t)$ $=\Lambda-\frac{c\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}+\omega V-(v+\mu)S,$ (1)
${}^{C}D_{t}^{\alpha}V(t)$ $=vS-(1-\varepsilon)\frac{c\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}-(\mu+\omega)V,$
${}^{C}D_{t}^{\alpha}E(t)$ $=\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}-(\mu+\sigma)E,$
${}^{C}D_{t}^{\alpha}I_{s}(t)$ $=p\sigma E-(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s},$
${}^{C}D_{t}^{\alpha}I_{a}(t)$ $=(1-p)\sigma E-(\gamma_{a}+\mu)I_{a},$
${}^{C}D_{t}^{\alpha}H(t)$ $=h_{s}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H,$
${}^{C}D_{t}^{\alpha}D(t)$ $=\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D,$
${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R.$
The force of infection is given by:
$\lambda=\frac{c\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N},$
and for notational simplicity, we define:
$q_{0}$ $=(v+\mu),$ $q_{1}$ $=(\mu+\omega),$ $q_{2}$ $=(\mu+\sigma),$
$q_{3}$ $=(\gamma_{s}+\delta_{s}+h_{s}+\mu),$ $q_{4}$ $=(\gamma_{a}+\mu),$ $q_{5}$ $=(\gamma_{h}+\delta_{h}+\mu),$
$q_{6}$ $=\mu_{d}.$
The model parameters include: natural mortality ($\mu$) and recruitment ($\Lambda$); vaccination rate $v(t)$ with waning immunity ($\omega$); and force of infection $\lambda(t)$ driving exposures. Exposed individuals ($E$) progress to infectious states at rate $\sigma$, splitting into symptomatic ($I_{s}$, proportion $p$) and asymptomatic ($I_{a}$, 1p1-p) pathways. Symptomatic cases face hospitalization ($h_{s}$), disease-induced mortality ($\delta_{s}$), or recovery ($\gamma_{s}$). Asymptomatic and hospitalized individuals recover at rates $\gamma_{a}$ and $\gamma_{h}$ respectively, with hospital mortality ($\delta_{h}$). Deceased individuals ($D$) transmit infection until safe burial ($\mu_{d}$). Transmission coefficients $\beta$, $\eta_{a}$, and $\eta_{d}$ govern infectivity from symptomatic, asymptomatic, and post-mortem sources. The system is governed by:
${}^{C}D_{t}^{\alpha}S(t)$ $=\Lambda-\lambda S+\omega V-q_{0}S,$ (2)
${}^{C}D_{t}^{\alpha}V(t)$ $=vS-(1-\varepsilon)\lambda V-q_{1}V,$
${}^{C}D_{t}^{\alpha}E(t)$ $=\lambda[S+(1-\varepsilon)V]-q_{2}E,$
${}^{C}D_{t}^{\alpha}I_{s}(t)$ $=p\sigma E-q_{3}I_{s},$
${}^{C}D_{t}^{\alpha}I_{a}(t)$ $=(1-p)\sigma E-q_{4}I_{a},$
${}^{C}D_{t}^{\alpha}H(t)$ $=h_{s}I_{s}-q_{5}H,$
${}^{C}D_{t}^{\alpha}D(t)$ $=\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D,$
${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R.$
The fractional-order system (2) is analyzed with the following initial conditions:
$S(0)=S_{0}\geq 0,\quad V(0)=V_{0}\geq 0,\quad E(0)=E_{0}\geq 0,\quad I_{s}(0)=I_{s_{0}}\geq 0,$
$I_{a}(0)=I_{a_{0}}\geq 0,\quad H(0)=H_{0}\geq 0,\quad D(0)=D_{0}\geq 0,\quad R(0)=R_{0}\geq 0.$
In this Caputo fractional formulation, the system in (2) captures the dynamics of Ebola transmission in an eight-dimensional frame. This construction involves ‘epidemic’ components such as: Susceptible ($S$), Vaccinated ($V$), Exposed ($E$), Symptomatically ($I_{s}$) and Asymptomatically ($I_{a}$) infectious, Hospitalized ($H$), deceased ($D$), and Recovery ($R$) stages, thus integrating critical structure of underlying dynamics. Furthermore, the model participates the fractional-order system of differential equations, thus inserting memory effects.
[…]
Figure 1: Transmission architecture of the fractional Ebola model.
Table 1: Epidemiological state variables and parameter estimates for the fractional-order Ebola model.
Variable Epidemiological Definition Estimated Value
$S(t)$ Susceptible individuals
$V(t)$ Vaccinated individuals
$E(t)$ Exposed (infected, not infectious)
$I_{s}(t)$ Symptomatic infectious individuals
$I_{a}(t)$ Asymptomatic infectious individuals
$H(t)$ Hospitalized/isolated individuals
$D(t)$ Deceased (infectious pre-burial)
$R(t)$ Recovered individuals
$N(t)$ Total living population
$\beta$ Transmission rate 0.287
$\eta_{a}$ Relative infectiousness (asymptomatic) 0.523
$\eta_{d}$ Relative infectiousness (deceased) 0.734
$\sigma$ Incubation rate (day-1) 0.094
$p$ Proportion symptomatic 0.712
$\gamma_{s}$ Recovery rate (symptomatic) 0.068
$\delta_{s}$ Death rate (symptomatic) 0.103
$h_{s}$ Hospitalization rate 0.312
$\gamma_{a}$ Recovery rate (asymptomatic) 0.089
$\delta_{h}$ Death rate (hospitalized) 0.067
Table 2: Parameter ranges and biological interpretation for the fractional-order Ebola model.
Symbol Biological Interpretation Range Source
$\Lambda$ Recruitment rate 100-1000 [1]
$\mu$ Natural mortality rate $3.5\times 10^{-5}$ [51]
$c(t)$ Contact rate function 0.4-1.2 [9]
$v(t)$ Vaccination rate 0.005-0.08 [52]
$\varepsilon$ Vaccine efficacy 0.85-0.95 [52]
$\omega$ Waning immunity rate 1/365-1/270 [53]
$\mu_{d}$ Burial rate 0.15-0.35 [54]
$\alpha$ Fractional order 0.75-0.95 [29]
Figure 2: Model validation: Curve fitting to confirmed Ebola cases ($R^{2}=0.96$, RMSE = 23.45).
4 Mathematical Analysis
4.1 Well-Posedness of the Model
We establish the mathematical well-posedness of the fractional-order Ebola system through existence and uniqueness analysis.
Theorem 4.1 (Existence and Uniqueness)
Given any $\alpha\in(0,1]$, and initial conditions
$\mathbf{X}_{0}=(S_{0},V_{0},E_{0},I_{s0},I_{a0},H_{0},D_{0},R_{0})\in\mathbb{R}_{+}^{8}$, the fractional-order system (2) has the unique solution $\mathbf{X}(t)=(S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t))$ for all $t\geq 0$.
Proof.  Define the state vector $\mathbf{X}(t)=(S,V,E,I_{s},I_{a},H,D,R)^{T}$ and express the system as:
$^{C}D_{t}^{\alpha}\mathbf{X}(t)=\mathbf{F}(t,\mathbf{X}(t)),\quad\mathbf{X}(0)=\mathbf{X}_{0}.$
The equivalent Volterra integral formulation is:
$\mathbf{X}(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds.$ (3)
Assume the operator $\mathcal{T}:C([0,T],\mathbb{R}^{8})\to C([0,T],\mathbb{R}^{8})$ defined by:
$(\mathcal{T}\mathbf{X})(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds.$
For $\mathbf{X},\mathbf{Y}\in C([0,T],\mathbb{R}^{8})$ with $\|\mathbf{X}\|=\sup_{t\in[0,T]}\|\mathbf{X}(t)\|_{\infty}$, the Lipschitz continuity of $\mathbf{F}$ follows from bounded parameters and finite state variables over compact intervals. Specifically, there exists $L>0$ such that:
$\|\mathbf{F}(t,\mathbf{X})-\mathbf{F}(t,\mathbf{Y})\|_{\infty}\leq L\|\mathbf{X}-\mathbf{Y}\|_{\infty}.$
The contraction estimate yields:
$\|\mathcal{T}\mathbf{X}-\mathcal{T}\mathbf{Y}\|\leq\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}\|\mathbf{X}-\mathbf{Y}\|.$
Selecting $\tau>0$ such that $\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}<1$ ensures $\mathcal{T}$ is a contraction on $C([0,\tau],\mathbb{R}^{8})$. Banach’s fixed point theorem guarantees a unique local solution $\mathbf{X}^{*}\in C([0,\tau],\mathbb{R}^{8})$.
Global extension to $0,\infty)$ follows from the linear growth condition $\|\mathbf{F}(t,\mathbf{X})\|_{\infty}\leq M(1+\|\mathbf{X}\|_{\infty})$ for some $M>0$, which precludes finite-time blow-up. Iterative continuation establishes global existence and uniqueness. $\square$
4.2 Invariant Region and Attractivity
The fractional-order Ebola system (2) evolves within the biologically feasible region:
$\Omega=\left\{(S,V,E,I_{s},I_{a},H,D,R)\in\mathbb{R}^{8}_{\geq 0}:\quad D\leq\frac{(\delta_{s}+\delta_{h})}{\mu_{d}}\left(\frac{\Lambda}{\mu}+\epsilon\right),N_{L}\leq(\frac{\Lambda}{\mu}+\epsilon)\right\},$ (4)
for $\epsilon>0$, where $N_{L}=S+V+E+I_{s}+I_{a}+H+R$ denotes the living population.
Theorem 4.2 (Invariance and Attractivity)
The region $\Omega$ is:
  1. (i)
    Positively invariant under (2),
  2. (ii)
    Globally attractive in $\mathbb{R}^{8}_{\geq 0}$.
Proof.  (i) Positivity and Invariance: Applying the fractional positivity lemma [29], we verify boundary behavior:
${}^{C}D_{t}^{\alpha}S|_{S=0}=\Lambda+\omega V\geq 0,\quad{}^{C}D_{t}^{\alpha}V|_{V=0}=vS\geq 0,$
${}^{C}D_{t}^{\alpha}E|_{E=0}=\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}\geq 0,$
${}^{C}D_{t}^{\alpha}I_{s}|_{I_{s}=0}=p\sigma E\geq 0,\quad{}^{C}D_{t}^{\alpha}I_{a}|_{I_{a}=0}=(1-p)\sigma E\geq 0,$
${}^{C}D_{t}^{\alpha}H|_{H=0}=h_{s}I_{s}\geq 0,\quad{}^{C}D_{t}^{\alpha}D|_{D=0}=\delta_{s}I_{s}+\delta_{h}H\geq 0,$
${}^{C}D_{t}^{\alpha}R|_{R=0}=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H\geq 0.$
Non-negativity of all boundary fluxes ensures $\mathbb{R}_{+}^{8}$ is invariant. (ii) Boundedness and Attractivity: For the living population $N_{L}$:
$^{C}D_{t}^{\alpha}N_{L}=\Lambda-\mu N_{L}-(\delta_{s}I_{s}+\delta_{h}H)\leq\Lambda-\mu N_{L}.$
Laplace transformation yields:
$N_{L}(s)\leq\frac{\Lambda}{s(s^{\alpha}+\mu)}+N_{L}(0)\frac{s^{\alpha-1}}{s^{\alpha}+\mu},$
with inverse transform:
$N_{L}(t)\leq N_{L}(0)E_{\alpha,1}(-\mu t^{\alpha})+\Lambda t^{\alpha}E_{\alpha,\alpha+1}(-\mu t^{\alpha}).$
As $t\to\infty$, $E_{\alpha,1}(-\mu t^{\alpha})\to 0$ and $t^{\alpha}E_{\alpha,\alpha+1}(-\mu t^{\alpha})\to 1/\mu$, giving:
$\limsup_{t\to\infty}N_{L}(t)\leq\frac{\Lambda}{\mu}.$
For deceased individuals:
$^{C}D_{t}^{\alpha}D\leq(\delta_{s}+\delta_{h})\left(\frac{\Lambda}{\mu}+\epsilon\right)-\mu_{d}D,$
implying:
$\limsup_{t\to\infty}D(t)\leq\frac{\delta_{s}+\delta_{h}}{\mu_{d}}\left(\frac{\Lambda}{\mu}+\epsilon\right).$
Thus, all trajectories are eventually absorbed into $\Omega$. $\square$
4.3 EVD-Free Equilibrium
The state wherein a population exists without any infections of the Ebola virus is referred to the Ebola-free equilibrium (DFE), at which point all infected compartments have completely dissipated:
$E=I_{s}=I_{a}=H=D=0.$
The system (2) reduces to:
$0$ $=\Lambda-q_{0}S+\omega V,$ (5)
$0$ $=vS-q_{1}V,$ (6)
$0$ $=-\mu R.$ (7)
From (7), $R^{*}=0$. Equations (5) and (6) form the linear system:
$q_{0}S-\omega V$ $=\Lambda,$ (8)
$-vS+q_{1}V$ $=0.$ (9)
The coefficient matrix
$A=\begin{pmatrix}q_{0}&-\omega\\ -v&q_{1}\end{pmatrix}$
has determinant
$\det(A)=q_{0}q_{1}-\omega v=\mu(q_{1}+v)>0,$
ensuring a unique solution. Applying Cramer’s rule:
$S^{*}=\frac{\Lambda q_{1}}{\mu(q_{1}+v)},\quad V^{*}=\frac{\Lambda v}{\mu(q_{1}+v)}.$
Thus, the disease-free equilibrium is:
$\mathcal{E}_{0}=\left(\frac{\Lambda q_{1}}{\mu(q_{1}+v)},\frac{\Lambda v}{\mu(q_{1}+v)},0,0,0,0,0,0\right).$
This biologically consistent DFE provides the foundation for stability analysis and reproduction number derivation.
[…]
$\mathcal{R}_{0}$ is computed via the next generation matrix method as expressed in [55]. In the case of the infected subsystem $X=(E,I_{s},I_{a},H,D)^{T}$, the new infection ($F$) and transition ($V$) matrices are as follows:
$F=\begin{pmatrix}0&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\end{pmatrix},\quad V=\begin{pmatrix}q_{2}&0&0&0&0\\ -p\sigma&q_{3}&0&0&0\\ -(1-p)\sigma&0&q_{4}&0&0\\ 0&-h_{s}&0&q_{5}&0\\ 0&-\delta_{s}&0&-\delta_{h}&q_{6}\end{pmatrix}$
where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$ with $N^{*}=\frac{\Lambda}{\mu}$.
The inverse transition matrix is:
$V^{-1}=\begin{pmatrix}\frac{1}{q_{2}}&0&0&0&0\\ \frac{p\sigma}{q_{2}q_{3}}&\frac{1}{q_{3}}&0&0&0\\ \frac{(1-p)\sigma}{q_{2}q_{4}}&0&\frac{1}{q_{4}}&0&0\\ \frac{p\sigma h_{s}}{q_{2}q_{3}q_{5}}&\frac{h_{s}}{q_{3}q_{5}}&0&\frac{1}{q_{5}}&0\\ \frac{p\sigma(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{2}q_{3}q_{5}q_{6}}&\frac{\delta_{s}q_{5}+\delta_{h}h_{s}}{q_{3}q_{5}q_{6}}&0&\frac{\delta_{h}}{q_{5}q_{6}}&\frac{1}{q_{6}}\end{pmatrix}$
The spectral radius of the next generation matrix is $K=FV^{-1}$ yields:
$\mathcal{R}_{0}=\beta\Psi\left[\frac{p\sigma}{q_{2}}\left(\frac{1}{q_{3}}+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)\sigma}{q_{2}q_{4}}\right]$
Substituting $\Psi$ provides the explicit form:
$\mathcal{R}_{0}=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\left[\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}\right]$
This threshold quantity governs disease invasion and extinction dynamics.
(a) Cumulative confirmed cases progression
(b) Cumulative deaths progression
(c) Active cases progression
(d) Comparative analysis of all metrics
Figure 3: Ebola Virus Disease outbreak analysis during the 2014-2016 West Africa outbreak showing: (a) cumulative confirmed cases, (b) cumulative deaths, (c) active cases, and (d) comparative analysis of all outbreak metrics.
4.5 Critical Vaccination Coverage
The critical vaccination coverage defines the minimum population immunity required for disease elimination, occurring when $\mathcal{R}_{0}=1$. Beginning with the reproduction number:
$\mathcal{R}_{0}(v)=\mathcal{R}_{0}(0)\cdot\frac{q_{1}+(1-\varepsilon)v}{q_{1}+v}$
where $\mathcal{R}_{0}(0)=K$ represents the vaccination-free reproduction number. Setting $\mathcal{R}_{0}(v_{c})=1$ yields:
$1=\mathcal{R}_{0}(0)\cdot\frac{q_{1}+(1-\varepsilon)v_{c}}{q_{1}+v_{c}}$
Solving for the critical vaccination rate $v_{c}$:
$q_{1}+v_{c}$ $=\mathcal{R}_{0}(0)[q_{1}+(1-\varepsilon)v_{c}]$
$v_{c}-\mathcal{R}_{0}(0)(1-\varepsilon)v_{c}$ $=\mathcal{R}_{0}(0)q_{1}-q_{1}$
$v_{c}[1-\mathcal{R}_{0}(0)(1-\varepsilon)]$ $=q_{1}[\mathcal{R}_{0}(0)-1]$
$v_{c}=\frac{q_{1}[\mathcal{R}_{0}(0)-1]}{1-(1-\varepsilon)\mathcal{R}_{0}(0)}$
The critical vaccination coverage $p_{c}$, representing the proportion vaccinated at equilibrium, is:
$p_{c}=\frac{v_{c}}{q_{1}+v_{c}}=\frac{\mathcal{R}_{0}(0)-1}{\varepsilon\mathcal{R}_{0}(0)}$
Thus, the herd immunity threshold for Ebola elimination is:
$p_{c}=\max\left(0,1-\frac{1}{\varepsilon\mathcal{R}_{0}(0)}\right)$
This establishes the minimal vaccination coverage required to interrupt disease transmission.
Figure 4: 3D surface: Vaccination and waning immunity impact on $\mathcal{R}_{0}$. The surface shows how the basic reproduction number varies with vaccination rate ($v$) and waning immunity rate ($\omega$).
4.6 Local Stability Analysis of the EVD-Free Equilibrium
Theorem 4.3
When $\mathcal{R}_{0}<1$, the EVD-free equilibrium $\mathcal{E}_{0}$ of the system (2) is achieved and remains locally stable; however, it becomes unsteady and $\mathcal{R}_{0}>1$ will prove it to be unstable.
Proof.  The Jacobian matrix evaluated at $\mathcal{E}_{0}$ takes the block triangular form:
$J(\mathcal{E}_{0})=\begin{pmatrix}J_{11}&J_{12}\\ 0&J_{22}\end{pmatrix}$
where $J_{11}$ corresponds to the uninfected compartments $(S,V,R)$ and $J_{22}$ to the infected compartments $(E,I_{s},I_{a},H,D)$. The uninfected subsystem matrix is:
$J_{11}=\begin{pmatrix}-(v+\mu)&\omega&0\\ v&-(\mu+\omega)&0\\ 0&0&-\mu\end{pmatrix}$
with eigenvalues $\lambda_{1}=-\mu$, $\lambda_{2}=-\mu$, $\lambda_{3}=-(\mu+v+\omega)$. All eigenvalues satisfy $|\arg(\lambda_{i})|=\pi>\frac{\alpha\pi}{2}$ for $\alpha\in(0,1]$, confirming stability of the uninfected subsystem. The infected subsystem Jacobian is:
$J_{22}=\begin{pmatrix}-q_{2}&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ p\sigma&-q_{3}&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0\\ 0&h_{s}&0&-q_{5}&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}\end{pmatrix}$
where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$. We decompose $J_{22}=F-V$, where $F$ and $V$ are the next-generation matrices previously defined. The stability of $J_{22}$ is determined by the spectral radius $\rho(FV^{-1})=\mathcal{R}_{0}$. According to the fractional Routh-Hurwitz criteria and properties of Metzler matrices, when $\mathcal{R}_{0}<1$, all eigenvalues of $J_{22}$ have negative real parts and satisfy the fractional stability condition $|\arg(\lambda)|>\frac{\alpha\pi}{2}$.
Conversely, when $\mathcal{R}_{0}>1$, the spectral abscissa of $J_{22}$ is positive, indicating at least one eigenvalue with positive real part. By the fractional stability theorem, this establishes instability of the disease-free equilibrium. $\square$
4.7 Global Stability of EVD-Free Equilibrium
Theorem 4.4
When $\mathcal{R}_{0}<1$, the EVD-free equilibrium $\mathcal{E}_{0}$ of the system (2) is globally asymptotically stable in $\Omega$.
Proof.  Consider the function for the infected compartments already introduced in the above section.
$L(E,I_{s},I_{a},H,D)=E+AI_{s}+BI_{a}+CH+FD,$
where the parameters A, B, C, and F are positive and not arbitrary. To investigate the function, we follow the trajectories described in the system and compute the Caputo derivative along each trajectory:
${}^{C}D_{t}^{\alpha}L=$ $\left[[\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}]-q_{2}E\right]$
$+A\left(p\sigma E-q_{3}I_{s}\right)+B((1-p)\sigma E-q_{4}I_{a})$
$+C\left(h_{s}I_{s}-q_{5}H\right)+F(\delta_{s}I_{s}+\delta_{h}H-q_{6}D)$
Rearranging terms by compartment:
${}^{C}D_{t}^{\alpha}L=$ $\left[-q_{2}+Ap\sigma+B(1-p)\sigma\right]E$
$+\left[\frac{c\beta[S+(1-\varepsilon)V]}{N}-Aq_{3}+Ch_{s}+F\delta_{s}\right]I_{s}$
$+\left[\frac{c\beta\eta_{a}[S+(1-\varepsilon)V]}{N}-Bq_{4}\right]I_{a}$
$+\left[-Cq_{5}+F\delta_{h}\right]H$
$+\left[\frac{c\beta\eta_{d}[S+(1-\varepsilon)V]}{N}-Fq_{6}\right]D$
We select coefficients to eliminate positive terms:
$C$ $=\frac{F\delta_{h}}{q_{5}},\quad B=\frac{\beta\eta_{a}\Psi}{q_{4}},\quad F=\frac{\beta\eta_{d}\Psi}{q_{6}}$
$A$ $=\frac{\beta\Psi}{q_{3}}+\frac{Ch_{s}}{q_{3}}+\frac{F\delta_{s}}{q_{3}}=\frac{\beta\Psi}{q_{3}}+\frac{F\delta_{h}h_{s}}{q_{3}q_{5}}+\frac{F\delta_{s}}{q_{3}}$
Using the bound $\frac{S+(1-\varepsilon)V}{N}\leq\Psi$ valid throughout $\Omega$, we obtain:
${}^{C}D_{t}^{\alpha}L\leq$ $\left[-q_{2}+\sigma\left(\frac{p\beta\Psi}{q_{3}}+\frac{pF\delta_{h}h_{s}}{q_{3}q_{5}}+\frac{pF\delta_{s}}{q_{3}}+\frac{(1-p)\beta\eta_{a}\Psi}{q_{4}}\right)\right]E$
$+\beta\Psi\left[1-1\right]I_{s}+\beta\eta_{a}\Psi\left[1-1\right]I_{a}+\beta\eta_{d}\Psi\left[1-1\right]D$
Substituting $F=\frac{\beta\eta_{d}\Psi}{q_{6}}$ yields:
$^{C}D_{t}^{\alpha}L\leq q_{2}\left[-1+\mathcal{R}_{0}\right]E$
For $\mathcal{R}_{0}\leq 1$, we have $^{C}D_{t}^{\alpha}L\leq 0$ throughout $\Omega$, with equality if and only if $E=0$. When $E=0$, the system dynamics ensure:
${}^{C}D_{t}^{\alpha}I_{s}$ $=-q_{3}I_{s}\Rightarrow I_{s}\to 0$
${}^{C}D_{t}^{\alpha}I_{a}$ $=-q_{4}I_{a}\Rightarrow I_{a}\to 0$
${}^{C}D_{t}^{\alpha}H$ $=h_{s}I_{s}-q_{5}H\Rightarrow H\to 0$
${}^{C}D_{t}^{\alpha}D$ $=\delta_{s}I_{s}+\delta_{h}H-q_{6}D\Rightarrow D\to 0$
Thus, the maximal invariant set where $^{C}D_{t}^{\alpha}L=0$ is precisely $\mathcal{E}_{0}$. Application of the fractional LaSalle invariance principle establishes global asymptotic stability when $\mathcal{R}_{0}\leq 1$. For $\mathcal{R}_{0}>1$, instability follows from local analysis, while uniform persistence can be demonstrated via established methods in fractional dynamical systems [48]. $\square$
4.8 Endemic Equilibrium Analysis
Theorem 4.5 (Existence of Endemic Equilibrium)
The fractional-order Ebola system (2) admits a unique endemic equilibrium point $\mathcal{E}^{*}=(S^{*},V^{*},E^{*},I_{s}^{*},I_{a}^{*},H^{*},D^{*},R^{*})$ with all components strictly positive if and only if $\mathcal{R}_{0}>1$.
Proof.  At equilibrium, the time derivatives vanish. The force of infection is:
$\lambda^{*}=\frac{c\beta(I_{s}^{*}+\eta_{a}I_{a}^{*}+\eta_{d}D^{*})}{N^{*}}$
The equilibrium equations provide:
$0$ $=\Lambda-\lambda^{*}S^{*}+\omega V^{*}-q_{0}S^{*}$ (10)
$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-q_{1}V^{*}$ (11)
$0$ $=\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]-q_{2}E^{*}$ (12)
$0$ $=p\sigma E^{*}-q_{3}I_{s}^{*}$ (13)
$0$ $=(1-p)\sigma E^{*}-q_{4}I_{a}^{*}$ (14)
$0$ $=h_{s}I_{s}^{*}-q_{5}H^{*}$ (15)
$0$ $=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}-q_{6}D^{*}$ (16)
$0$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}-\mu R^{*}$ (17)
Expressing infected states in terms of $E^{*}$ as:
$I_{s}^{*}$ $=\frac{p\sigma}{q_{3}}E^{*}\ and\quad I_{a}^{*}=\frac{(1-p)\sigma}{q_{4}}E^{*}$
$H^{*}$ $=\frac{p\sigma h_{s}}{q_{3}q_{5}}E^{*}and\quad D^{*}=\frac{p\sigma}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)E^{*}$
The force of infection becomes:
$\lambda^{*}=\frac{c\beta\sigma E^{*}}{N^{*}}\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right]$
Define the composite parameter:
$M=c\beta\sigma\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right]$
yielding $\lambda^{*}=ME^{*}/N^{*}$. From the exposed compartment equilibrium:
$E^{*}=\frac{\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]}{q_{2}},$
Uniting these relations gives the key equation:
$1=\frac{M}{q_{2}N^{*}}[S^{*}+(1-\varepsilon)V^{*}]$
Solving the susceptible-vaccinated subsystem:
$S^{*}$ $=\frac{\Lambda(q_{1}+(1-\varepsilon)\lambda^{*})}{\Delta(\lambda^{*})},\quad V^{*}=\frac{\Lambda v}{\Delta(\lambda^{*})}$
where $\Delta(\lambda^{*})=\left[(\omega+\mu+(1-\varepsilon)\lambda^{*})(\mu+v+\lambda^{*})\right]-\omega v$. This leads to $N^{*}=\Lambda/\mu$ hence the key equation becomes:
$1=\left[\frac{M}{q_{2}(\Lambda/\mu)}\right][S^{*}+(1-\varepsilon)V^{*}]$
At the disease-free state ($\lambda^{*}=0$), this expression equals $\mathcal{R}_{0}$. The monotonic decrease of $S^{*}+(1-\varepsilon)V^{*}]$ with $\lambda^{*}$ guarantees a unique positive solution $\lambda^{*}>0$ when $\mathcal{R}_{0}>1$, establishing the endemic equilibrium. $\square$
4.9 Local Stability Analysis of the Endemic Equilibrium
Theorem 4.6
Consider system (2). When $\mathcal{R}_{0}>1$, the endemic equilibrium $\mathcal{E}^{*}$ is locally asymptotically stable.
Proof.  The Jacobian matrix evaluated at $\mathcal{E}^{*}$ exhibits the block structure:
$J(\mathcal{E}^{*})=\begin{pmatrix}J_{11}&J_{12}\\ J_{21}&J_{22}\end{pmatrix}$
with components:
$J_{11}$ $=\begin{pmatrix}-(\lambda^{*}+q_{0})&\omega\\ v&-\lambda^{*}+\varepsilon\lambda^{*}-q_{1})\end{pmatrix}$,
$J_{12}$ $=\begin{pmatrix}0&-\beta S^{*}\Phi&-\beta\eta_{a}S^{*}\Phi&0&-\beta\eta_{d}S^{*}\Phi&0\\ 0&-\beta(1-\varepsilon)V^{*}\Phi&-\beta\eta_{a}(1-\varepsilon)V^{*}\Phi&0&-\beta\eta_{d}(1-\varepsilon)V^{*}\Phi&0\end{pmatrix}$,
$J_{21}$ $=\begin{pmatrix}\lambda^{*}&(1-\varepsilon)\lambda^{*}\\ 0&0\\ 0&0\\ 0&0\\ 0&0\\ 0&0\end{pmatrix}$,
$J_{22}$ $=\begin{pmatrix}-q_{2}&\beta\Psi\Phi&\beta\eta_{a}\Psi\Phi&0&\beta\eta_{d}\Psi\Phi&0\\ p\sigma&-q_{3}&0&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0&0\\ 0&h_{s}&0&-q_{5}&0&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}&0\\ 0&\gamma_{s}&\gamma_{a}&\gamma_{h}&0&-\mu\end{pmatrix}$
where $\Psi=S^{*}+(1-\varepsilon)V^{*}$, $\Phi=c/N^{*}$, and $\lambda^{*}$ denotes the equilibrium force of infection. The characteristic polynomial is an eighth-degree equation:
$P(\lambda)=\lambda^{8}+a_{7}\lambda^{7}+a_{6}\lambda^{6}+a_{5}\lambda^{5}+a_{4}\lambda^{4}+a_{3}\lambda^{3}+a_{2}\lambda^{2}+a_{1}\lambda+a_{0}=0$
Consider the Metzler matrix $Q=-J(\mathcal{E}^{*})$. For $\mathcal{R}_{0}>1$, $Q$ is irreducible and satisfies:
  • Diagonal positivity: $q_{ii}>0$
  • Diagonal dominance: $q_{ii}\geq\sum_{j\neq i}|q_{ij}|$
  • Existence of positive vector $v>0$ with $Qv>0$
According to the Perron-Frobenius theorem, the eigenvalues of $Q$ have positive real parts, and thus the eigenvalues of $J(\mathcal{E}^{*})=-Q$ have negative real parts. The fractional stability condition $|\arg(\lambda_{i})|>\alpha\pi/2$ is satisfied, because $|\arg(\lambda_{i})|=\pi>\alpha\pi/2$ for $\alpha\in(0,1]$. Stability conditions are confirmed via Routh-Hurwitz’s analysis technique if all the principal minors of the Hurwitz matrix associated with the system are strictly positive. This guarantees that all eigenvalues of the characteristic equation have negative real parts. The result stems from the fractional-order stability theorem [48], which proves that the system is asymptotically stable. $\square$
4.10 Global Stability of the Endemic Equilibrium
Theorem 4.7
For values of $\mathcal{R}_{0}$ greater than 1, the endemic equilibrium point $\mathcal{E}^{*}$ of system (2) is globally asymptotically stable in the interior of the invariant region $\Omega$.
Proof.  A Lyapunov function can be constructed with the use of the logarithmic Volterra-type form as follows:
$\mathcal{V}=$ $\left(S-S^{*}-S^{*}\ln\frac{S}{S^{*}}\right)+\left(V-V^{*}-V^{*}\ln\frac{V}{V^{*}}\right)$
$+K_{1}\left(E-E^{*}-E^{*}\ln\frac{E}{E^{*}}\right)+K_{2}\left(I_{s}-I_{s}^{*}-I_{s}^{*}\ln\frac{I_{s}}{I_{s}^{*}}\right)$
$+K_{3}\left(I_{a}-I_{a}^{*}-I_{a}^{*}\ln\frac{I_{a}}{I_{a}^{*}}\right)+K_{4}\left(H-H^{*}-H^{*}\ln\frac{H}{H^{*}}\right)$
$+K_{5}\left(D-D^{*}-D^{*}\ln\frac{D}{D^{*}}\right)+K_{6}\left(R-R^{*}-R^{*}\ln\frac{R}{R^{*}}\right)$
with $K_{1},K_{2},K_{3},K_{4},K_{5},K_{6}>0$ to be determined subsequently Capital, along the system trajectories, the Caputo derivative produces:
${}^{C}D_{t}^{\alpha}\mathcal{V}=$ $\left(1-\frac{S^{*}}{S}\right){}^{C}D_{t}^{\alpha}S+\left(1-\frac{V^{*}}{V}\right){}^{C}D_{t}^{\alpha}V+K_{1}\left(1-\frac{E^{*}}{E}\right){}^{C}D_{t}^{\alpha}E$
$+K_{2}\left(1-\frac{I_{s}^{*}}{I_{s}}\right){}^{C}D_{t}^{\alpha}I_{s}+K_{3}\left(1-\frac{I_{a}^{*}}{I_{a}}\right){}^{C}D_{t}^{\alpha}I_{a}+K_{4}\left(1-\frac{H^{*}}{H}\right){}^{C}D_{t}^{\alpha}H$
$+K_{5}\left(1-\frac{D^{*}}{D}\right){}^{C}D_{t}^{\alpha}D+K_{6}\left(1-\frac{R^{*}}{R}\right){}^{C}D_{t}^{\alpha}R$
Replacing the model equations while utilizing endemic equilibrium conditions:
$\Lambda$ $=\lambda^{*}S^{*}+\omega V^{*}-(\mu+v)S^{*}$
$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-(\omega+\mu)V^{*}$
$q_{2}E^{*}$ $=\lambda^{*}(S^{*}+(1-\varepsilon)V^{*})$
$q_{3}I_{s}^{*}$ $=p\sigma E^{*},\quad q_{4}I_{a}^{*}=(1-p)\sigma E^{*}$
$q_{5}H^{*}$ $=h_{s}I_{s}^{*},\quad\mu_{d}D^{*}=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}$
$\mu R^{*}$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}$
We take coefficients as:
$K_{1}$ $=1,\quad K_{2}=\frac{\lambda^{*}S^{*}}{q_{3}I_{s}^{*}},\quad K_{3}=\frac{\lambda^{*}S^{*}\eta_{a}}{q_{4}I_{a}^{*}},$
$K_{4}$ $=\frac{\lambda^{*}S^{*}\eta_{d}\delta_{h}}{q_{5}q_{6}H^{*}},\quad K_{5}=\frac{\lambda^{*}S^{*}\eta_{d}}{q_{6}D^{*}},\quad K_{6}=1$
By means of grouping the terms accordingly, we end up with the inequality:
$^{C}D_{t}^{\alpha}\mathcal{V}\leq\lambda^{*}S^{*}\left[6-\frac{S^{*}}{S}-\frac{S}{S^{*}}-\frac{EI_{s}^{*}}{E^{*}I_{s}}-\frac{I_{s}E^{*}}{I_{s}^{*}E}-\frac{DH^{*}}{D^{*}H}-\frac{HD^{*}}{H^{*}D}\right]+\Psi(S,V,E,I_{s},I_{a},H,D,R)$
where $\Psi$ indicates a negative definite function. From the arithmetic-geometric mean inequality:
$\frac{S^{*}}{S}+\frac{S}{S^{*}}+\frac{EI_{s}^{*}}{E^{*}I_{s}}+\frac{I_{s}E^{*}}{I_{s}^{*}E}+\frac{DH^{*}}{D^{*}H}+\frac{HD^{*}}{H^{*}D}\geq 6$
Equality is achieved only when $S=S^{*},E=E^{*},I_{s}=I_{s}^{*},H=H^{*},D=D^{*}$. We can conclude that $^{C}D_{t}^{\alpha}\mathcal{V}\leq 0$ holds true for all points in the feasible region $\Omega$. It is only at the endemic equilibrium $\mathcal{E}^{*}$ that $^{C}D_{t}^{\alpha}\mathcal{V}=0$. This allows us to conclude the global asymptotic stability of $\mathcal{E}^{*}$ using the fractional LaSalle invariance principle [48]. $\square$
Corollary 4.8
The fractional-order Ebola system (2) experiences a forward bifurcation at $\mathcal{R}_{0}=1$, without showing a backward bifurcation and without exhibiting multiple endemic equilibria.
Proof.  Global stability properties—disease-free equilibrium whenever $\mathcal{R}_{0}\leq 1$ and unique endemic equilibrium when $\mathcal{R}_{0}>1$—contradict the possibility of backward bifurcation. The epidemiological dynamics displays classical threshold behavior: disease elimination befalls when $\mathcal{R}_{0}\leq 1$ and $\mathcal{R}_{0}>1$ indicates endemic tenacity. $\square$
5 Sensitivity Analysis
5.1 Global Sensitivity Analysis and Parameter Influence
From the data, Latin Hypercube Sampling (LHS) estimates $\mathcal{R}_{0}$ and other critical epidemiological characteristics and examines the correlation of the parameters using PRCC. The PRCC examined global sensitivity for 10,000 samples within the reasonable biological boundaries. The outcomes are summarized in Table 3.
Table 3: Global sensitivity analysis: Parameter ranges and PRCC values for $\mathcal{R}_{0}$ and peak infection dynamics
Parameter Lower Bound Upper Bound PRCC ($\mathcal{R}_{0}$) PRCC (Peak $I_{s}$) p-value
$\beta$ 0.15 0.35 0.92 0.88 $p<0.001$
$\sigma$ 1/12 1/8 0.85 0.81 $p<0.001$
$\eta_{d}$ 0.4 0.9 0.68 0.65 $p<0.001$
pp 0.6 0.85 0.52 0.49 $p<0.001$
$\eta_{a}$ 0.3 0.7 0.41 0.38 $p<0.001$
$\varepsilon$ 0.7 0.95 -0.61 -0.58 $p<0.001$
$\gamma_{s}$ 1/16 1/12 -0.45 -0.42 $p<0.001$
$v$ 0.01 0.08 -0.35 -0.32 $p<0.001$
$\mu_{d}$ 0.1 0.3 -0.28 -0.25 $p<0.001$
$\alpha$ 0.7 0.95 -0.25 -0.23 $p<0.001$
$h_{s}$ 0.2 0.6 -0.21 -0.19 $p<0.001$
$\delta_{s}$ 0.05 0.12 0.15 0.18 $p=0.002$
5.2 Analytical Sensitivity Framework
The normalized sensitivity index quantifying $\mathcal{R}_{0}$ dependence on parameter $p$ is defined as:
$\Upsilon_{p}^{\mathcal{R}_{0}}=\frac{\partial\mathcal{R}_{0}}{\partial p}\times\frac{p}{\mathcal{R}_{0}}$ (18)
Theorem 5.1 (Analytical Sensitivity Indices)
The sensitivity indices for dominant parameters are given by:
Proof.  Analytical differentiation of the reproduction number yields.
$\text{Sensitivity indices for }\mathcal{R}_{0}:$ (19)
$\Upsilon_{\beta}^{\mathcal{R}_{0}}=+1$
$\Upsilon_{\sigma}^{\mathcal{R}_{0}}=1+\frac{\mathcal{R}_{0}^{(1)}}{\mathcal{R}_{0}}$
$\Upsilon_{\eta_{a}}^{\mathcal{R}_{0}}=\frac{\mathcal{R}_{0}^{(I_{a})}}{\mathcal{R}_{0}}$
$\Upsilon_{\eta_{d}}^{\mathcal{R}_{0}}=\frac{\mathcal{R}_{0}^{(D)}}{\mathcal{R}_{0}}$
$\Upsilon_{\varepsilon}^{\mathcal{R}_{0}}=-\frac{V^{*}}{S^{*}+(1-\varepsilon)V^{*}}$
$\Upsilon_{v}^{\mathcal{R}_{0}}=\frac{\beta\sigma}{\mathcal{R}_{0}(\Lambda/\mu)q_{2}}\cdot\frac{\partial}{\partial v}[S^{*}+(1-\varepsilon)V^{*}]\cdot K,$
The following equations illustrate the different components that contribute to the basic reproduction number:
$\mathcal{R}_{0}^{(I_{a})}$ $=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\cdot\frac{\eta_{a}(1-p)}{q_{4}}$ (20)
$\mathcal{R}_{0}^{(D)}$ $=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\cdot\frac{p\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}$ (21)
$\frac{\partial}{\partial v}[S^{*}+(1-\varepsilon)V^{*}]$ $=-\frac{\Lambda\varepsilon(\omega+\mu)}{\mu(q_{1}+v)^{2}}$ (22)
with the composite transmission factor given by
$K=\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}.$ (23)
From equations (20) and (21) one may discern the individual contributions of asymptomatic infections and post-mortem transmission to the total basic reproduction number, $\mathcal{R}_{0}$. $\square$
5.3 Sensitivity Analysis and Numerical Results
Remark 5.1 (Parameter Influence Hierarchy)
Analytical and PRCC sensitivity analyses establish the following parameter influence hierarchy on $\mathcal{R}_{0}$:
$\text{High Positive: }\beta>\sigma>\eta_{d}$
$\text{Moderate Positive: }p>\eta_{a}$
$\text{High Negative: }\varepsilon$
$\text{Moderate Negative: }\gamma_{s}>v$
$\text{Low Negative: }\mu_{d}>\alpha>h_{s}$
Classification follows PRCC thresholds: highly influential ($|\text{PRCC}|>0.7$), moderately influential ($0.5<|\text{PRCC}|\leq 0.7$), and less influential ($|\text{PRCC}|\leq 0.5$).
Incorporating targeted integrated strategies aimed at reducing transmission ($\beta$), targeted rapid detection ($\sigma$), and safe burial ($\eta_{d}$) achieve an 85% reduction in outbreaks according to numerical simulations which is a significant increase from vaccination-only strategies that only attain a 45% reduction.
(a) Threshold dynamics verification
(b) Parameter sensitivity analysis
(c) PRCC ranking results
Figure 5: Numerical simulations validating theoretical findings and quantifying parameter impacts
Table 4: Evidence-based intervention prioritization from sensitivity analysis
Priority Interventions Impact
Primary Contact reduction, Safe burial, Early detection 60-85% $\mathcal{R}_{0}$ reduction
Secondary High-efficacy vaccination, Treatment access 25-45% $\mathcal{R}_{0}$ reduction
Tertiary Mass vaccination, Hospital capacity 10-20% $\mathcal{R}_{0}$ reduction
Figure 6: Contours of parameter sensitivity: (a) $\beta,\sigma$ (most sensitive), (b) $\eta_{d},p$ (intermediate), (c) $\varepsilon,\gamma_{s}$ (least sensitive), (d) $\eta_{a},v$ (mixed sensitivity)
The results further affirm the developed theory by demonstrating very strong threshold behavior at $\mathcal{R}_{0}=1$. Such confirmation also supports the parameter sweep results which assert the dominance of the transmission parameters $\beta$, $\sigma$, and $\eta_{d}$, in the dynamics of the outbreaks. Contour plots illustrate the nonlinear interactions which add to the ever-growing complexity of the transmission dynamics of the Ebola virus.
The primary community-based switch strategy employed to control the spread of the virus focuses on the interruption of transmission. Thus, the safe burial practices endorsed by the World Health Organization are critical in a community’s efforts to disturb transmission. Evidence of memory in the Ebola virus, captured by the sophisticated model elements, is indicated by the strong effect of the fractional order $\alpha$. Such results, in turn, offer the mathematical foundation for controlled community-centric Ebola response which is often lacking. This model control will inform the aggressive response necessary for targeted intervention strategies to control the Ebola virus.
6 Optimal Control Analysis
6.1 Problem Formulation
In this section, we delineate four time-dependent control variables for a complete management of the Ebola virus. Control variables: $u_{1}$ protection [0,0.8], $u_{2}$ vaccination [0,0.15], $u_{3}$ treatment [0,0.8], $u_{4}$ burial [0,0.5] (day-1). The controlled fractional-order system now becomes:
${}^{C}D_{t}^{\alpha}S(t)$ $=\Lambda-(1-u_{1})\lambda S+\omega V-(\mu+u_{2})S,$ (24)
${}^{C}D_{t}^{\alpha}V(t)$ $=u_{2}S-(1-\varepsilon)(1-u_{1})\lambda V-(\mu+\omega)V,$
${}^{C}D_{t}^{\alpha}E(t)$ $=(1-u_{1})\lambda[S+(1-\varepsilon)V]-(\mu+\sigma)E,$
${}^{C}D_{t}^{\alpha}I_{s}(t)$ $=p\sigma E-(\gamma_{s}+\delta_{s}+u_{3}+\mu)I_{s},$
${}^{C}D_{t}^{\alpha}I_{a}(t)$ $=\sigma E(1-p)-(\gamma_{a}+\mu)I_{a},$
${}^{C}D_{t}^{\alpha}H(t)$ $=u_{3}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H,$
${}^{C}D_{t}^{\alpha}D(t)$ $=\delta_{s}I_{s}+\delta_{h}H-(u_{4}+\mu_{d})D,$
${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R.$
with modified force of infection $\lambda(t)=\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)/N$. The objective functional minimizes total cost over $[0,T]$:
$J(\mathbf{u})=\int_{0}^{T}\left[\sum_{i=1}^{4}A_{i}X_{i}(t)+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}(t)\right]dt$ (25)
Where $X=[I_{s},I_{a},H,D]$ indicates the infectious compartments, diseased burden costs are denoted by $A_{i}>0$, and the costs related to the implementation of controls are denoted as $B_{j}>0$ in quadratic form.
We look to find the optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that fulfill the requirements of:
$J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u})$ (26)
for the controlled system (24), where the admissible set is:
$\mathcal{U}=\left\{\mathbf{u}\in L^{\infty}(0,T):0\leq u_{i}(t)\leq u_{i}^{\max}\right\}$ (27)
6.2 Necessary Conditions for Optimality
Theorem 6.1 (Existence of Optimal Controls)
Let $J(\mathbf{u})$ be the objective functional relative to the controlled fractional-order system (24) with controllable initial conditions $\mathbf{X}(0)\in\mathbb{R}^{8}_{\geq 0}$, and with $\mathbf{u}\in\mathcal{U}$ being the admissible controls. An optimal control pair $(\mathbf{X}^{*},\mathbf{u}^{*})$ such that $J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u})$ exists.
Proof.  Existence has been demonstrated by verifying the fundamental conditions in [56]. The state space $\mathbb{R}^{8}_{\geq 0}$ is closed and the control set $\mathcal{U}$ is closed, convex and bounded (as per its definition). The system dynamics $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ has the required properties as follows:
The system dynamics $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ are continuous in $(t,\mathbf{X},\mathbf{u})$, is bounded such that $\|\mathbf{f}\|\leq C(1+\|\mathbf{X}\|)$, and is Lipschitz continuous $\|\mathbf{f}(\mathbf{X}_{1})-\mathbf{f}(\mathbf{X}_{2})\|\leq L\|\mathbf{X}_{1}-\mathbf{X}_{2}\|.$ The cost integrand $\mathcal{L}(t,\mathbf{X},\mathbf{u})$ exhibits convexity in $\mathbf{u}$ through quadratic terms and coercivity $\mathcal{L}\geq\frac{1}{2}\min B_{j}\|\mathbf{u}\|^{2}-K$.
The inequality $^{C}D_{t}^{\alpha}N(t)\leq\Lambda-\mu N$ leads to population boundedness, ensuring that $N(t)\leq\max\{N(0),\Lambda/\mu\}$, and thus the state trajectories are bounded. Together, these conditions guarantee the existence of an optimal control pair. $\square$
6.3 Pontryagin’s Maximum Principle for Fractional Systems
Theorem 6.2
Assuming that $(\mathbf{X}^{*},\mathbf{u}^{*})$ is optimal for the control problem, there exist absolutely continuous adjoint variables satisfying the system of fractional differential equations:
${}^{C}D_{t}^{\alpha}\lambda_{S}$ $=-\frac{\partial H}{\partial S},$ ${}^{C}D_{t}^{\alpha}\lambda_{V}$ $=-\frac{\partial H}{\partial V},$ ${}^{C}D_{t}^{\alpha}\lambda_{E}$ $=-\frac{\partial H}{\partial E},$
${}^{C}D_{t}^{\alpha}\lambda_{I_{s}}$ $=-\frac{\partial H}{\partial I_{s}},$ ${}^{C}D_{t}^{\alpha}\lambda_{I_{a}}$ $=-\frac{\partial H}{\partial I_{a}},$ ${}^{C}D_{t}^{\alpha}\lambda_{H}$ $=-\frac{\partial H}{\partial H},$
${}^{C}D_{t}^{\alpha}\lambda_{D}$ $=-\frac{\partial H}{\partial D},$ ${}^{C}D_{t}^{\alpha}\lambda_{R}$ $=-\frac{\partial H}{\partial R},$
The optimal control $\mathbf{u}^{*}(t)$ is described by
$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}\mathcal{H}(t,\mathbf{x}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{a.e. on }[0,T],$
supplemented with the transversality condition
$\bm{\lambda}(T)=\mathbf{0},$
with $\bm{\lambda}=(\lambda_{S},\lambda_{V},\lambda_{E},\lambda_{I_{s}},\lambda_{I_{a}},\lambda_{H},\lambda_{D},\lambda_{R})^{\top}$.
Proof.  The proof follows the variational approach for fractional optimal control systems. Consider a perturbation $\mathbf{u}^{\epsilon}=\mathbf{u}^{*}+\epsilon\mathbf{v}$ with $\mathbf{v}\in L^{\infty}([0,T],\mathbb{R}^{4})$ such that $\mathbf{u}^{\epsilon}\in\mathcal{U}$ for sufficiently small $\epsilon>0$. Let $\mathbf{X}^{\epsilon}$ denote the corresponding state trajectory.
The first variation of the objective functional is given by:
$\delta J=\lim_{\epsilon\to 0}\frac{J(\mathbf{u}^{\epsilon})-J(\mathbf{u}^{*})}{\epsilon}=\int_{0}^{T}\left[\frac{\partial\mathcal{L}}{\partial\mathbf{X}}\cdot\delta\mathbf{X}+\frac{\partial\mathcal{L}}{\partial\mathbf{u}}\cdot\mathbf{v}\right]dt,$
where $\delta\mathbf{X}$ satisfies the linearized state equation derived from the controlled system. We have defined adjoint variables $\bm{\lambda}$ which are defined by the following fractional differential equations:
$^{C}D_{t}^{\alpha}\bm{\lambda}=-\frac{\partial H}{\partial\mathbf{X}},\quad\bm{\lambda}(T)=\mathbf{0}.$
Using integration by parts for Caputo fractional derivatives leads to:
$\delta J=\int_{0}^{T}\frac{\partial H}{\partial\mathbf{u}}\cdot\mathbf{v}dt.$
For the problem to be optimally solved, it is necessary to have $\delta J\geq 0$ for every admissible variations $\mathbf{v}$, which leads to:
$\frac{\partial H}{\partial\mathbf{u}}\cdot\mathbf{v}\geq 0\quad\text{a.e. on }[0,T].$
From this inequality, we derive the pointwise minimization condition. Consolidating the minimization condition, we have:
$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}H(t,\mathbf{X}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{a.e.}$
We have thus completed the derivation of necessary optimality conditions. $\square$
6.4 Optimal Control Characterization
Theorem 6.3
The optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that minimize the Hamiltonian are characterized by:
$u_{1}^{*}(t)$ $=\min\left\{u_{1}^{\max},\max\left\{0,\frac{[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)]\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{B_{1}N}\right\}\right\},$
$u_{2}^{*}(t)$ $=\min\left\{u_{2}^{\max},\max\left\{0,\frac{(\lambda_{S}-\lambda_{V})S}{B_{2}}\right\}\right\},$
$u_{3}^{*}(t)$ $=\min\left\{u_{3}^{\max},\max\left\{0,\frac{(\lambda_{I_{s}}-\lambda_{H})I_{s}}{B_{3}}\right\}\right\},$
$u_{4}^{*}(t)$ $=\min\left\{u_{4}^{\max},\max\left\{0,\frac{\lambda_{D}D}{B_{4}}\right\}\right\},$
where all state and adjoint variables follow the optimal trajectory.
Proof.  The characterization derives from solving the first-order optimality conditions $\partial H/\partial u_{i}=0$ for each control variable. For personal protection control $u_{1}^{*}$, the optimality condition yields:
$\frac{\partial H}{\partial u_{1}}=B_{1}u_{1}+\left[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-(S+(1-\varepsilon)V)\lambda_{E}\right]\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}=0.$
Solving algebraically provides the unconstrained optimum:
$u_{1}=\frac{[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)]\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{B_{1}N}.$
The last description arises from the projection onto the last interval $0,u_{1}^{\max}]$. For vaccination control $u_{2}^{*}$, the condition:
$\frac{\partial H}{\partial u_{2}}=B_{2}u_{2}+(\lambda_{V}-\lambda_{S})S=0$
gives $u_{2}=(\lambda_{S}-\lambda_{V})S/B_{2}$, with projection ensuring admissibility. For treatment control $u_{3}^{*}$, the optimality condition:
$\frac{\partial H}{\partial u_{3}}=B_{3}u_{3}+(\lambda_{H}-\lambda_{I_{s}})I_{s}=0$
yields $u_{3}=(\lambda_{I_{s}}-\lambda_{H})I_{s}/B_{3}$. For burial control $u_{4}^{*}$, the condition:
$\frac{\partial H}{\partial u_{4}}=B_{4}u_{4}-\lambda_{D}D=0$
provides $u_{4}=\lambda_{D}D/B_{4}$. The satisfying all controls within set projection established within the admis round case the set $\mathcal{U}$ which fulfills the distinctive classification. $\square$
6.5 Adjoint System Specification
Theorem 6.4
The fractional differential equations outlined below are satisfied by the adjoint variables:
${}^{C}D_{t}^{\alpha}\lambda_{S}$ $=(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(V+E+I_{s}+I_{a}+H+R)}{N^{2}}\right]$
$\quad+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$\quad+\lambda_{S}(\mu+u_{2})-\lambda_{V}u_{2},$ (28)
${}^{C}D_{t}^{\alpha}\lambda_{V}$ $=-\lambda_{S}\omega+\lambda_{V}(\omega+\mu)+(\lambda_{S}-\lambda_{E})(1-u_{1})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$\quad+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(S+E+I_{s}+I_{a}+H+R)}{N^{2}}\right],$ (29)
${}^{C}D_{t}^{\alpha}\lambda_{E}$ $=\lambda_{E}(\sigma+\mu)-\lambda_{I_{s}}p\sigma-(1-p)\lambda_{I_{a}}\sigma$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$\quad+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right],$ (30)
${}^{C}D_{t}^{\alpha}\lambda_{I_{s}}$ $=-A_{1}+\lambda_{I_{s}}(\gamma_{s}+\delta_{s}+u_{3}+\mu)-\lambda_{H}u_{3}-\lambda_{D}\delta_{s}-\lambda_{R}\gamma_{s}$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\frac{\beta S}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right]$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\frac{\beta V}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right],$ (31)
${}^{C}D_{t}^{\alpha}\lambda_{I_{a}}$ $=-A_{2}+\lambda_{I_{a}}(\gamma_{a}+\mu)-\lambda_{R}\gamma_{a}$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\frac{\beta S\eta_{a}}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right]$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\frac{\beta V\eta_{a}}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right],$ (32)
${}^{C}D_{t}^{\alpha}\lambda_{H}$ $=-A_{3}+\lambda_{H}(\gamma_{h}+\delta_{h}+\mu)-\lambda_{D}\delta_{h}-\lambda_{R}\gamma_{h}$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right],$ (33)
${}^{C}D_{t}^{\alpha}\lambda_{D}$ $=-A_{4}+\lambda_{D}(u_{4}+\mu_{d})$
$\quad+(\lambda_{S}-\lambda_{E})(1-u_{1})\frac{\beta S\eta_{d}}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right]$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\frac{\beta V\eta_{d}}{N}\left[1-\frac{I_{s}+\eta_{a}I_{a}+\eta_{d}D}{N}\right],$ (34)
${}^{C}D_{t}^{\alpha}\lambda_{R}$ $=\lambda_{R}\mu+(\lambda_{S}-\lambda_{E})(1-u_{1})\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}$
$\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right].$ (35)
Proof.  The calculation of the partial derivative of the Hamiltonian yields the equations of the adjoint each corresponding to the state variable of consideration. For $\lambda_{S}$, we have:
$\frac{\partial\mathcal{H}}{\partial S}=$ $(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(V+E+I_{s}+I_{a}+H+R)}{N^{2}}\right]$
$+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$+\lambda_{S}(\mu+u_{2})-\lambda_{V}u_{2}.$
Thus, (28) follows from $^{C}D_{t}^{\alpha}\lambda_{S}=-\partial\mathcal{H}/\partial S$. For $\lambda_{V}$, the derivative gives:
$\frac{\partial\mathcal{H}}{\partial V}=$ $-\lambda_{S}\omega+\lambda_{V}(\omega+\mu)$
$+(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
$+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(S+E+I_{s}+I_{a}+H+R)}{N^{2}}\right],$
which gives (29) under $^{C}D_{t}^{\alpha}\lambda_{V}=-\partial\mathcal{H}/\partial V$. The remaining adjoint equations (30)–(35) can be derived in a similar manner by calculating the respective partial derivatives $\partial\mathcal{H}/\partial x_{i}$ and using the fundamental relationship $^{C}D_{t}^{\alpha}\lambda_{i}=-\partial\mathcal{H}/\partial x_{i}$. The arrangement continues to embody the epidemiological relations where the transmission mechanisms are influenced by the cosate differentials and the passage rates determined by the compartments. $\square$
Remark 6.1
Each part of the adjoint system can be seen recursively decomposed in the sequence structured by the equations $\lambda_{R}$, $\lambda_{H}$, and $\lambda_{D}$, which become independent once the other adjoint variables are fixed. This characteristic feature of the system corresponds to the compartmentalized configuration of the epidemiological dynamics and facilitates the computation of an efficient numerical solution.
6.6 Numerical Results of Optimal Control
Figure 7: Optimal Control Analysis : (a) Algorithm Convergence, (b) Control Path, (c) Cost Control Effectiveness, (d) Control of Epidemics and Containment
Using Python, NumPy, and SciPy, the faster forward-backward sweep algorithm achieved convergence, exhibiting epidemic containment (peak: 170.0, final: 0.0) in 50 iterations within 30 seconds, and at a cost of 8750\$ Of the burial controls, which were the most cost-effective at 4\$/case, treatment with a 92% reduction in transmission, sequentially distilled intervention timing.
7 Efficiency Evaluation of Optimal Interventions
The numerical simulation of the fractional-order Ebola system employs the RK45 method. This method features adaptive step-size control which allows for a proper trade-off between precision and computational efficiency.
Algorithm 1 Runge-Kutta-Fehlberg (RKF45)
[…]
3:Initialize: $t\leftarrow t_{0}$, $\mathbf{y}\leftarrow\mathbf{y}_{0}$
4:Set parameters: $h_{min}\leftarrow 10^{-8}$, $h_{max}\leftarrow 0.1$, $h\leftarrow 0.01$
5:while $t<T$ do
6:  Compute RK coefficients $\mathbf{k}_{1}$ through $\mathbf{k}_{6}$
7:  Calculate 4th/5th order solutions $\mathbf{y}_{4}$, $\mathbf{y}_{5}$
8:  Estimate error $\delta=\|\mathbf{y}_{5}-\mathbf{y}_{4}\|$
9:  Compute optimal step $h_{opt}=0.9h(\epsilon_{rel}/\delta)^{0.2}$
10:  if $\delta\leq\epsilon_{rel}$ then
11:   Accept step: $t\leftarrow t+h$, $\mathbf{y}\leftarrow\mathbf{y}_{5}$
12:   Update $h\leftarrow\min(h_{max},\max(h_{min},h_{opt}))$
13:  else
14:   Reject step: $h\leftarrow\max(h_{min},h_{opt})$
15:  end if
16:end while
17:return Interpolated solution
The algorithm maintains 4th-order accuracy with 5th-order error control, employing adaptive step-size selection based on local truncation error estimates. Implementation uses relative tolerance $\epsilon_{rel}=10^{-6}$ and absolute tolerance $\epsilon_{abs}=10^{-8}$ for high-precision epidemiological simulations.
7.1 Results and Discussion of Intervention Strategies
Figure 8: Dynamics in compartments under Treatment Intervention Strategy ($u_{3}\neq 0$)
Figure 9: Foundational epidemiological principles regarding the impact of the treatment intervention
Figure 10: Effectiveness assessment of treatment strategy
Figure 11: compartmatic dynamics with safe burial intervention $u_{4}\neq 0$
Figure 12: Comparison of key metrics for safe burial intervention.
Figure 13: Analysis of Safeguard Burial Approach Efficiency
Figure 14: Dynamics of compartments for the strategy of instantaneous vaccination and treatment ($u_{2}\neq 0,u_{3}\neq 0$)
Figure 15: Key Metrics for the combined treatment and vaccination
Figure 16: Effectiveness assessment of combined vaccination and treatment
Figure 17: Compartmental dynamics for combined treatment and safe burial strategy ($u_{3}\neq 0,u_{4}\neq 0$)
Figure 18: Key metrics comparison for treatment-safe burial combination
Figure 19: Effectiveness assessment of treatment and safe burial strategy
Figure 20: Compartmental dynamics for three-intervention combination ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0$)
Figure 21: Key metrics for personal protection, vaccination, and treatment combination
Figure 22: Effectiveness assessment of three-intervention strategy
Figure 23: Compartmental dynamics for vaccination, treatment, and safe burial combination ($u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)
Figure 24: Key metrics comparison for clinical and preventive combination
Figure 25: Effectiveness assessment of vaccination, treatment, and safe burial strategy
Figure 26: Compartmental dynamics under comprehensive intervention strategy ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)
Figure 27: Key metrics for comprehensive intervention approach
Figure 28: Effectiveness assessment of comprehensive strategy
7.2 Discussion
Evaluating strategies for controlling pandemics has unequally advanced in mortality reduction and has repeatedly failed in controlling transmission. In each case studied, no intervention profoundly altered the fundamental dynamics of spread, evidenced by the persistent values for peak infection rates and, in controlling the intervention to the baseline, a net nonexistence of infection post intervention. Mortality figures, on the other hand, were markedly disparate with strategies demonstrating either complementary or overlapping impacts when aggregated.
The recorded treatment intervention ($u_{3}\neq 0$) from integration of treatment with other controlling appliances yielded a 60.4% decrease in final mortality, and the safe burial ($u_{4}\neq 0$) showed 73.8% reduction in final mortality, thereby demonstrating the safe burial of deceased individuals treatment in burial both alone and in combination. The rationale for the incorporated benefits from Figures 1013 compartmental dynamics are on a shift in disease with no less than transmission chain breakage which possibly led to the two outcomes being flawed.
There were significant interactions when different types of interventions were used together. From figures $16$ vaccination-treatment combination ($u_{2}\neq 0,u_{3}\neq 0$) achieved a 60.4% reduction in mortality which is the same as treatment done alone, therefore, illustrating that vaccination is not additive. In contrast, treatment with safe burial ($u_{3}\neq 0,u_{4}\neq 0$) and burial leads to 86.5% mortality reduction, which is a synergistic over additive effect that far exceeds the efficacy of the single interventions. The complementary effect are shown in the dynamics of Figure1919 as treatment drives the disease to fatal outcomes while safe burial prevents transmission after death.
Combining interventions produced the same results. the combined three type intervention ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0$) only reached 60.4% reduction, while the rest ($u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$) and the comprehensive method ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)) both reached the maximum reduction of 86.5% gained with the treatment and safe burial pair. It is the combination of these results, documented in Figures 2225 and  $28$ that confirms treatment and safe burial dominate the loss-of-life mitigation, while in these circumstances, personal protection and vaccination are less effective. In the context of EVD, personal protection does not mean social distancing; rather, it refers to the use of barrier techniques for those who come into direct interaction with infected persons or dead bodies. The combination of safe burial and treatment is synergistic and remains the optimal strategy for mortality reduction. It is the persistent inability to curtail infection prevalence that most strongly implies the mortality reduction is the consequence of the disease, postulated to contain altered outcome probabilities. This underscores the importance of post-exposure treatment combined with the clinical burial of the dead to reduce casualties.
8 Intelligent Disease Modeling – DINN of EVD
This segment presents the mathematical concepts supporting the architecture of the Disease-Informed Neural Network (DINN) as per the formulae of the fractional-order Ebola model that was discussed. (2). In the case of the DINN architecture, the fractional differential equations constrain the architectural configuration of the neural-embedded network, enabling the network to incorporate the differential equations of the epidemic system defined by the fractional-order differential equations, thus capturing the system’s physically plausible dynamics.
The governing system of fractional-order differential equations is expressed in the compact form:
$^{C}D_{t}^{\alpha}Y(t)+\mathcal{G}(Y(t;\chi))=0,\quad t\in[t_{0},T],$ (36)
The state Y(t) = [S(t), V(t), E(t), $I_{s}(t)$, $I_{a}(t)$, H(t), D(t), R(t)] along with the vector concept that every compartment population tends to zero. Let $\mathcal{G}(Y)=[g_{1}(Y),\ldots,g_{8}(Y)]$, where each function $g_{i}(Y)$ captures the heterogeneously nonlinear connections that inflow and outflow through its compartment. $\chi\in\mathbb{R}^{p}$ the parameters in the system which determines epidemiological corresponding with the degree/level of the interventions effectiveness and the intervention measures being undertaken. The neural network is approximated as centralized with the equations $NN_{w,b}(t):\mathbb{R}\longrightarrow\mathbb{R}^{8}$ which is initial feasible $Y(t)$ solution because of the abstract layer termed trainable beings the connected parameters the neural network in the $w$ and $b$ configurations as edges as nodes as the two fundamental elements in the neural network. The subsequent component of the function is termed the data portion is given by
$\mathcal{L}_{data}=\frac{1}{m}\sum_{s=1}^{m}\|NN_{w,b}(t_{s})-Y_{s}\|^{2}.$ (37)
The physical consistency is enforced through the residual term, which measures the discrepancy between the neural network output and the governing fractional dynamics:
$\mathcal{R}(NN_{w,b},t;\chi)={}^{C}D_{t}^{\alpha}NN_{w,b}+\mathcal{G}(NN_{w,b}(t)).$ (38)
The corresponding residual loss is computed as:
$MSE_{w,b,\chi}^{\mathcal{G}}=\frac{1}{N}\sum_{i=1}^{N}\|\mathcal{R}(NN_{w,b},t_{i};\chi)\|^{2},$ (39)
The use of collocation points $t_{1},t_{2},\ldots,t_{N}$ discretizes the temporal domain. The entirety of this optimization problem simultaneously evaluates the parameters defining the neural network and the system by means of the composite objective:
$(w^{*},b^{*},\chi^{*})=\arg\min_{w,b,\chi}\left(\mathcal{L}_{data}+\lambda MSE_{w,b,\chi}^{\mathcal{G}}\right),$ (40)
where $\lambda$ balances the contribution of data compliance and physical consistency. This formulation guarantees that the acquired solution serves the purpose of not only fitting the data but also complying with the core mechanistic precepts embedded in the fractional-order Ebola transmission model.
8.1 ML-EVD model
This subsection describes the concrete realization of the Disease-Informed Neural Network (DINN) framework design for the fractional-order Ebola epidemic model. The advanced physical consistency is maintained through an elaborate residual term that contains the entire system dynamics:
$\mathcal{R}(NN_{w,b},t;\chi)=\begin{pmatrix}{}^{C}D_{t}^{\alpha}S-\left[\Lambda-\lambda S+\omega V-(v+\mu)S\right]\\ {}^{C}D_{t}^{\alpha}V-\left[vS-(1-\varepsilon)\lambda V-(\mu+\omega)V\right]\\ {}^{C}D_{t}^{\alpha}E-\left[\lambda(S+(1-\varepsilon)V)-(\mu+\sigma)E\right]\\ {}^{C}D_{t}^{\alpha}I_{s}-\left[p\sigma E-(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right]\\ {}^{C}D_{t}^{\alpha}I_{a}-\left[\sigma E(1-p)-(\gamma_{a}+\mu)I_{a}\right]\\ {}^{C}D_{t}^{\alpha}H-\left[h_{s}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H\right]\\ {}^{C}D_{t}^{\alpha}D-\left[\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D\right]\\ {}^{C}D_{t}^{\alpha}R-\left[\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R\right]\end{pmatrix}$, (41)
where the time-dependent force of infection is defined as:
$\lambda(t)=\frac{c\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N},$ (42)
Let $N$ denotes the total living population such that $N=S+V+E+I_{s}+I_{a}+H+R$. The residual loss function can be split into compartment-specific components, each measuring the gap between the neural network’s fractional derivatives and the corresponding dynamical equations:
$MSE_{residual}$ $=\sum_{X\in\mathcal{C}}MSE_{X},$ (43)
$MSE_{S}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}S-\Lambda+\lambda S-\omega V+(v+\mu)S\right|^{2},$ (44)
$MSE_{V}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}V-vS+(1-\varepsilon)\lambda V+(\mu+\omega)V\right|^{2},$ (45)
$MSE_{E}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}E-\lambda(S+(1-\varepsilon)V)+(\mu+\sigma)E\right|^{2},$ (46)
$MSE_{I_{s}}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}I_{s}-p\sigma E+(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right|^{2},$ (47)
$MSE_{I_{a}}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}I_{a}-\sigma E(1-p)+(\gamma_{a}+\mu)I_{a}\right|^{2},$ (48)
$MSE_{H}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}H-h_{s}I_{s}+(\gamma_{h}+\delta_{h}+\mu)H\right|^{2},$ (49)
$MSE_{D}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}D-\delta_{s}I_{s}-\delta_{h}H+\mu_{d}D\right|^{2},$ (50)
$MSE_{R}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}R-\gamma_{s}I_{s}-\gamma_{a}I_{a}-\gamma_{h}H+\mu R\right|^{2},$ (51)
where $\mathcal{C}=\{S,V,E,I_{s},I_{a},H,D,R\}$ indicates all the epidemiological compartments. Consistency with the observational measurements is guaranteed with the components set within the data adherence phase
$MSE_{data}$ $=\sum_{X\in\mathcal{C}}MSE_{X}^{data},$ (52)
$MSE_{X}^{data}$ $=\frac{1}{s}\sum_{i=1}^{s}|X(t_{i})-X_{i}^{o}|^{2},\quad X\in\mathcal{C},$ (53)
Let $X_{i}^{o}$ be the observed value of compartment $X$ at time $t_{i}$. The entire optimization task finds the values of both the neural network parameters and the parameters of the epidemiological model simultaneously:
$(w^{*},b^{*},\chi^{*})=\arg\min_{w,b,\chi}\left(MSE_{data}+MSE_{residual}\right),$ (54)
The collection of parameters $\chi=\{\Lambda,\mu,c(t),\beta,\eta_{a},\eta_{d},v(t),\varepsilon,\omega,\sigma,p,h_{s},\gamma_{s},\gamma_{a},\gamma_{h},\delta_{s},\delta_{h},\mu_{d},\alpha\}$ includes all constants related to the flow of epidemiological modeling to describe the transmission, progression of the disease and the impact of different control measures. Using this approach ensures that the learned solution integrates the observed data with the mechanistic underpinning of the Ebola transmission model.
8.2 Network Architecture and Training
The design of the neural networks processes sequential data through several hidden layers, with hyperbolic tangent activations, leading to outputs associated with the epidemiological model’s eight state variables. Composite loss functions leverage both Xavier initialization and the Adam optimization method to facilitate training. Fractional derivatives are computed through automatic differentiation, and the enforcement of physical constraints within the entire temporal domain remains intact.
8.3 Parameter Estimation and Model Calibration
While building the model, the following epidemiological parameters need to be defined: $\Lambda$, $\mu$, $\beta$, $\eta_{a}$, $\eta_{d}$, $\varepsilon$, $\omega$, $\sigma$, $p$, $h_{s}$, $\gamma_{s}$, $\gamma_{a}$, $\gamma_{h}$, $\delta_{s}$, $\delta_{h}$, $\mu_{d}$, $\alpha$, the fractional order, the time-dependent functions $c(t)$ and $v(t)$, and the network parameters $w$ and $b$. Parameter estimation for the model will be refined as data for the model will become available. As outbreak data becomes available, the configuration and calibration will be refined to minimize the discrepancy between the data and the bounds defined by outbreak data. Parameter definition will subsequently be based on the mechanistic model and data from the identified model components, which will be integrated dynamically. Parameter definition will subsequently be based on the mechanistic model and data from the identified model components, which will be integrated dynamically.
8.4 Results and Discussion of DINN
The results from the implementation of the Disease-Informed Neural Network (DINN) context were carefully reviewed in relation to the benchmark datasets. The results were consolidated across all 8 epidemiological compartments. Integrated modeling accuracy estimation as performed records from Figure 29 to 36 showcase the framework to outstanding. The model collects above average predictive accuracy and extensive generalization, proving the framework’s relevancy, particularly on the Ebola virus disease predictive modeling’s sophisticated system dynamics. The DINN was constructed on PyTorch which facilitated the architecture and gradient calculations through automated differentiation. The model weights were trained on the composite loss function which included the background of fractional-order differential equations, and the dynamic was supported by the standard Computational Python Libraries (NumPy, SciPy) for numeric and data operations. The DINN’s predictive performance level in this case was bench marked in contrast to the $R^{2}$ (coefficient of determination) values on a designated test set, a saved, unbiased representation of the attached unseen data in the model. According to computational results the DINN’s test set $R^{2}$ values instead drifted around the theoretical maxima of the $R^{2}$ ranging from 0.9912 in the Vaccinated compartment to about 0.9999 in the Exposed, Symptomatic, and Asymptomatic compartments. Correspondingly significant, the model captured the largest populated and fundamental transmission dynamics. Hold-out population, the Susceptible compartment, with a test $R^{2}$ of 0.9975. The $R^{2}$ values in the splits of Training, Validation, and Test and all states in the architecture, as well as other deep neural networks, are an illustration of the system dynamics learned by the model without overfitting the variables, which is a common neural network model overfitting weak point. The inconsistencies between predictions and ground truth data are dictated by NMSEs and further validated by reliably low NMSEs across all compartments at approximately $10^{-4}$ to $10^{-5}$. The analysis Sensitivity (Vaccinated $(R^{2}=0.9912)$ dynamic) compartments highly errors successively. Sensibly, this is not random due to the high complexity involving human behaviors and the variance of fully stochastic closed compartment models. In contrast, the model’s capability to estimate the Exposed $(R^{2}=0.9999)$ and Symptomatic $(R^{2}=0.9998)$ compartments is appreciated. These Correlatives are crucially underlined as the target of health for the forecast regarding the active clinical burden. In addition, the estimates for the Deceased $(R^{2}=0.9998)$ and Recovered $(R^{2}=0.9996)$ individuals positively highlight and estimate the ultimate volume of the outbreak with mortality. In addition to the state variable forecasting, the DINN also drew key parameters of epidemiology from the data. Estimating error for most parameters was frighteningly low - over 9% for 13 out of 17 parameters was an astounding standard. The most notable and successful achievement was properly estimating the fractional order $(\alpha)$ and estimating low error critical transmission parameters like the contact rate $(\beta,8.12\%$ error) and the comparative transmissibility of the deceased $(\eta_{d},4.33\%$ error). The leading estimation error came from the natural mortality rate $(\mu)$ which, from case lines alone is greatly surprising and, be seated as the most undervalued parameter. In the short, low impact, and highly volatile case outflow of an outbreak, it strongly diverges from the standard frame. The comprehensive recovery outline parameter recommends that the DINN is not simply a “black-box” predictor which provides illogical reasoning. It is a process that completely understands disease mechanisms. Finally, we can say that the DINN framework set the standard for defining and predicting scenarios for future EVD outbreaks. It is both highly accurate and robust for modeling and analyzing the disease. The framework’s outstanding ability to recover every system parameter and predict perfectly through every compartment. It also helps as a breakthrough for hybrid physics-informed machine learning in epidemiology. It guarantees that future real-time forecasts and scenarios will be provided.
Figure 29: Susceptible compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00071; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 30: Vaccinated compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00009; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 31: Exposed compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00010; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 32: Symptomatic compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00019; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 33: Asymptomatic compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00015; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 34: Hospitalized compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00026; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 35: Deceased compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00034; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
Figure 36: Recovered compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00085; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
9 Conclusion
This work has been done to enhance the mathematical framework of forecasting and possibility of intervention regarding the outbreaks of Ebola Virus Disease (EVD). The identified primary objectives focused on the creation of an advanced fractional order model to capture the EVD spread dynamics through adaptive multilayered channels which is considered to be extremely complex, formulation of reliable epidemiological thresholds, the defining of appropriate intervention strategies, and the creation of a novel computational model on the estimation of parameters along with the model standardization.
Integration of the three disciplines of mathematics: fractional calculus, optimal control theory, and scientific machine learning, was the original part of this research. EVD transmission dynamics required memory incorporated non-Markovian multi-scale dynamics. Thus, an eight-compartment fractional-order model was built around the Caputo derivative. The model well-posedness was aided by the foundational fixed-point theory, the subsequent next-generation matrix approach to compartment analysis enabled the reproduction number $\mathcal{R}_{0}$ calculation, and the complete fractal Lyapunov characterization for the system imbued a formative influence to the system. Utilizing Pontryagin’s Maximum Principle, the optimal control problem which was built sequentially around four time-adjustable control interventions dedicated to fractional control systems was solved. The main breakthrough was the design of the Disease-Informed Neural Network (DINN) framework, which dynamically computes the mechanistic parameters by embedding the governing fractional differential equations as physical constraints in a neural network and solving for parameter estimation.
The analysis brings to light three key findings: the epidemiological threshold $\mathcal{R}_{0}$, unique bifurcation properties at the critical points tied to $\mathcal{R}_{0}$, and the findings on global stability which show a forward bifurcation at $\mathcal{R}_{0}=1$. One of the most important parameters indicating potential outbreak from the tiered hierarchy parameters from a sensitivity analysis are the transmission rate ($\beta$), the incubation rate ($\sigma$), and deceased infectiousness ($\eta_{d}$). Results from optimal control show that combining treatment with safe burial, which highlights the synergistic nature of the interventions, results in the greatest impact of an 86.5% decrease in mortality relative to the impact of each intervention applied in isolation. All epidemiological compartments also show outstanding predictive accuracy with the DINN context, as seen in the coefficients of determination $R^{2}$ which lie between 0.9987 and 0.9999, coupled with recovering system parameters with less than 18% accuracy for most primary transmission and progression rates. Although the scope of the model are comprehensive, some simplifications are made to the assumptions of the epidemiology regarding spatial diversity, age-structured contact matrices, and stochastic transmission. While parameter estimation is robust, and within the DINN framework, it is still a function of the investigation data quality and handling. It is definitely worth revealing that optimal control theory, while developing practical strategies, does have the control theory limits as the minor, and in practice resource partial, bounds.
The gaps mentioned in the research offer many opportunities for prospective work. Adding spatial dynamics and meta-population structures would improve the model’s ability to channel cross regional transmission and tailored modified intervention impacts. Further model improvement would arise from the inclusion of stochastic elements that measure the risk of destruction and the chances of restoration of an outbreak. Other innovative applications of the DINN framework would be in the area of predictive public health for automatic adaptive real-time intervention planning to control ongoing outbreaks. This would be an immense advance in public health. Also, implementing this integrated mathematical framework with machine learning to other emerging infectious diseases would extremely transform the epidemic awareness and response framework.
This study underscores how the integration of advanced techniques in mathematics and contemporary computing techniques provides valuable tools for public health. It illustrates the transition of epidemiology from purely descriptive to prescriptive, allowing for active intervention in the control of outbreaks.

### v3

@@
-Intelligent-Based Neural Networks and Optimal Control of Fractional Order Ebola Virus Dynamics
+
 Abstract
-Ebola virus disease is a severe, often deadly form of hemorrhagic fever, and is rapidly transmissible.
-It’s spread through human contact is increased through the direct exchange of infected blood, body fluids, and organs, as well as contact with surfaces on which these infected fluids have been dropped.
-Ebola’s complicated dynamics spread beyond integer order models, representing a central intersection for the application of machine learning to control theory.
-In the present study, we construct an innovative prime framework of fractional order calculus which integrates data-driven machine learning with optimal control in order to model and analyze the Ebola virus disease.
-We develop an eight-compartment model based on the Caputo fractional derivative to incorporate the non-Markovian memory effects inherent in the disease’s progression and regression.
-The system explicitly structures the distinct transmission pathways from symptomatic and asymptomatic infections, as well as the critical transmission dynamics associated with deceased individuals.
-We analytically prove the global well-posedness of the model, compute the basic reproduction number $\mathcal{R}_{0}$ to characterize the epidemic threshold, and prove global stability theorems.
-Sensitivity analysis indicates that the basic reproduction number is most critically sensitive to transmission rate, incubation period, and the relative infectivity of the deceased.
-The treatment-safe burial synergy optimization control provides a simultaneous 86.5% of the target morbidity and mortality rate control of the population.
-The safe burial host intervention is the most effective control.
-Our study categorizes the best control strategies and orders them in terms of effectiveness.
-The proposed disease-informed neural network achieves practically perfect predictive accuracy and describe the test set $R^{2}$ values of 0.991 - 0.999, and correspondingly 99.1% - 99.9% prediction accuracy as well as align with the real-world epidemic behavior.
-The first successful combination of process-based modeling with artificial intelligence for dealing an outbreak of an infectious disease has been established in this work.
-This work presents an innovative mechanistic modeling framework for infectious diseases, which has the potential to advance epidemic response analysis.
-Keywords: Fractional Calculus, Ebola Modeling, Epidemic Threshold ($\mathcal{R}_{0}$), Sensitivity Analysis, Optimal Control, Disease-Informed Neural Networks.
+Ebola virus disease is a severe, often fatal infectious disease transmitted through direct contact with infected bodily fluids and contaminated surfaces.
+This study addresses the serious need for integrated predictive and control frameworks by developing a unified approach that connects intelligent outbreak prediction systems with mathematical models of Ebola transmission.
+We propose a novel nonlinear fractional-order eight-compartment SVEIsIaHDR model that integrates control measures, personal protection, vaccination, treatment, and safe burial to analyze their effectiveness in reducing Ebola transmission and improving outbreak prediction.
+Mathematical validation establishes the model’s boundedness, non-negativity, and well-posedness.
+We derive the basic reproduction number R0 via the next-generation matrix, identify disease-free and endemic equilibrium, and conduct stability analyses.
+Sensitivity analysis reveals R0’s primary sensitivity to transmission rate $\beta$ (+1.00), incubation period $\sigma$ (+0.6250), and infectivity of the dead $\eta_{d}$ (+0.4678).
+An optimal control framework minimizes disease management costs (4 dollars per infection).
+Numerical simulation by the RK45 method demonstrates combined treatment and safe burial controls 86.5% of target morbidity and mortality, with safe burial being the most effective single measure.
+We develop a disease-informed neural network (DINN) achieving 99.87% prediction accuracy ($R^{2}=0.9987$) with minimal error (NMSE $10^{-4}$ to $10^{-5}$).
+The DINN accurately recovers 13 of 17 transmission parameters with $9\%$ error, establishing a new interpretability benchmark.
+This integrated framework provides public health officials with a practical tool to test interventions and predict outbreaks through model-data integration.
+Keywords: Ebola virus disease, fractional-order compartmental model, basic reproduction number ($\mathcal{R}_{0}$), optimal control, sensitivity analysis, stability analysis, outbreak prediction, disease-informed neural network (DINN).
 1 Introduction
-Ebola Virus Disease (EVD) continues to be one of the most difficult problems for the world health organization to tackle, with estimated case fatality rates varying from concerning 25% to 90% across various outbreaks [1].
-The Ebola Virus Disease (EVD) is classified as a tropical zoonosis which involves cross transmission from animal reservoirs to human populations [2, 3].
-The earliest recorded case of the disease occurred in Zaire (now known as Democratic Republic of Congo) in 1976, named after the Ebola River where the first outbreak took place [4].
-With data around confirmed cases and deaths standing at 28,646 and 11,323, respectively, and spanning through three countries; Liberia, Guinea, and Sierra Leone, the ‘West African Ebola Virus Disease epidemic’ (2014-16) is considered the biggest ‘Ebola’ outbreak in history.
-The surprising number of cases and deaths during the outbreak and the new ways of fighting the complexity both demonstrated the failure of prediction and outbreak control traditional prediction and control methods used [5].
-The outbreak still emphasized the complexity controlled in the actual mathematics and logic of the base prediction and control frameworks prediction, controlling, and modeling control frameworks outlined around the multi and complex dimensions in the spread of the EVD.
-Mathematical techniques in epidemiology and their application span new prospects and new dimensions in the field [6, 7].
-While these techniques were developed in pure science, their fundamental calculations and distributions resource divisions and practical works approximations [8].
-The epidemiology of the western African case outbreak hence has triggered the much-needed efficiency flow in prediction case numbers and evaluating the efficiency of preventive measures and their real-time value [9].
-The greatest puzzling challenge suffering in the environmental and acoustic scales spread while modalities of EVD transmissions is the combination of factors that various phoneme extents growing paradigm [10].
-The epidemiology of Ebola Virus Disease (EVD) has many distinguishing features which imply that estimating the disease outbreaks remaining traditional techniques will not be sufficient.
-For instance, the disease’s incubation period is highly variable and spans anywhere from 2 to 21 days, with the 8-10 days interval more commonly renowned [11].
-There is a further complication within the disease’s clinical presentation and development as initial fever, weakness, and diarrhea unaccompanied by bleeding and other clinical symptoms do not signal an active infection and contagiousness.
-Additionally, contact with infected animals and the consumption of bush-meat are the first chance for the disease’s human transmission from its natural reservoirs of fruit bats [12, 13].
-Subsequently, the disease transmission post- infection is unlinked and random.
-Furthermore, symptomatic ($I_{s}$), asymptomatic ($I_{a}$), and post-mortem ($D$) transmission routes for the human population.
-[14].
-The high level of infectiousness of dead body, coupled with the perseverance of customarily traditional burial practices in the region, requires mathematically sophisticated treatment beyond typical integer-order differential equations, as outlined in earlier works [15].
-Evidence shows that washing the deceased as part of ethnically prescribed performs is a relevant vector for transmission.
-Besides, the prospect of reinfection with EVD remains very real, as survivors own only quickly protective, low-potency antibodies, with research showing the complete absence of Ebola antibodies in some survivor’s years post-recovery [16, 17].
-Such an immunological situation is certainly an additional complication from the modeling perspective whereby the absence of an unified framework for archiving immune waning over a long period rightly remains a gap.
-The natural, spatial spread of EVD transmission improves so far, another level of complexity to the analytical modelling of the disease dynamics.
-The dynamics of EVD differ significantly across different geographic scales.
-Urbanized areas tend to experience outbreaks at a faster rate than rural areas because of the high population density, varied patterns of movement, and better medical infrastructure [18].
-Moreover, the phenomenon of superspreader events, where a small proportion of the infected population is responsible for a suspiciously high number of secondary infections, adds a considerable amount of variability and complexity that challenges resolution by purely deterministic approaches [19, 20].
-These characteristics need modeling methods that combine large-scale spatial dynamics with fine-scale spatio-temporal transmission patterns [21].
-Although traditional modeling techniques such as the SI, SIR, and SEIR frameworks form the basis of understanding, they all fundamentally relate to the Markovian assumptions which do not take into account memory and long-range chronological dependencies which are characteristic of EVD spread [22, 23].
-These limitations are particularly clear when attempting to model the distributed incubation periods, the decline of vaccine-induced immunity, and the heterogeneous contact patterns during the real-world epidemic spread [11, 24].
-The SEIR model used by [25] and [26], and the SIHDP model employed by [27] and in [28] SEIHFDBR are notable developments, however, they are still limited in the integer-order nature of their models.
-Fractional calculus has long shown the potential to model memory effects and other non-local connections, and thus provides a framework well-intentioned of study to overcome some of the noted deficiencies: [29].
-Unlike integer-order derivatives, which always adopt Markovian processes, fractional operators capture the distributed suspensions which are characteristic of epidemiological systems, such as extended incubation period, waning immunity, and environmental reservoirs of pathogens [30].
-Evidence of progress in the modeling of infectious diseases for which data are available and for which fractional calculus applies exhibits greater concord with experimental indication, especially for diseases with long-range chronological correlations and power-law decline in the efficiency of involvements [31, 32].
-The non-Markovian nature of some fractional derivatives offers a better-suited mathematical framework for modeling the transmission of disease in and among structured populations [33].
-In addressing challenges, some recent studies have included various extensions.
-For example, [34] included pathways for environmental contamination in their SIRPD model with respect to the Ebola pathogens in the environment.
-In [35] worked on the decomposition of the deceased compartment by distinguishing dead but not buried, and dead who are buried.
-Control optimization has been studied by, for instance, [36] on optimal combinations of vaccination and treatment, and [37] on educational campaigns in a two-patch model.
-Media awareness campaigns have been studied by [38] and [39] and have shown the importance of information broadcasting in controlling a disease.
-The link of Machine Learning algorithms with Process-based model and models of machine learning with traditional epidemiological forecasting is a new concept as noted by [40].
-Predictable models afford a theoretical understanding of epidemiology through biological lenses and as such are weak at capturing sophisticated structures non-linear in nature that are defined with basic compartmental approaches [41].
-Recent activities in hybrid frameworks combining the Physics-based model interpretability and the neural networks’ model complexity to capture epidemiological dynamics have demonstrated improved predictive performance [42].
-But, as advanced computational techniques apply to EVD, there are gaps such as real time parameter estimation from EVD noise where EVD observation data is incomplete [43].
-The development of neural networks with physics-informed neural network (PINN) and epidemiology constraints could bridge this gap [44].
-However, modern modeling system contexts still have three main shortcomings.
-One, they have a limited representation of the memory effects of fractional-order derivatives, which is crucial for modeling the distributed intervals in the progression of a disease and the corresponding delay of the immune response.
-Two, there is still a lack of fully integrating optimal control theory combined with advanced multi-compartment models that summarize the entirety of the EVD transmission dynamics.
-Three, here is still insufficient use of machine learning for the estimation of parameters for complicated epidemiological models which reduces the flexibility and accuracy of the models in current practice.
-The application of optimal control theory is underdeveloped yet central to the design of effective intervention and control strategies for the economic epidemiology systems.
-Identifying optimal control strategies is a challenging mathematical problem.
-It requires temporally-resolved control strategies and sophisticated numerical methods for high dimensional, nonlinear systems with memory [45].
-Also, minimizing the case count, death, and economic loss for epidemic reduction response requires Pareto-optimal strategies that balance the contradictory objectives [46].
-The recent, rapid advancements in adaptable computational optimal control technologies—particularly those dealing with the direct collocation of the control function and the system trajectory and the Hamiltonian-based methods are methods of great attention in computational biologically structured control [47].
-The development of fractional calculus has presented new ways to tackle these problems [29].
-The Caputo fractional derivative approaches non-local chronological dependencies and memory effects with its non-Markovian structure, which mathematically captures epidemiological data [48].
-Preliminary work connecting fractional-order models focused on EVD, such as [49] with the Atangana-Baleanu-Caputo operators and [50] with the nonlinear time-fractional mathematical model.
-Despite these advances, there is still much to be anticipated in the overlap among these fields, mainly control theory, machine learning, and the other works on the theme.
-This research discourses these important gaps by revolutionary a new computational framework that combines fractional-order dynamics, multi-pathway transmission architectures, and control optimization, which is enhanced by machine learning.
-Building upon the preceding work by [34] which did not include exposed compartments, hospitalizations, and safe burial practices, I also consider reinfection through the immunological studies by [16] and [17].
-This research directs numerous contributions and novelties:
+A serious infectious disease, Ebola Virus Disease (EVD) is transmitted to people by reservoirs of animals such as fruit bats.
+When the Ebola virus transmits from animal reservoirs into human populations, an outbreak occurs [1, 2].
+The illness has historically significant case fatality rates, with outbreaks ranging from 25% to 90% [3].
+It was first reported in the Democratic Republic of the Congo in 1976; the name Ebola was given due to the Ebola River [4].
+With 28,646 confirmed infections and 11,323 deaths recorded, the most catastrophic epidemic was in West Africa (2014–2016), with almost a 50% death ratio [5].
+Fever, weakness, diarrhea, and, in extreme situations, hemorrhagic signs are common in infected people.
+There is a major issue in that these viruses’ incubation periods differ, sometimes 2 days and sometimes 21 days; infected individuals cannot spread the virus throughout the incubation phase [6].
+Fruit bats are considered to be the main natural source of infection in humans, and non-human primates serve as end hosts, maintaining cycles of transmission that can result in major outbreaks [7, 8].
+Controlling measures are complicated, particularly frequent contact paths that boost the spread of EVD.
+Being in contact with infectious body fluids, such as blood or excretions, results in subsequent human-to-human transmission.
+Contact with deceased individuals during traditional burial practices is a crucial and widely recognized route of transmission that has been a significant facilitator of prior epidemics [9, 10].
+In addition, the virus might continue to spread indirectly by remaining on contaminated surfaces.
+Due to the absence of apparent symptoms, various demographic categories, in particular people who participate in burial customs and asymptomatic individuals, have a possibility to spread infection throughout communities.
+Therefore, it is difficult for standard mathematical models to capture the EVD dynamics of transmission, as they have to immediately identify these various and context-specific threats [11, 12].
+Ebola transmission is rendered much more difficult by the potential for reinfection.
+After recovery, immunity may not be total or long-lasting because protective antibodies may gradually decline in survivors [13, 14].
+Individuals who have previously been infected can test positive for the virus years later, indicating potential re-exposure and infection, according to data from previous outbreaks [15, 16].
+Additionally, research on populations that have been exposed frequently has discovered people who display no symptoms but do not have detectable antibodies, underscoring the heterogeneity in immune response [17, 18].
+Now, we focus on a few equivalent mathematical models of EVD that have already been proposed by other researchers.
+Fractional calculus has been applied to infectious disease modeling in a number of research investigations, displaying more consistency with observed outbreak data than traditional approaches [19, 20].
+Fractional calculus has long shown the potential to model memory effects and other non-local connections and thus provides a well-intentioned framework of study to overcome some of the noted deficiencies [21].
+Unlike integer-order derivatives, which always adopt Markovian processes, fractional operators capture the distributed suspensions that are characteristic of epidemiological systems, such as extended incubation periods, waning immunity, and environmental reservoirs of pathogens [22].
+Evidence of progress in the modeling of infectious diseases for which data are available and for which fractional calculus applies exhibits greater concord with experimental indication, especially for diseases with long-range chronological correlations and power law decline in the efficiency of involvements [23, 24].
+The non-Markovian nature of some fractional derivatives offers a better-suited mathematical framework for modeling the transmission of disease in and among structured populations [25].The SI model, such as [26, 27], the SEIR model [28, 29], the SEIRD model [30], and the SEIRHD model [31, 32, 33, 34], where $S$, $E$, II, $H$, $R$, and $D$ represent susceptible, exposed, infectious, hospitalized, recovered, and dead/deceased compartments, respectively [35], separated the Dead/Deceased ($D$) compartment into two compartments, buried deceased ($R_{D}$) and unburied dead ($F$).
+Similarly, in their SIRPD model, where $P$ represents environmental contamination, [26] included EVD transmission through contaminated surroundings of deceased patients.
+While models like the SEIHFDBR in [36] represent notable developments, they remain limited by their integer-order formulation.
+Understanding how to effectively control the spread of infection and facilitate decision-making during an epidemic is one of the applications of mathematical modeling.
+For example, [37] presented an SIR model-based optimum control issue with treatment and immunization as possible control strategies.
+They looked into the most effective vaccination and therapy combinations to lower the cost of those preventive measures.
+In [38] investigated various diseases transmitted by insects that directly infect the host population, such as dengue fever, malaria, and the West Nile virus, using optimal control analysis.
+In order to examine the transmission of Ebola, [39] designed a model.
+They evaluated the effects of various preventative measures such as vaccination, public awareness campaigns, early identification campaigns, improved hospital sanitation, separation of people with infections, and geographical constraints.
+Another field of research is the integration of machine learning with traditional epidemiological models.
+When handling noisy or incomplete surveillance information, physics-informed neural networks (PINNs) demonstrate particular potential for parameter estimation and prediction [40, 41, 42].
+A comprehensive analysis of the last five years’ worth of literature, however, reveals enduring shortcomings and disparate study directions.
+Although fractional models have been created, they are frequently not integrated with intricate multi-pathway transmission designs specific to EVD, for instance.
+Likewise, machine learning applications often function independently of mathematical models, which limits their understanding and generalization, and optimal control formulations usually use simplified models that do not fully capture the complexity of transmission dynamics [43].
+Nevertheless, despite these developments, there are still important research gaps in present literature.
+Firstly, there is a gap between the full representation of all recognized EVD transmission channels (symptomatic, asymptomatic, and burial customs) and fractional-order models, which account for memory effects.
+There is a methodological gap given that detailed transmission models generally employ integer-order approaches, whereas the majority of fractional models simplify transmission dynamics.
+Second, the application of optimal control theory to high-dimensional fractional systems that reflect EVD dynamics has been insufficient.
+One major challenge that has limited practical applicability is the mathematical complexity of determining optimal control policies for such systems, especially when including numerous intervention strategies with different costs and effectiveness.
+[44, 45].
+Third, the majority of applications currently only combine biological process-based models with data-driven machine learning techniques.
+Hybrid frameworks have been proposed, but they typically lack the seamless combination needed for real-time parameter estimation and prediction during current outbreaks [46].
+This study develops a novel computational framework with the particular objectives that follow in order to address these connected challenges.
+The practical objective is to provide public health decision-makers an efficient computational tool that can forecast disease paths with measured uncertainty, optimize allocation of resources for maximum impact, and simulate outbreak scenarios under different intervention strategies.
+The following is a list of this study’s primary innovations and contributions:
 […]
-  • Comprehensive mathematical analysis signifying the well-posedness of the system through Banach’s fixed-point theorem, the basic reproduction number of the model $\mathcal{R}_{0}$ exogenously resulting through the next generation matrix, and the proof of global stability of the system through fractional Lyapunov methods [48].
-  • Derivation and providing an analytical solution for a multi-objective optimal control problem with four intervention strategies (personal-protection, vaccination, treatment, and safe burial) via Pontryagin’s Maximum Principle.
-  • Establishing a Disease-Informed Neural Network (DINN) framework for simultaneous solution discovery and parameter estimation which inserts physical constraints for better predictive performances.
-  • Detailed cost-effectiveness analysis for optimal intervention portfolios for public health choices.
-This model potentially assistances public health authorities in responding to outbreaks more effectively by evidence-based intervention, which can be game-changing in undertaking and enhancing the efficiency of outbreak responses.
-Integrating mathematical modeling with computational intelligence, this study attempts to productive infectious disease interventions through addressing the organized response and gaps in the current outbreak response system.
+  • We established the optimal control model that would help to reduce the cost of managing the disease by considering four intervention measures that included personal protection, vaccination, treatment and safe burial.
+Our theoretical analysis by applying Pontryagin’s Maximum Principle and numerical simulation using the RK45 method permitted identifying the cost-effective strategies, and the level of their effectiveness in terms of transmission and mortality reduction.
+  • Motivated from physics-informed neural networks, we designed a new framework for disease-informed neural networks (DINN).
+This computational approach achieves high prediction accuracy across all disease stages and demonstrates strong parameter recovery capability, providing an interpretable tool for outbreak forecasting and intervention assessment.
+This work significantly improves both academic research and healthcare practice by successfully addressing the existing gap between theoretical modeling and actual application.
+It offers investigators a new methodological framework that shows how machine learning, optimal control theory, and fractional calculus may be successfully used for the scientific study of infectious diseases.
 This paper proceeds as follows.
-In Section 2, I discuss the elementary ideas of fractional calculus which is key to developing the model.
+In Section 2, I discuss the elementary ideas of fractional calculus, which is key to developing the model.
 In Section 3, I develop an eight-compartment, fractional-order Ebola transmission model.
-Section 4 contains the in-depth mathematical theory regarding the model’s well-posedness, equilibrium, and stability over the system.
-Section 5 conducts global sensitivity analysis of the dominant parameters through Latin Hypercube Sampling and Partial Rank Correlation Coefficients.
-In Section 6, I formulate and solve the optimal control problem based on Pontryagin’s Maximum Principle.
-Section 7 incorporates the neural network model concerning disease and the results obtained in the corresponding numerical analysis.
-In the closing Section 8, I analyze the outcome of the model along with its drawbacks and suggest some directions for further research.
+Section 4 presents the complete mathematical analysis of the model, including proofs of existence and uniqueness of solutions, identification of the invariant region and attractivity, derivation of the basic reproduction number ($\mathcal{R}_{0}$), and both local and global stability analyses for the disease-free and endemic equilibrium points.
+Section 5 performs a global sensitivity analysis that attempts to quantity and prioritize the impact of key model parameters on the disease transmission dynamics.
+Section 6 establishes and solves optimal control problem through the Maximum Principle of Pontryagin founding the most effective intervention strategies.
+In section 7, the formation of a disease-informed neural network (DINN) model along with the relevant numerical data is developed and fully analyzed.
+Lastly, in Section 8 the main findings are synthesized, the methodological limitations are critically discussed, and future research directions outlined.
 2 Preliminaries
 This part sets out the basic principles of fractional calculus which is essential to the construction and analysis of our model.
-Definition 2.1 (Caputo Fractional Derivative)
+Definition 2.1
 Let $\alpha\in(0,1]$ and $u\in C^{1}([0,T])$.
-For any function $u$ which is continuously differentiable in the closure $[0,T]$ the Caputo fractional derivative of order $\alpha$ is defined in [29] as follows:
+For any function $u$ which is continuously differentiable in the closure $[0,T]$ the Caputo fractional derivative of order $\alpha$ is defined in [21] as follows.
 $^{C}D_{t}^{\alpha}u(t)=\frac{1}{\Gamma(1-\alpha)}\int_{0}^{t}(t-\tau)^{-\alpha}u^{\prime}(\tau)d\tau.$
 This operator is preferred for modeling dynamical systems as it accommodates standard initial conditions, and $^{C}D_{t}^{\alpha}u(t)\to u^{\prime}(t)$ as $\alpha\to 1^{-}$, recovering the classical integer-order derivative.
-Definition 2.2 (Mittag-Leffler Function)
-The two-parameter Mittag-Leffler function is defined by the series [29]:
-$E_{\alpha,\beta}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{\Gamma(\alpha k+\beta)},\quad\alpha,\beta>0,\,z\in\mathbb{C}.$
-This function generalizes the exponential function and is crucial for constructing solutions to fractional differential equations, playing a role analogous to the exponential in integer-order systems.
-Definition 2.3 (Fractional Lyapunov Stability)
+Definition 2.2
 The equilibrium point $y^{*}$ of the fractional-order system $^{C}D_{t}^{\alpha}y(t)=f(t,y(t))$ is said to be Mittag-Leffler stable if
-$\|y(t)-y^{*}\|\leq\left[m(y(0)-y^{*})E_{\alpha,1}(-\lambda t^{\alpha})\right]^{b},$
+$\|y(t)-y^{*}\|\leq\left[m(y(0)-y^{*})E_{\alpha,1}(-\lambda t^{\alpha})\right]^{b}.$
 where $\alpha\in(0,1)$, $\lambda\geq 0$, $b>0$, $m(0)=0$, and $m(y)\geq 0$. Mittag-Leffler stability implies asymptotic stability [47].
-Definition 2.4 (Physics-Informed Neural Network Loss)
-For a system governed by $^{C}D_{t}^{\alpha}\mathbf{Y}(t)-\mathbf{F}(t,\mathbf{Y}(t))=0$, the loss function for a Physics-Informed Neural Network (PINN) $NN(t;\theta)$ approximating $\mathbf{Y}(t)$ is defined as:
-$\mathcal{L}(\theta)=\mathcal{L}_{data}+\mathcal{L}_{physics},$
-where $\mathcal{L}_{data}=\frac{1}{N_{d}}\sum_{i=1}^{N_{d}}|NN(t_{i};\theta)-\mathbf{Y}_{i}|^{2}$ enforces data fidelity, and $\mathcal{L}_{physics}=\frac{1}{N_{c}}\sum_{j=1}^{N_{c}}|{}^{C}D_{t}^{\alpha}NN(t_{j};\theta)-\mathbf{F}(t_{j},NN(t_{j};\theta))|^{2}$ enforces the physical law described by the fractional differential equation.
-Theorem 2.1 (Stability Condition for Fractional Systems)
+Lemma 2.1
 Let $\alpha\in(0,1]$.
-An equilibrium point $y^{*}$ of the autonomous system $^{C}D_{t}^{\alpha}y(t)=f(y(t))$ is asymptotically stable if all eigenvalues $\lambda_{i}$ of the Jacobian matrix $J=\partial f/\partial y|_{y^{*}}$ satisfy the condition [48]:
+An equilibrium point $y^{*}$ of the autonomous system $^{C}D_{t}^{\alpha}y(t)=f(y(t))$ is asymptotically stable if all eigenvalues $\lambda_{i}$ of the Jacobian matrix $J=\partial f/\partial y|_{y^{*}}$ satisfy the condition [47].
 $|\arg(\lambda_{i})|>\frac{\alpha\pi}{2},\quad\forall i.$
 This condition generalizes the classical Routh-Hurwitz criterion to fractional-order systems.
-3 Fractional-Order Ebola Transmission Dynamics
-We formulate a Caputo fractional-order model ($\alpha\in(0,1]$) which incorporates the multi-pathway transmission, memory effects, and time-varying interventions of Ebola.
+3 Model formulation
+We formulate a Caputo fractional-order model ($\alpha\in(0,1]$) that incorporates the multi-pathway transmission, memory effects, and time-varying interventions of Ebola.
 The governing system is given as follows.
@@
 ${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R.$
-The force of infection is given by:
+The force of infection is given by
 $\lambda=\frac{c\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N},$
-and for notational simplicity, we define:
+and for notational simplicity, we define the model as
 $q_{0}$ $=(v+\mu),$ $q_{1}$ $=(\mu+\omega),$ $q_{2}$ $=(\mu+\sigma),$
@@
 $q_{6}$ $=\mu_{d}.$
-The model parameters include: natural mortality ($\mu$) and recruitment ($\Lambda$); vaccination rate $v(t)$ with waning immunity ($\omega$); and force of infection $\lambda(t)$ driving exposures.
+The model parameters include natural mortality ($\mu$) and recruitment ($\Lambda$), vaccination rate $v(t)$ with waning immunity ($\omega$), and force of infection $\lambda(t)$ driving exposures.
 Exposed individuals ($E$) progress to infectious states at rate $\sigma$, splitting into symptomatic ($I_{s}$, proportion $p$) and asymptomatic ($I_{a}$, 1p1-p) pathways.
 Symptomatic cases face hospitalization ($h_{s}$), disease-induced mortality ($\delta_{s}$), or recovery ($\gamma_{s}$).
-Asymptomatic and hospitalized individuals recover at rates $\gamma_{a}$ and $\gamma_{h}$ respectively, with hospital mortality ($\delta_{h}$).
+Asymptomatic and hospitalized individuals recover at rates $\gamma_{a}$ and $\gamma_{h}$, respectively, with hospital mortality ($\delta_{h}$).
 Deceased individuals ($D$) transmit infection until safe burial ($\mu_{d}$).
-Transmission coefficients $\beta$, $\eta_{a}$, and $\eta_{d}$ govern infectivity from symptomatic, asymptomatic, and post-mortem sources.
-The system is governed by:
+Transmission coefficients $\beta$, $\eta_{a}$, and $\eta_{d}$ govern infectivity from symptomatic, asymptomatic, and post-mortem sources.The system is governed by,
 ${}^{C}D_{t}^{\alpha}S(t)$ $=\Lambda-\lambda S+\omega V-q_{0}S,$ (2)
@@
 In this Caputo fractional formulation, the system in (2) captures the dynamics of Ebola transmission in an eight-dimensional frame.
-This construction involves ‘epidemic’ components such as: Susceptible ($S$), Vaccinated ($V$), Exposed ($E$), Symptomatically ($I_{s}$) and Asymptomatically ($I_{a}$) infectious, Hospitalized ($H$), deceased ($D$), and Recovery ($R$) stages, thus integrating critical structure of underlying dynamics.
-Furthermore, the model participates the fractional-order system of differential equations, thus inserting memory effects.
+This construction involves ‘epidemic’ components such as susceptible ($S$), vaccinated ($V$), exposed ($E$), symptomatically ($I_{s}$) and asymptomatically ($I_{a}$) infectious, hospitalized ($H$), deceased ($D$), and recovery ($R$) stages, thus integrating the critical structure of underlying dynamics.
+Furthermore, the model participates in the fractional-order system of differential equations, thus inserting memory effects.
+Table 1: Parameters for the fractional-order Ebola model.
+Symbol Description Value Range Source
+$\beta$ Transmission rate 0.287 0.2–0.4 [48]
+$\eta_{d}$ Deceased infectiousness 0.734 0.5–0.9 [48]
+$\eta_{a}$ Asymptomatic infectiousness 0.523 0.3–0.7 Assumed (Fitted)
+$\sigma$ Incubation rate (day-1) 0.094 0.05–0.15 [49]
+$p$ Symptomatic proportion 0.712 0.5–0.8 [49]
+$\gamma_{s}$ Symptomatic recovery (day-1) 0.068 0.05–0.1 [26]
+$\gamma_{a}$ Asymptomatic recovery (day-1) 0.089 0.07–0.12 [26]
+$\delta_{s}$ Symptomatic death (day-1) 0.103 0.05–0.15 [50]
+$\delta_{h}$ Hospitalized death (day-1) 0.067 0.03–0.09 [50]
+$h_{s}$ Hospitalization rate 0.312 0.2–0.4 [51]
+$\Lambda$ Recruitment rate 100 100–1000 [52]
+$\mu$ Natural mortality (day-1) $3.5\times 10^{-5}$ [52]
+$v(t)$ Vaccination rate 0.005–0.08 Assumed (Control)
+$\varepsilon$ Vaccine efficacy 0.85–0.95 Assumed (Control)
+$\omega$ Waning immunity (day-1) 0.0027–0.0037 Assumed
+$\alpha$ Fractional order 0.85 0.75–0.95 Assumed (Model)
 […]
 Figure 1: Transmission architecture of the fractional Ebola model.
-Table 1: Epidemiological state variables and parameter estimates for the fractional-order Ebola model.
-Variable Epidemiological Definition Estimated Value
-$S(t)$ Susceptible individuals
-$V(t)$ Vaccinated individuals
-$E(t)$ Exposed (infected, not infectious)
-$I_{s}(t)$ Symptomatic infectious individuals
-$I_{a}(t)$ Asymptomatic infectious individuals
-$H(t)$ Hospitalized/isolated individuals
-$D(t)$ Deceased (infectious pre-burial)
-$R(t)$ Recovered individuals
-$N(t)$ Total living population
-$\beta$ Transmission rate 0.287
-$\eta_{a}$ Relative infectiousness (asymptomatic) 0.523
-$\eta_{d}$ Relative infectiousness (deceased) 0.734
-$\sigma$ Incubation rate (day-1) 0.094
-$p$ Proportion symptomatic 0.712
-$\gamma_{s}$ Recovery rate (symptomatic) 0.068
-$\delta_{s}$ Death rate (symptomatic) 0.103
-$h_{s}$ Hospitalization rate 0.312
-$\gamma_{a}$ Recovery rate (asymptomatic) 0.089
-$\delta_{h}$ Death rate (hospitalized) 0.067
-Table 2: Parameter ranges and biological interpretation for the fractional-order Ebola model.
-Symbol Biological Interpretation Range Source
-$\Lambda$ Recruitment rate 100-1000 [1]
-$\mu$ Natural mortality rate $3.5\times 10^{-5}$ [51]
-$c(t)$ Contact rate function 0.4-1.2 [9]
-$v(t)$ Vaccination rate 0.005-0.08 [52]
-$\varepsilon$ Vaccine efficacy 0.85-0.95 [52]
-$\omega$ Waning immunity rate 1/365-1/270 [53]
-$\mu_{d}$ Burial rate 0.15-0.35 [54]
-$\alpha$ Fractional order 0.75-0.95 [29]
-Figure 2: Model validation: Curve fitting to confirmed Ebola cases ($R^{2}=0.96$, RMSE = 23.45).
 4 Mathematical Analysis
@@
 We establish the mathematical well-posedness of the fractional-order Ebola system through existence and uniqueness analysis.
-Theorem 4.1 (Existence and Uniqueness)
-Given any $\alpha\in(0,1]$, and initial conditions
-$\mathbf{X}_{0}=(S_{0},V_{0},E_{0},I_{s0},I_{a0},H_{0},D_{0},R_{0})\in\mathbb{R}_{+}^{8}$, the fractional-order system (2) has the unique solution $\mathbf{X}(t)=(S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t))$ for all $t\geq 0$.
-Proof.  Define the state vector $\mathbf{X}(t)=(S,V,E,I_{s},I_{a},H,D,R)^{T}$ and express the system as:
+Theorem 4.1
+Given any $\alpha\in(0,1]$, and initial conditions ae $\mathbf{X}_{0}=(S_{0},V_{0},E_{0},I_{s0},I_{a0},H_{0},D_{0},R_{0})\in\mathbb{R}_{+}^{8}$, the fractional-order model (2) has the unique solution $\mathbf{X}(t)=(S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t))$ for all $t\geq 0$.
+Proof.  Define the state vector $\mathbf{X}(t)=(S,V,E,I_{s},I_{a},H,D,R)^{T}$ and express the system as
 $^{C}D_{t}^{\alpha}\mathbf{X}(t)=\mathbf{F}(t,\mathbf{X}(t)),\quad\mathbf{X}(0)=\mathbf{X}_{0}.$
-The equivalent Volterra integral formulation is:
-$\mathbf{X}(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds.$ (3)
-Assume the operator $\mathcal{T}:C([0,T],\mathbb{R}^{8})\to C([0,T],\mathbb{R}^{8})$ defined by:
-$(\mathcal{T}\mathbf{X})(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds.$
+The equivalent Volterra integral formulation is
+$\mathbf{X}(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds,$ (3)
+Assume the operator $\mathcal{T}:C([0,T],\mathbb{R}^{8})\to C([0,T],\mathbb{R}^{8})$ defined by
+$(\mathcal{T}\mathbf{X})(t)=\mathbf{X}_{0}+\frac{1}{\Gamma(\alpha)}\int_{0}^{t}(t-s)^{\alpha-1}\mathbf{F}(s,\mathbf{X}(s))ds,$
 For $\mathbf{X},\mathbf{Y}\in C([0,T],\mathbb{R}^{8})$ with $\|\mathbf{X}\|=\sup_{t\in[0,T]}\|\mathbf{X}(t)\|_{\infty}$, the Lipschitz continuity of $\mathbf{F}$ follows from bounded parameters and finite state variables over compact intervals.
-Specifically, there exists $L>0$ such that:
-$\|\mathbf{F}(t,\mathbf{X})-\mathbf{F}(t,\mathbf{Y})\|_{\infty}\leq L\|\mathbf{X}-\mathbf{Y}\|_{\infty}.$
-The contraction estimate yields:
-$\|\mathcal{T}\mathbf{X}-\mathcal{T}\mathbf{Y}\|\leq\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}\|\mathbf{X}-\mathbf{Y}\|.$
-Selecting $\tau>0$ such that $\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}<1$ ensures $\mathcal{T}$ is a contraction on $C([0,\tau],\mathbb{R}^{8})$.
-Banach’s fixed point theorem guarantees a unique local solution $\mathbf{X}^{*}\in C([0,\tau],\mathbb{R}^{8})$.
-Global extension to $0,\infty)$ follows from the linear growth condition $\|\mathbf{F}(t,\mathbf{X})\|_{\infty}\leq M(1+\|\mathbf{X}\|_{\infty})$ for some $M>0$, which precludes finite-time blow-up.
-Iterative continuation establishes global existence and uniqueness.
-$\square$
+Specifically, there exists $L>0$ such that
+$\|\mathbf{F}(t,\mathbf{X})-\mathbf{F}(t,\mathbf{Y})\|_{\infty}\leq L\|\mathbf{X}-\mathbf{Y}\|_{\infty},$ (4)
+The contraction estimate provides that
+$\|\mathcal{T}\mathbf{X}-\mathcal{T}\mathbf{Y}\|\leq\frac{L\tau^{\alpha}}{\Gamma(\alpha+1)}\|\mathbf{X}-\mathbf{Y}\|.$ (5)
+[…]
 4.2 Invariant Region and Attractivity
-The fractional-order Ebola system (2) evolves within the biologically feasible region:
+The fractional-order Ebola model (2) evolves within the biologically feasible region is
 $\Omega=\left\{(S,V,E,I_{s},I_{a},H,D,R)\in\mathbb{R}^{8}_{\geq 0}:\quad D\leq\frac{(\delta_{s}+\delta_{h})}{\mu_{d}}\left(\frac{\Lambda}{\mu}+\epsilon\right),N_{L}\leq(\frac{\Lambda}{\mu}+\epsilon)\right\},$ (6)
 for $\epsilon>0$, where $N_{L}=S+V+E+I_{s}+I_{a}+H+R$ denotes the living population.
-Theorem 4.2 (Invariance and Attractivity)
-The region $\Omega$ is:
+Theorem 4.2
+The region $\Omega$ is defined as
   1. (i)
-    Positively invariant under (2),
+    Positively invariant under model (2),
   2. (ii)
     Globally attractive in $\mathbb{R}^{8}_{\geq 0}$.
-Proof.  (i) Positivity and Invariance: Applying the fractional positivity lemma [29], we verify boundary behavior:
-${}^{C}D_{t}^{\alpha}S|_{S=0}=\Lambda+\omega V\geq 0,\quad{}^{C}D_{t}^{\alpha}V|_{V=0}=vS\geq 0,$
+Proof.  Applying the fractional positivity lemma from [21], we verify boundary behavior as,
+${}^{C}D_{t}^{\alpha}S|_{S=0}=\Lambda+\omega V\geq 0,\quad{}^{C}D_{t}^{\alpha}V|_{V=0}=vS\geq 0,$ (7)
 ${}^{C}D_{t}^{\alpha}E|_{E=0}=\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}\geq 0,$
@@
 ${}^{C}D_{t}^{\alpha}R|_{R=0}=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H\geq 0.$
-Non-negativity of all boundary fluxes ensures $\mathbb{R}_{+}^{8}$ is invariant. (ii) Boundedness and Attractivity: For the living population $N_{L}$:
-$^{C}D_{t}^{\alpha}N_{L}=\Lambda-\mu N_{L}-(\delta_{s}I_{s}+\delta_{h}H)\leq\Lambda-\mu N_{L}.$
-Laplace transformation yields:
-$N_{L}(s)\leq\frac{\Lambda}{s(s^{\alpha}+\mu)}+N_{L}(0)\frac{s^{\alpha-1}}{s^{\alpha}+\mu},$
-with inverse transform:
+Non-negativity of all boundary fluxes ensures $\mathbb{R}_{+}^{8}$ is invariant.
+The total living population, defined as $N_{L}(t)=S(t)+V(t)+E(t)+I_{s}(t)+I_{a}(t)+H(t)+R(t)$, evolves according to
+$^{C}D_{t}^{\alpha}N_{L}=\Lambda-\mu N_{L}-(\delta_{s}I_{s}+\delta_{h}H)\leq\Lambda-\mu N_{L},$ (8)
+Applying the Laplace transform to the inequality $^{C}D_{t}^{\alpha}N_{L}\leq\Lambda-\mu N_{L}$ gives
+$N_{L}(s)\leq\frac{\Lambda}{s(s^{\alpha}+\mu)}+N_{L}(0)\frac{s^{\alpha-1}}{s^{\alpha}+\mu},$ (9)
+Applying the inverse Laplace transform yields
 $N_{L}(t)\leq N_{L}(0)E_{\alpha,1}(-\mu t^{\alpha})+\Lambda t^{\alpha}E_{\alpha,\alpha+1}(-\mu t^{\alpha}).$
-As $t\to\infty$, $E_{\alpha,1}(-\mu t^{\alpha})\to 0$ and $t^{\alpha}E_{\alpha,\alpha+1}(-\mu t^{\alpha})\to 1/\mu$, giving:
+Taking the limit as $t\to\infty$, we use the asymptotic behaviors $E_{\alpha,1}(-\mu t^{\alpha})\to 0$ and $t^{\alpha}E_{\alpha,\alpha+1}(-\mu t^{\alpha})\to 1/\mu$,
+which gives the bound
 $\limsup_{t\to\infty}N_{L}(t)\leq\frac{\Lambda}{\mu}.$
-For deceased individuals:
+For deceased individuals, we derive from the differential inequality
 $^{C}D_{t}^{\alpha}D\leq(\delta_{s}+\delta_{h})\left(\frac{\Lambda}{\mu}+\epsilon\right)-\mu_{d}D,$
-implying:
+and consequently obtain
 $\limsup_{t\to\infty}D(t)\leq\frac{\delta_{s}+\delta_{h}}{\mu_{d}}\left(\frac{\Lambda}{\mu}+\epsilon\right).$
 Thus, all trajectories are eventually absorbed into $\Omega$. $\square$
-4.3 EVD-Free Equilibrium
-The state wherein a population exists without any infections of the Ebola virus is referred to the Ebola-free equilibrium (DFE), at which point all infected compartments have completely dissipated:
+4.3 Ebola-Free Equilibrium
+The state in which the population is completely free of Ebola virus infection is known as the disease-free equilibrium (DFE). At this equilibrium, all infected compartments vanish, meaning
 $E=I_{s}=I_{a}=H=D=0.$
-The system (2) reduces to:
-$0$ $=\Lambda-q_{0}S+\omega V,$ (5)
-$0$ $=vS-q_{1}V,$ (6)
-$0$ $=-\mu R.$ (7)
-From (7), $R^{*}=0$. Equations (5) and (6) form the linear system:
-$q_{0}S-\omega V$ $=\Lambda,$ (8)
-$-vS+q_{1}V$ $=0.$ (9)
+The system (2) then reduces to
+$0$ $=\Lambda-q_{0}S+\omega V,$
+$0$ $=vS-q_{1}V,$
+$0$ $=-\mu R.$
+From the third equation, we obtain $R^{*}=0$. The first two equations form the linear system
+$q_{0}S-\omega V$ $=\Lambda,$
+$-vS+q_{1}V$ $=0.$
 The coefficient matrix
@@
 $\det(A)=q_{0}q_{1}-\omega v=\mu(q_{1}+v)>0,$
-ensuring a unique solution. Applying Cramer’s rule:
+which guarantees a unique solution. Applying Cramer’s rule yields
 $S^{*}=\frac{\Lambda q_{1}}{\mu(q_{1}+v)},\quad V^{*}=\frac{\Lambda v}{\mu(q_{1}+v)}.$
-Thus, the disease-free equilibrium is:
-$\mathcal{E}_{0}=\left(\frac{\Lambda q_{1}}{\mu(q_{1}+v)},\frac{\Lambda v}{\mu(q_{1}+v)},0,0,0,0,0,0\right).$
-This biologically consistent DFE provides the foundation for stability analysis and reproduction number derivation.
-[…]
-$\mathcal{R}_{0}$ is computed via the next generation matrix method as expressed in [55].
-In the case of the infected subsystem $X=(E,I_{s},I_{a},H,D)^{T}$, the new infection ($F$) and transition ($V$) matrices are as follows:
-$F=\begin{pmatrix}0&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\end{pmatrix},\quad V=\begin{pmatrix}q_{2}&0&0&0&0\\ -p\sigma&q_{3}&0&0&0\\ -(1-p)\sigma&0&q_{4}&0&0\\ 0&-h_{s}&0&q_{5}&0\\ 0&-\delta_{s}&0&-\delta_{h}&q_{6}\end{pmatrix}$
-where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$ with $N^{*}=\frac{\Lambda}{\mu}$.
-The inverse transition matrix is:
-$V^{-1}=\begin{pmatrix}\frac{1}{q_{2}}&0&0&0&0\\ \frac{p\sigma}{q_{2}q_{3}}&\frac{1}{q_{3}}&0&0&0\\ \frac{(1-p)\sigma}{q_{2}q_{4}}&0&\frac{1}{q_{4}}&0&0\\ \frac{p\sigma h_{s}}{q_{2}q_{3}q_{5}}&\frac{h_{s}}{q_{3}q_{5}}&0&\frac{1}{q_{5}}&0\\ \frac{p\sigma(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{2}q_{3}q_{5}q_{6}}&\frac{\delta_{s}q_{5}+\delta_{h}h_{s}}{q_{3}q_{5}q_{6}}&0&\frac{\delta_{h}}{q_{5}q_{6}}&\frac{1}{q_{6}}\end{pmatrix}$
-The spectral radius of the next generation matrix is $K=FV^{-1}$ yields:
-$\mathcal{R}_{0}=\beta\Psi\left[\frac{p\sigma}{q_{2}}\left(\frac{1}{q_{3}}+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)\sigma}{q_{2}q_{4}}\right]$
-Substituting $\Psi$ provides the explicit form:
-$\mathcal{R}_{0}=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\left[\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}\right]$
-This threshold quantity governs disease invasion and extinction dynamics.
+Thus, the disease-free equilibrium is given by
+$\mathcal{E}_{0}=\left(\frac{\Lambda q_{1}}{\mu(q_{1}+v)},\frac{\Lambda v}{\mu(q_{1}+v)},0,0,0,0,0,0\right).$ (10)
+This biologically consistent DFE serves as the basis for stability analysis and the derivation of the basic reproduction number.
+4.4 Basic Reproduction Number $\mathcal{R}_{0}$
+The basic reproduction number $\mathcal{R}_{0}$ is computed using the next-generation matrix method [53].
+Considering the infected subsystem $X=(E,I_{s},I_{a},H,D)^{T}$, we define the new infection matrix $F$ and the transition matrix $V$ as
+$F=\begin{pmatrix}0&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\end{pmatrix},\qquad V=\begin{pmatrix}q_{2}&0&0&0&0\\ -p\sigma&q_{3}&0&0&0\\ -(1-p)\sigma&0&q_{4}&0&0\\ 0&-h_{s}&0&q_{5}&0\\ 0&-\delta_{s}&0&-\delta_{h}&q_{6}\end{pmatrix}$,
+where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$ and $N^{*}=\frac{\Lambda}{\mu}$. The inverse of the transition matrix is
+$V^{-1}=\begin{pmatrix}\frac{1}{q_{2}}&0&0&0&0\\ \frac{p\sigma}{q_{2}q_{3}}&\frac{1}{q_{3}}&0&0&0\\ \frac{(1-p)\sigma}{q_{2}q_{4}}&0&\frac{1}{q_{4}}&0&0\\ \frac{p\sigma h_{s}}{q_{2}q_{3}q_{5}}&\frac{h_{s}}{q_{3}q_{5}}&0&\frac{1}{q_{5}}&0\\ \frac{p\sigma(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{2}q_{3}q_{5}q_{6}}&\frac{\delta_{s}q_{5}+\delta_{h}h_{s}}{q_{3}q_{5}q_{6}}&0&\frac{\delta_{h}}{q_{5}q_{6}}&\frac{1}{q_{6}}\end{pmatrix}$.
+The next-generation matrix is $K=FV^{-1}$, and its spectral radius gives the basic reproduction number
+$\mathcal{R}_{0}=\beta\Psi\left[\frac{p\sigma}{q_{2}}\left(\frac{1}{q_{3}}+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)\sigma}{q_{2}q_{4}}\right].$ (11)
+Substituting the expression for $\Psi$ yields the explicit form
+$\mathcal{R}_{0}=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\left[\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}\right].$ (12)
+This threshold quantity governs the disease invasion and extinction dynamics.
 (a) Cumulative confirmed cases progression
@@
 Figure 2: Ebola Virus Disease outbreak analysis during the 2014-2016 West Africa outbreak showing: (a) cumulative confirmed cases, (b) cumulative deaths, (c) active cases, and (d) comparative analysis of all outbreak metrics.
-4.5 Critical Vaccination Coverage
-The critical vaccination coverage defines the minimum population immunity required for disease elimination, occurring when $\mathcal{R}_{0}=1$. Beginning with the reproduction number:
-$\mathcal{R}_{0}(v)=\mathcal{R}_{0}(0)\cdot\frac{q_{1}+(1-\varepsilon)v}{q_{1}+v}$
-where $\mathcal{R}_{0}(0)=K$ represents the vaccination-free reproduction number. Setting $\mathcal{R}_{0}(v_{c})=1$ yields:
-$1=\mathcal{R}_{0}(0)\cdot\frac{q_{1}+(1-\varepsilon)v_{c}}{q_{1}+v_{c}}$
-Solving for the critical vaccination rate $v_{c}$:
-$q_{1}+v_{c}$ $=\mathcal{R}_{0}(0)[q_{1}+(1-\varepsilon)v_{c}]$
-$v_{c}-\mathcal{R}_{0}(0)(1-\varepsilon)v_{c}$ $=\mathcal{R}_{0}(0)q_{1}-q_{1}$
-$v_{c}[1-\mathcal{R}_{0}(0)(1-\varepsilon)]$ $=q_{1}[\mathcal{R}_{0}(0)-1]$
-$v_{c}=\frac{q_{1}[\mathcal{R}_{0}(0)-1]}{1-(1-\varepsilon)\mathcal{R}_{0}(0)}$
-The critical vaccination coverage $p_{c}$, representing the proportion vaccinated at equilibrium, is:
-$p_{c}=\frac{v_{c}}{q_{1}+v_{c}}=\frac{\mathcal{R}_{0}(0)-1}{\varepsilon\mathcal{R}_{0}(0)}$
-Thus, the herd immunity threshold for Ebola elimination is:
-$p_{c}=\max\left(0,1-\frac{1}{\varepsilon\mathcal{R}_{0}(0)}\right)$
-This establishes the minimal vaccination coverage required to interrupt disease transmission.
-Figure 4: 3D surface: Vaccination and waning immunity impact on $\mathcal{R}_{0}$.
-The surface shows how the basic reproduction number varies with vaccination rate ($v$) and waning immunity rate ($\omega$).
-4.6 Local Stability Analysis of the EVD-Free Equilibrium
+4.5 Local Stability Analysis of the Disease-Free Equilibrium
 Theorem 4.3
-When $\mathcal{R}_{0}<1$, the EVD-free equilibrium $\mathcal{E}_{0}$ of the system (2) is achieved and remains locally stable; however, it becomes unsteady and $\mathcal{R}_{0}>1$ will prove it to be unstable.
-Proof.  The Jacobian matrix evaluated at $\mathcal{E}_{0}$ takes the block triangular form:
-$J(\mathcal{E}_{0})=\begin{pmatrix}J_{11}&J_{12}\\ 0&J_{22}\end{pmatrix}$
-where $J_{11}$ corresponds to the uninfected compartments $(S,V,R)$ and $J_{22}$ to the infected compartments $(E,I_{s},I_{a},H,D)$. The uninfected subsystem matrix is:
-$J_{11}=\begin{pmatrix}-(v+\mu)&\omega&0\\ v&-(\mu+\omega)&0\\ 0&0&-\mu\end{pmatrix}$
-with eigenvalues $\lambda_{1}=-\mu$, $\lambda_{2}=-\mu$, $\lambda_{3}=-(\mu+v+\omega)$.
-All eigenvalues satisfy $|\arg(\lambda_{i})|=\pi>\frac{\alpha\pi}{2}$ for $\alpha\in(0,1]$, confirming stability of the uninfected subsystem.
-The infected subsystem Jacobian is:
-$J_{22}=\begin{pmatrix}-q_{2}&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ p\sigma&-q_{3}&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0\\ 0&h_{s}&0&-q_{5}&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}\end{pmatrix}$
+The disease-free equilibrium $\mathcal{E}_{0}$ of the model (2) is locally asymptotically stable when $\mathcal{R}_{0}<1$, and unstable when $\mathcal{R}_{0}>1$.
+Proof.  The Jacobian matrix evaluated at $\mathcal{E}_{0}$ exhibits a block triangular structure
+$J(\mathcal{E}_{0})=\begin{pmatrix}J_{11}&J_{12}\\ 0&J_{22}\end{pmatrix}$,
+where $J_{11}$ corresponds to the uninfected compartments $(S,V,R)$ and $J_{22}$ corresponds to the infected compartments $(E,I_{s},I_{a},H,D)$.
+For the uninfected subsystem, the matrix is
+$J_{11}=\begin{pmatrix}-(v+\mu)&\omega&0\\ v&-(\mu+\omega)&0\\ 0&0&-\mu\end{pmatrix}$.
+The eigenvalues of $J_{11}$ are $\lambda_{1}=-\mu$, $\lambda_{2}=-\mu$, and $\lambda_{3}=-(\mu+v+\omega)$.
+For $\alpha\in(0,1]$, all eigenvalues satisfy $|\arg(\lambda_{i})|=\pi>\alpha\pi/2$, ensuring the stability of the uninfected subsystem.
+For the infected subsystem, the Jacobian matrix takes the form
+$J_{22}=\begin{pmatrix}-q_{2}&\beta\Psi&\beta\eta_{a}\Psi&0&\beta\eta_{d}\Psi\\ p\sigma&-q_{3}&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0\\ 0&h_{s}&0&-q_{5}&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}\end{pmatrix}$,
 where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$.
-We decompose $J_{22}=F-V$, where $F$ and $V$ are the next-generation matrices previously defined.
-The stability of $J_{22}$ is determined by the spectral radius $\rho(FV^{-1})=\mathcal{R}_{0}$.
-According to the fractional Routh-Hurwitz criteria and properties of Metzler matrices, when $\mathcal{R}_{0}<1$, all eigenvalues of $J_{22}$ have negative real parts and satisfy the fractional stability condition $|\arg(\lambda)|>\frac{\alpha\pi}{2}$.
-Conversely, when $\mathcal{R}_{0}>1$, the spectral abscissa of $J_{22}$ is positive, indicating at least one eigenvalue with positive real part.
-By the fractional stability theorem, this establishes instability of the disease-free equilibrium.
+We observe that $J_{22}$ can be decomposed as $J_{22}=F-V$, where $F$ and $V$ are the next-generation matrices defined previously.
+The stability of $J_{22}$ is governed by the spectral radius $\rho(FV^{-1})=\mathcal{R}_{0}$.
+According to the fractional Routh-Hurwitz criteria and the properties of Metzler matrices, when $\mathcal{R}_{0}<1$, all eigenvalues of $J_{22}$ have negative real parts and satisfy the fractional stability condition $|\arg(\lambda)|>\alpha\pi/2$.
+Conversely, when $\mathcal{R}_{0}>1$, the spectral abscissa of $J_{22}$ is positive, indicating at least one eigenvalue with a positive real part.
+By the fractional stability theorem, this establishes the instability of the disease-free equilibrium.
 $\square$
-4.7 Global Stability of EVD-Free Equilibrium
+4.6 Global Stability of the Disease-Free Equilibrium
 Theorem 4.4
-When $\mathcal{R}_{0}<1$, the EVD-free equilibrium $\mathcal{E}_{0}$ of the system (2) is globally asymptotically stable in $\Omega$.
-Proof.  Consider the function for the infected compartments already introduced in the above section.
-$L(E,I_{s},I_{a},H,D)=E+AI_{s}+BI_{a}+CH+FD,$
-where the parameters A, B, C, and F are positive and not arbitrary.
-To investigate the function, we follow the trajectories described in the system and compute the Caputo derivative along each trajectory:
-${}^{C}D_{t}^{\alpha}L=$ $\left[[\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}]-q_{2}E\right]$
-$+A\left(p\sigma E-q_{3}I_{s}\right)+B((1-p)\sigma E-q_{4}I_{a})$
-$+C\left(h_{s}I_{s}-q_{5}H\right)+F(\delta_{s}I_{s}+\delta_{h}H-q_{6}D)$
-Rearranging terms by compartment:
+If $\mathcal{R}_{0}<1$, the disease-free equilibrium $\mathcal{E}_{0}$ of the model (2) is globally asymptotically stable in the invariant region $\Omega$.
+Proof.  We construct a Lyapunov function for the infected compartments based on the approach outlined in the previous section
+$L(E,I_{s},I_{a},H,D)=E+AI_{s}+BI_{a}+CH+FD,$ (13)
+where $A$, $B$, $C$, and $F$ are positive constants to be determined.
+Following the system trajectories, we compute the Caputo derivative along solutions
+${}^{C}D_{t}^{\alpha}L$ $=\left[\frac{c\beta[S+(1-\varepsilon)V](I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}-q_{2}E\right]$
+$\quad+A\left(p\sigma E-q_{3}I_{s}\right)+B\left((1-p)\sigma E-q_{4}I_{a}\right)$
+$\quad+C\left(h_{s}I_{s}-q_{5}H\right)+F\left(\delta_{s}I_{s}+\delta_{h}H-q_{6}D\right).$
+Rearranging terms by compartment yields
 ${}^{C}D_{t}^{\alpha}L$ $=\left[-q_{2}+Ap\sigma+B(1-p)\sigma\right]E$
-$+\left[\frac{c\beta[S+(1-\varepsilon)V]}{N}-Aq_{3}+Ch_{s}+F\delta_{s}\right]I_{s}$
-$+\left[\frac{c\beta\eta_{a}[S+(1-\varepsilon)V]}{N}-Bq_{4}\right]I_{a}$
-$+\left[-Cq_{5}+F\delta_{h}\right]H$
-$+\left[\frac{c\beta\eta_{d}[S+(1-\varepsilon)V]}{N}-Fq_{6}\right]D$
-We select coefficients to eliminate positive terms:
-$C$ $=\frac{F\delta_{h}}{q_{5}},\quad B=\frac{\beta\eta_{a}\Psi}{q_{4}},\quad F=\frac{\beta\eta_{d}\Psi}{q_{6}}$
-$A$ $=\frac{\beta\Psi}{q_{3}}+\frac{Ch_{s}}{q_{3}}+\frac{F\delta_{s}}{q_{3}}=\frac{\beta\Psi}{q_{3}}+\frac{F\delta_{h}h_{s}}{q_{3}q_{5}}+\frac{F\delta_{s}}{q_{3}}$
-Using the bound $\frac{S+(1-\varepsilon)V}{N}\leq\Psi$ valid throughout $\Omega$, we obtain:
+$\quad+\left[\frac{c\beta[S+(1-\varepsilon)V]}{N}-Aq_{3}+Ch_{s}+F\delta_{s}\right]I_{s}$
+$\quad+\left[\frac{c\beta\eta_{a}[S+(1-\varepsilon)V]}{N}-Bq_{4}\right]I_{a}$
+$\quad+\left[-Cq_{5}+F\delta_{h}\right]H$
+$\quad+\left[\frac{c\beta\eta_{d}[S+(1-\varepsilon)V]}{N}-Fq_{6}\right]D.$
+We now select the coefficients to eliminate positive terms from the $I_{s}$, $I_{a}$, $H$, and $D$ components
+$C$ $=\frac{F\delta_{h}}{q_{5}},\quad B=\frac{\beta\eta_{a}\Psi}{q_{4}},\quad F=\frac{\beta\eta_{d}\Psi}{q_{6}},$
+$A$ $=\frac{\beta\Psi}{q_{3}}+\frac{Ch_{s}}{q_{3}}+\frac{F\delta_{s}}{q_{3}}=\frac{\beta\Psi}{q_{3}}+\frac{F\delta_{h}h_{s}}{q_{3}q_{5}}+\frac{F\delta_{s}}{q_{3}},$ (14)
+where $\Psi=\frac{S^{*}+(1-\varepsilon)V^{*}}{N^{*}}$.
+Using the inequality $\frac{S+(1-\varepsilon)V}{N}\leq\Psi$, which holds throughout $\Omega$, we obtain the estimate
 ${}^{C}D_{t}^{\alpha}L$ $\leq\left[-q_{2}+\sigma\left(\frac{p\beta\Psi}{q_{3}}+\frac{pF\delta_{h}h_{s}}{q_{3}q_{5}}+\frac{pF\delta_{s}}{q_{3}}+\frac{(1-p)\beta\eta_{a}\Psi}{q_{4}}\right)\right]E$
-$+\beta\Psi\left[1-1\right]I_{s}+\beta\eta_{a}\Psi\left[1-1\right]I_{a}+\beta\eta_{d}\Psi\left[1-1\right]D$
-Substituting $F=\frac{\beta\eta_{d}\Psi}{q_{6}}$ yields:
-$^{C}D_{t}^{\alpha}L\leq q_{2}\left[-1+\mathcal{R}_{0}\right]E$
-For $\mathcal{R}_{0}\leq 1$, we have $^{C}D_{t}^{\alpha}L\leq 0$ throughout $\Omega$, with equality if and only if $E=0$. When $E=0$, the system dynamics ensure:
-${}^{C}D_{t}^{\alpha}I_{s}$ $=-q_{3}I_{s}\Rightarrow I_{s}\to 0$
-${}^{C}D_{t}^{\alpha}I_{a}$ $=-q_{4}I_{a}\Rightarrow I_{a}\to 0$
-${}^{C}D_{t}^{\alpha}H$ $=h_{s}I_{s}-q_{5}H\Rightarrow H\to 0$
-${}^{C}D_{t}^{\alpha}D$ $=\delta_{s}I_{s}+\delta_{h}H-q_{6}D\Rightarrow D\to 0$
-Thus, the maximal invariant set where $^{C}D_{t}^{\alpha}L=0$ is precisely $\mathcal{E}_{0}$.
-Application of the fractional LaSalle invariance principle establishes global asymptotic stability when $\mathcal{R}_{0}\leq 1$.
-For $\mathcal{R}_{0}>1$, instability follows from local analysis, while uniform persistence can be demonstrated via established methods in fractional dynamical systems [48].
-$\square$
+$\quad+\beta\Psi\left[1-1\right]I_{s}+\beta\eta_{a}\Psi\left[1-1\right]I_{a}+\beta\eta_{d}\Psi\left[1-1\right]D.$
+Substituting $F=\frac{\beta\eta_{d}\Psi}{q_{6}}$ and simplifying leads to the key inequality
+$^{C}D_{t}^{\alpha}L\leq q_{2}\left(-1+\mathcal{R}_{0}\right)E.$ (15)
+When $\mathcal{R}_{0}\leq 1$, we have $^{C}D_{t}^{\alpha}L\leq 0$ throughout $\Omega$, with equality holding only when $E=0$. If $E=0$, the system dynamics imply
+${}^{C}D_{t}^{\alpha}I_{s}$ $=-q_{3}I_{s}\quad\Rightarrow\quad I_{s}\to 0,$
+${}^{C}D_{t}^{\alpha}I_{a}$ $=-q_{4}I_{a}\quad\Rightarrow\quad I_{a}\to 0,$
+${}^{C}D_{t}^{\alpha}H$ $=h_{s}I_{s}-q_{5}H\quad\Rightarrow\quad H\to 0,$
+${}^{C}D_{t}^{\alpha}D$ $=\delta_{s}I_{s}+\delta_{h}H-q_{6}D\quad\Rightarrow\quad D\to 0.$
+Therefore, the maximal invariant set where $^{C}D_{t}^{\alpha}L=0$ is exactly the disease-free equilibrium $\mathcal{E}_{0}$.
+By the fractional LaSalle invariance principle, $\mathcal{E}_{0}$ is globally asymptotically stable in $\Omega$ when $\mathcal{R}_{0}\leq 1$.
+For the case $\mathcal{R}_{0}>1$, instability follows from the local analysis, while uniform persistence can be established using standard methods for fractional dynamical systems. $\square$
 4.7 Endemic Equilibrium Analysis
-Theorem 4.5 (Existence of Endemic Equilibrium)
-The fractional-order Ebola system (2) admits a unique endemic equilibrium point $\mathcal{E}^{*}=(S^{*},V^{*},E^{*},I_{s}^{*},I_{a}^{*},H^{*},D^{*},R^{*})$ with all components strictly positive if and only if $\mathcal{R}_{0}>1$.
-Proof.  At equilibrium, the time derivatives vanish. The force of infection is:
-$\lambda^{*}=\frac{c\beta(I_{s}^{*}+\eta_{a}I_{a}^{*}+\eta_{d}D^{*})}{N^{*}}$
-The equilibrium equations provide:
-$0$ $=\Lambda-\lambda^{*}S^{*}+\omega V^{*}-q_{0}S^{*}$ (10)
-$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-q_{1}V^{*}$ (11)
-$0$ $=\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]-q_{2}E^{*}$ (12)
-$0$ $=p\sigma E^{*}-q_{3}I_{s}^{*}$ (13)
-$0$ $=(1-p)\sigma E^{*}-q_{4}I_{a}^{*}$ (14)
-$0$ $=h_{s}I_{s}^{*}-q_{5}H^{*}$ (15)
-$0$ $=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}-q_{6}D^{*}$ (16)
-$0$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}-\mu R^{*}$ (17)
-Expressing infected states in terms of $E^{*}$ as:
-$I_{s}^{*}$ $=\frac{p\sigma}{q_{3}}E^{*}\ and\quad I_{a}^{*}=\frac{(1-p)\sigma}{q_{4}}E^{*}$
-$H^{*}$ $=\frac{p\sigma h_{s}}{q_{3}q_{5}}E^{*}and\quad D^{*}=\frac{p\sigma}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)E^{*}$
-The force of infection becomes:
-$\lambda^{*}=\frac{c\beta\sigma E^{*}}{N^{*}}\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right]$
-Define the composite parameter:
-$M=c\beta\sigma\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right]$
-yielding $\lambda^{*}=ME^{*}/N^{*}$. From the exposed compartment equilibrium:
-$E^{*}=\frac{\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]}{q_{2}},$
-Uniting these relations gives the key equation:
-$1=\frac{M}{q_{2}N^{*}}[S^{*}+(1-\varepsilon)V^{*}]$
-Solving the susceptible-vaccinated subsystem:
-$S^{*}$ $=\frac{\Lambda(q_{1}+(1-\varepsilon)\lambda^{*})}{\Delta(\lambda^{*})},\quad V^{*}=\frac{\Lambda v}{\Delta(\lambda^{*})}$
-where $\Delta(\lambda^{*})=\left[(\omega+\mu+(1-\varepsilon)\lambda^{*})(\mu+v+\lambda^{*})\right]-\omega v$. This leads to $N^{*}=\Lambda/\mu$ hence the key equation becomes:
-$1=\left[\frac{M}{q_{2}(\Lambda/\mu)}\right][S^{*}+(1-\varepsilon)V^{*}]$
-At the disease-free state ($\lambda^{*}=0$), this expression equals $\mathcal{R}_{0}$.
-The monotonic decrease of $S^{*}+(1-\varepsilon)V^{*}]$ with $\lambda^{*}$ guarantees a unique positive solution $\lambda^{*}>0$ when $\mathcal{R}_{0}>1$, establishing the endemic equilibrium.
+Theorem 4.5
+The fractional-order Ebola model (2) admits a unique endemic equilibrium point $\mathcal{E}^{*}=(S^{*},V^{*},E^{*},I_{s}^{*},I_{a}^{*},H^{*},D^{*},R^{*})$ with all components strictly positive if and only if $\mathcal{R}_{0}>1$.
+Proof.  At equilibrium, all time derivatives vanish. We define the force of infection at endemic equilibrium as
+$\lambda^{*}=\frac{c\beta(I_{s}^{*}+\eta_{a}I_{a}^{*}+\eta_{d}D^{*})}{N^{*}}.$ (16)
+The equilibrium conditions yield the following system of equations
+$0$ $=\Lambda-\lambda^{*}S^{*}+\omega V^{*}-q_{0}S^{*},$ (17)
+$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-q_{1}V^{*},$ (18)
+$0$ $=\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]-q_{2}E^{*},$ (19)
+$0$ $=p\sigma E^{*}-q_{3}I_{s}^{*},$ (20)
+$0$ $=(1-p)\sigma E^{*}-q_{4}I_{a}^{*},$ (21)
+$0$ $=h_{s}I_{s}^{*}-q_{5}H^{*},$ (22)
+$0$ $=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}-q_{6}D^{*},$ (23)
+$0$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}-\mu R^{*}.$ (24)
+From equations (20)–(23), we express the infected compartments in terms of $E^{*}$
+$I_{s}^{*}$ $=\frac{p\sigma}{q_{3}}E^{*},$ (25)
+$I_{a}^{*}$ $=\frac{(1-p)\sigma}{q_{4}}E^{*},$ (26)
+$H^{*}$ $=\frac{p\sigma h_{s}}{q_{3}q_{5}}E^{*},$ (27)
+$D^{*}$ $=\frac{p\sigma}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)E^{*}.$ (28)
+Substituting these expressions into (16) gives
+$\lambda^{*}=\frac{c\beta\sigma E^{*}}{N^{*}}\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right].$ (29)
+We define the composite parameter
+$M=c\beta\sigma\left[\frac{p}{q_{3}}+\frac{\eta_{a}(1-p)}{q_{4}}+\frac{\eta_{d}p}{q_{3}q_{6}}\left(\delta_{s}+\frac{\delta_{h}h_{s}}{q_{5}}\right)\right],$ (30)
+so that $\lambda^{*}=ME^{*}/N^{*}$.
+From equation (19), we have
+$E^{*}=\frac{\lambda^{*}[S^{*}+(1-\varepsilon)V^{*}]}{q_{2}}.$ (31)
+Combining (29) and (31) yields the relation
+$1=\frac{M}{q_{2}N^{*}}[S^{*}+(1-\varepsilon)V^{*}].$ (32)
+Solving equations (17) and (18) for $S^{*}$ and $V^{*}$ in terms of $\lambda^{*}$ gives
+$S^{*}$ $=\frac{\Lambda(q_{1}+(1-\varepsilon)\lambda^{*})}{\Delta(\lambda^{*})},$ (33)
+$V^{*}$ $=\frac{\Lambda v}{\Delta(\lambda^{*})},$ (34)
+where $\Delta(\lambda^{*})=\left[(\omega+\mu+(1-\varepsilon)\lambda^{*})(\mu+v+\lambda^{*})\right]-\omega v$.
+At equilibrium, the total living population satisfies $N^{*}=\Lambda/\mu$. Substituting this and expressions (33) and (34) into (32) produces an equation in $\lambda^{*}$
+$1=\frac{M}{q_{2}(\Lambda/\mu)}\left[S^{*}+(1-\varepsilon)V^{*}\right].$ (35)
+When $\lambda^{*}=0$ (disease-free state), the left-hand side of (35) equals $\mathcal{R}_{0}$.
+Since the function $S^{*}+(1-\varepsilon)V^{*}]$ decreases monotonically with increasing $\lambda^{*}$, there exists a unique positive solution $\lambda^{*}>0$ if and only if $\mathcal{R}_{0}>1$.
+This establishes the existence and uniqueness of the endemic equilibrium when $\mathcal{R}_{0}>1$.
 $\square$
@@
 Theorem 4.6
-Consider system (2). When $\mathcal{R}_{0}>1$, the endemic equilibrium $\mathcal{E}^{*}$ is locally asymptotically stable.
-Proof.  The Jacobian matrix evaluated at $\mathcal{E}^{*}$ exhibits the block structure:
-$J(\mathcal{E}^{*})=\begin{pmatrix}J_{11}&J_{12}\\ J_{21}&J_{22}\end{pmatrix}$
-with components:
-$J_{11}$ $=\begin{pmatrix}-(\lambda^{*}+q_{0})&\omega\\ v&-\lambda^{*}+\varepsilon\lambda^{*}-q_{1})\end{pmatrix}$,
+When $\mathcal{R}_{0}>1$, the endemic equilibrium $\mathcal{E}^{*}$ of Ebola model (2) is locally asymptotically stable.
+Proof.  To analyze the local stability of $\mathcal{E}^{*}$, we evaluate the Jacobian matrix at this equilibrium point. The Jacobian exhibits the block structure
+$J(\mathcal{E}^{*})=\begin{pmatrix}J_{11}&J_{12}\\ J_{21}&J_{22}\end{pmatrix}$,
+where the submatrices are defined as follows
+$J_{11}$ $=\begin{pmatrix}-(\lambda^{*}+q_{0})&\omega\\ v&-((1-\varepsilon)\lambda^{*}+q_{1})\end{pmatrix}$,
 $J_{12}$ $=\begin{pmatrix}0&-\beta S^{*}\Phi&-\beta\eta_{a}S^{*}\Phi&0&-\beta\eta_{d}S^{*}\Phi&0\\ 0&-\beta(1-\varepsilon)V^{*}\Phi&-\beta\eta_{a}(1-\varepsilon)V^{*}\Phi&0&-\beta\eta_{d}(1-\varepsilon)V^{*}\Phi&0\end{pmatrix}$,
 $J_{21}$ $=\begin{pmatrix}\lambda^{*}&(1-\varepsilon)\lambda^{*}\\ 0&0\\ 0&0\\ 0&0\\ 0&0\\ 0&0\end{pmatrix}$,
-$J_{22}$ $=\begin{pmatrix}-q_{2}&\beta\Psi\Phi&\beta\eta_{a}\Psi\Phi&0&\beta\eta_{d}\Psi\Phi&0\\ p\sigma&-q_{3}&0&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0&0\\ 0&h_{s}&0&-q_{5}&0&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}&0\\ 0&\gamma_{s}&\gamma_{a}&\gamma_{h}&0&-\mu\end{pmatrix}$
-where $\Psi=S^{*}+(1-\varepsilon)V^{*}$, $\Phi=c/N^{*}$, and $\lambda^{*}$ denotes the equilibrium force of infection. The characteristic polynomial is an eighth-degree equation:
-$P(\lambda)=\lambda^{8}+a_{7}\lambda^{7}+a_{6}\lambda^{6}+a_{5}\lambda^{5}+a_{4}\lambda^{4}+a_{3}\lambda^{3}+a_{2}\lambda^{2}+a_{1}\lambda+a_{0}=0$
-Consider the Metzler matrix $Q=-J(\mathcal{E}^{*})$. For $\mathcal{R}_{0}>1$, $Q$ is irreducible and satisfies:
-  • Diagonal positivity: $q_{ii}>0$
-  • Diagonal dominance: $q_{ii}\geq\sum_{j\neq i}|q_{ij}|$
-  • Existence of positive vector $v>0$ with $Qv>0$
-According to the Perron-Frobenius theorem, the eigenvalues of $Q$ have positive real parts, and thus the eigenvalues of $J(\mathcal{E}^{*})=-Q$ have negative real parts.
-The fractional stability condition $|\arg(\lambda_{i})|>\alpha\pi/2$ is satisfied, because $|\arg(\lambda_{i})|=\pi>\alpha\pi/2$ for $\alpha\in(0,1]$.
-Stability conditions are confirmed via Routh-Hurwitz’s analysis technique if all the principal minors of the Hurwitz matrix associated with the system are strictly positive.
-This guarantees that all eigenvalues of the characteristic equation have negative real parts.
-The result stems from the fractional-order stability theorem [48], which proves that the system is asymptotically stable.
+$J_{22}$ $=\begin{pmatrix}-q_{2}&\beta\Psi\Phi&\beta\eta_{a}\Psi\Phi&0&\beta\eta_{d}\Psi\Phi&0\\ p\sigma&-q_{3}&0&0&0&0\\ (1-p)\sigma&0&-q_{4}&0&0&0\\ 0&h_{s}&0&-q_{5}&0&0\\ 0&\delta_{s}&0&\delta_{h}&-q_{6}&0\\ 0&\gamma_{s}&\gamma_{a}&\gamma_{h}&0&-\mu\end{pmatrix}$.
+Here $\Psi=S^{*}+(1-\varepsilon)V^{*}$, $\Phi=c/N^{*}$, and $\lambda^{*}$ denotes the equilibrium force of infection defined in (16).
+The characteristic polynomial of $J(\mathcal{E}^{*})$ takes the form
+$P(\lambda)=\lambda^{8}+a_{7}\lambda^{7}+a_{6}\lambda^{6}+a_{5}\lambda^{5}+a_{4}\lambda^{4}+a_{3}\lambda^{3}+a_{2}\lambda^{2}+a_{1}\lambda+a_{0}=0,$ (36)
+where the coefficients $a_{i}$ ($i=0,\dots,7$) are determined by the model parameters.
+Consider the Metzler matrix $Q=-J(\mathcal{E}^{*})$. For $\mathcal{R}_{0}>1$, the matrix $Q$ is irreducible and satisfies the following properties
+  • All diagonal entries are strictly positive,
+  • $Q$ is row diagonally dominant with $q_{ii}\geq\sum_{j\neq i}|q_{ij}|$ for each row $i$,
+  • There exists a positive vector $\mathbf{v}>\mathbf{0}$ such that $Q\mathbf{v}>\mathbf{0}$.
+By the Perron-Frobenius theorem for Metzler matrices [54], all eigenvalues of $Q$ have positive real parts.
+Consequently, the eigenvalues of $J(\mathcal{E}^{*})=-Q$ have negative real parts.
+For the fractional-order system with $\alpha\in(0,1]$, the stability condition requires $|\arg(\lambda_{i})|>\alpha\pi/2$ for each eigenvalue $\lambda_{i}$ of $J(\mathcal{E}^{*})$.
+Since all eigenvalues have negative real parts, their arguments satisfy $|\arg(\lambda_{i})|=\pi>\alpha\pi/2$.
+This conclusion is further supported by the Routh-Hurwitz criterion applied to the characteristic polynomial (36).
+When $\mathcal{R}_{0}>1$, all principal minors of the associated Hurwitz matrix are strictly positive, ensuring that every eigenvalue of $J(\mathcal{E}^{*})$ has a negative real part.
+By the fractional-order stability theorem, the endemic equilibrium $\mathcal{E}^{*}$ is therefore locally asymptotically stable.
 $\square$
-4.10 Global Stability of the Endemic Equilibrium
+4.9 GAS of Endemic Equilibrium
 Theorem 4.7
-For values of $\mathcal{R}_{0}$ greater than 1, the endemic equilibrium point $\mathcal{E}^{*}$ of system (2) is globally asymptotically stable in the interior of the invariant region $\Omega$.
-Proof.  A Lyapunov function can be constructed with the use of the logarithmic Volterra-type form as follows:
-$\mathcal{V}=$ $\left(S-S^{*}-S^{*}\ln\frac{S}{S^{*}}\right)+\left(V-V^{*}-V^{*}\ln\frac{V}{V^{*}}\right)$
-$+K_{1}\left(E-E^{*}-E^{*}\ln\frac{E}{E^{*}}\right)+K_{2}\left(I_{s}-I_{s}^{*}-I_{s}^{*}\ln\frac{I_{s}}{I_{s}^{*}}\right)$
-$+K_{3}\left(I_{a}-I_{a}^{*}-I_{a}^{*}\ln\frac{I_{a}}{I_{a}^{*}}\right)+K_{4}\left(H-H^{*}-H^{*}\ln\frac{H}{H^{*}}\right)$
-$+K_{5}\left(D-D^{*}-D^{*}\ln\frac{D}{D^{*}}\right)+K_{6}\left(R-R^{*}-R^{*}\ln\frac{R}{R^{*}}\right)$
-with $K_{1},K_{2},K_{3},K_{4},K_{5},K_{6}>0$ to be determined subsequently Capital, along the system trajectories, the Caputo derivative produces:
-${}^{C}D_{t}^{\alpha}\mathcal{V}=$ $\left(1-\frac{S^{*}}{S}\right){}^{C}D_{t}^{\alpha}S+\left(1-\frac{V^{*}}{V}\right){}^{C}D_{t}^{\alpha}V+K_{1}\left(1-\frac{E^{*}}{E}\right){}^{C}D_{t}^{\alpha}E$
-$+K_{2}\left(1-\frac{I_{s}^{*}}{I_{s}}\right){}^{C}D_{t}^{\alpha}I_{s}+K_{3}\left(1-\frac{I_{a}^{*}}{I_{a}}\right){}^{C}D_{t}^{\alpha}I_{a}+K_{4}\left(1-\frac{H^{*}}{H}\right){}^{C}D_{t}^{\alpha}H$
-$+K_{5}\left(1-\frac{D^{*}}{D}\right){}^{C}D_{t}^{\alpha}D+K_{6}\left(1-\frac{R^{*}}{R}\right){}^{C}D_{t}^{\alpha}R$
-Replacing the model equations while utilizing endemic equilibrium conditions:
-$\Lambda$ $=\lambda^{*}S^{*}+\omega V^{*}-(\mu+v)S^{*}$
-$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-(\omega+\mu)V^{*}$
-$q_{2}E^{*}$ $=\lambda^{*}(S^{*}+(1-\varepsilon)V^{*})$
-$q_{3}I_{s}^{*}$ $=p\sigma E^{*},\quad q_{4}I_{a}^{*}=(1-p)\sigma E^{*}$
-$q_{5}H^{*}$ $=h_{s}I_{s}^{*},\quad\mu_{d}D^{*}=\delta_{s}I_{s}^{*}+\delta_{h}H^{*}$
-$\mu R^{*}$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}$
-We take coefficients as:
-$K_{1}$ $=1,\quad K_{2}=\frac{\lambda^{*}S^{*}}{q_{3}I_{s}^{*}},\quad K_{3}=\frac{\lambda^{*}S^{*}\eta_{a}}{q_{4}I_{a}^{*}},$
-$K_{4}$ $=\frac{\lambda^{*}S^{*}\eta_{d}\delta_{h}}{q_{5}q_{6}H^{*}},\quad K_{5}=\frac{\lambda^{*}S^{*}\eta_{d}}{q_{6}D^{*}},\quad K_{6}=1$
-By means of grouping the terms accordingly, we end up with the inequality:
-$^{C}D_{t}^{\alpha}\mathcal{V}\leq\lambda^{*}S^{*}\left[6-\frac{S^{*}}{S}-\frac{S}{S^{*}}-\frac{EI_{s}^{*}}{E^{*}I_{s}}-\frac{I_{s}E^{*}}{I_{s}^{*}E}-\frac{DH^{*}}{D^{*}H}-\frac{HD^{*}}{H^{*}D}\right]+\Psi(S,V,E,I_{s},I_{a},H,D,R)$
-where $\Psi$ indicates a negative definite function. From the arithmetic-geometric mean inequality:
-$\frac{S^{*}}{S}+\frac{S}{S^{*}}+\frac{EI_{s}^{*}}{E^{*}I_{s}}+\frac{I_{s}E^{*}}{I_{s}^{*}E}+\frac{DH^{*}}{D^{*}H}+\frac{HD^{*}}{H^{*}D}\geq 6$
-Equality is achieved only when $S=S^{*},E=E^{*},I_{s}=I_{s}^{*},H=H^{*},D=D^{*}$.
-We can conclude that $^{C}D_{t}^{\alpha}\mathcal{V}\leq 0$ holds true for all points in the feasible region $\Omega$.
-It is only at the endemic equilibrium $\mathcal{E}^{*}$ that $^{C}D_{t}^{\alpha}\mathcal{V}=0$.
-This allows us to conclude the global asymptotic stability of $\mathcal{E}^{*}$ using the fractional LaSalle invariance principle [48].
+If $\mathcal{R}_{0}>1$, the endemic equilibrium $\mathcal{E}^{*}$ of the model (2) is globally asymptotically stable in the interior of the invariant region $\Omega$.
+Proof.  We construct a Lyapunov function of logarithmic Volterra type
+$\mathcal{V}$ $=\left(S-S^{*}-S^{*}\ln\frac{S}{S^{*}}\right)+\left(V-V^{*}-V^{*}\ln\frac{V}{V^{*}}\right)$ (37)
+$\quad+K_{1}\left(E-E^{*}-E^{*}\ln\frac{E}{E^{*}}\right)+K_{2}\left(I_{s}-I_{s}^{*}-I_{s}^{*}\ln\frac{I_{s}}{I_{s}^{*}}\right)$
+$\quad+K_{3}\left(I_{a}-I_{a}^{*}-I_{a}^{*}\ln\frac{I_{a}}{I_{a}^{*}}\right)+K_{4}\left(H-H^{*}-H^{*}\ln\frac{H}{H^{*}}\right)$
+$\quad+K_{5}\left(D-D^{*}-D^{*}\ln\frac{D}{D^{*}}\right)+K_{6}\left(R-R^{*}-R^{*}\ln\frac{R}{R^{*}}\right),$
+where the constants $K_{1},K_{2},K_{3},K_{4},K_{5},K_{6}>0$ will be determined.
+Computing the Caputo derivative along the system trajectories yields
+${}^{C}D_{t}^{\alpha}\mathcal{V}$ $=\left(1-\frac{S^{*}}{S}\right){}^{C}D_{t}^{\alpha}S+\left(1-\frac{V^{*}}{V}\right){}^{C}D_{t}^{\alpha}V+K_{1}\left(1-\frac{E^{*}}{E}\right){}^{C}D_{t}^{\alpha}E$ (38)
+$\quad+K_{2}\left(1-\frac{I_{s}^{*}}{I_{s}}\right){}^{C}D_{t}^{\alpha}I_{s}+K_{3}\left(1-\frac{I_{a}^{*}}{I_{a}}\right){}^{C}D_{t}^{\alpha}I_{a}+K_{4}\left(1-\frac{H^{*}}{H}\right){}^{C}D_{t}^{\alpha}H$
+$\quad+K_{5}\left(1-\frac{D^{*}}{D}\right){}^{C}D_{t}^{\alpha}D+K_{6}\left(1-\frac{R^{*}}{R}\right){}^{C}D_{t}^{\alpha}R.$
+We substitute the model equations from (2) and use the endemic equilibrium conditions
+$\Lambda$ $=\lambda^{*}S^{*}+\omega V^{*}-(\mu+v)S^{*},$ (39)
+$0$ $=vS^{*}-(1-\varepsilon)\lambda^{*}V^{*}-(\omega+\mu)V^{*},$ (40)
+$q_{2}E^{*}$ $=\lambda^{*}(S^{*}+(1-\varepsilon)V^{*}),$ (41)
+$q_{3}I_{s}^{*}$ $=p\sigma E^{*},\quad q_{4}I_{a}^{*}=(1-p)\sigma E^{*},$ (42)
+$q_{5}H^{*}$ $=h_{s}I_{s}^{*},\quad\mu_{d}D^{*}=\delta_{s}I_{s}^{*}+\delta_{h}H^{*},$ (43)
+$\mu R^{*}$ $=\gamma_{s}I_{s}^{*}+\gamma_{a}I_{a}^{*}+\gamma_{h}H^{*}.$ (44)
+After extensive algebraic manipulation and grouping of terms, we select the coefficients
+$K_{1}$ $=1,\quad K_{2}=\frac{\lambda^{*}S^{*}}{q_{3}I_{s}^{*}},\quad K_{3}=\frac{\lambda^{*}S^{*}\eta_{a}}{q_{4}I_{a}^{*}},$ (45)
+$K_{4}$ $=\frac{\lambda^{*}S^{*}\eta_{d}\delta_{h}}{q_{5}q_{6}H^{*}},\quad K_{5}=\frac{\lambda^{*}S^{*}\eta_{d}}{q_{6}D^{*}},\quad K_{6}=1.$
+With these choices, the derivative simplifies to
+$^{C}D_{t}^{\alpha}\mathcal{V}\leq\lambda^{*}S^{*}\left[6-\frac{S^{*}}{S}-\frac{S}{S^{*}}-\frac{EI_{s}^{*}}{E^{*}I_{s}}-\frac{I_{s}E^{*}}{I_{s}^{*}E}-\frac{DH^{*}}{D^{*}H}-\frac{HD^{*}}{H^{*}D}\right]+\Psi(S,V,E,I_{s},I_{a},H,D,R),$ (46)
+where $\Psi$ is a negative definite function.
+Applying the arithmetic-geometric mean inequality gives
+$\frac{S^{*}}{S}+\frac{S}{S^{*}}+\frac{EI_{s}^{*}}{E^{*}I_{s}}+\frac{I_{s}E^{*}}{I_{s}^{*}E}+\frac{DH^{*}}{D^{*}H}+\frac{HD^{*}}{H^{*}D}\geq 6,$ (47)
+with equality holding only when $S=S^{*}$, $E=E^{*}$, $I_{s}=I_{s}^{*}$, $H=H^{*}$, and $D=D^{*}$.
+Consequently, $^{C}D_{t}^{\alpha}\mathcal{V}\leq 0$ throughout $\Omega$, and $^{C}D_{t}^{\alpha}\mathcal{V}=0$ only at the endemic equilibrium $\mathcal{E}^{*}$.
+By the fractional LaSalle invariance principle [47], $\mathcal{E}^{*}$ is globally asymptotically stable in the interior of $\Omega$.
 $\square$
 Corollary 4.8
-The fractional-order Ebola system (2) experiences a forward bifurcation at $\mathcal{R}_{0}=1$, without showing a backward bifurcation and without exhibiting multiple endemic equilibria.
-Proof.
-Global stability properties—disease-free equilibrium whenever $\mathcal{R}_{0}\leq 1$ and unique endemic equilibrium when $\mathcal{R}_{0}>1$—contradict the possibility of backward bifurcation.
-The epidemiological dynamics displays classical threshold behavior: disease elimination befalls when $\mathcal{R}_{0}\leq 1$ and $\mathcal{R}_{0}>1$ indicates endemic tenacity.
-$\square$
-5 Sensitivity Analysis
-5.1 Global Sensitivity Analysis and Parameter Influence
-From the data, Latin Hypercube Sampling (LHS) estimates $\mathcal{R}_{0}$ and other critical epidemiological characteristics and examines the correlation of the parameters using PRCC.
-The PRCC examined global sensitivity for 10,000 samples within the reasonable biological boundaries.
-The outcomes are summarized in Table 3.
-Table 3: Global sensitivity analysis: Parameter ranges and PRCC values for $\mathcal{R}_{0}$ and peak infection dynamics
-Parameter Lower Bound Upper Bound PRCC ($\mathcal{R}_{0}$) PRCC (Peak $I_{s}$) p-value
-$\beta$ 0.15 0.35 0.92 0.88 $p<0.001$
-$\sigma$ 1/12 1/8 0.85 0.81 $p<0.001$
-$\eta_{d}$ 0.4 0.9 0.68 0.65 $p<0.001$
-pp 0.6 0.85 0.52 0.49 $p<0.001$
-$\eta_{a}$ 0.3 0.7 0.41 0.38 $p<0.001$
-$\varepsilon$ 0.7 0.95 -0.61 -0.58 $p<0.001$
-$\gamma_{s}$ 1/16 1/12 -0.45 -0.42 $p<0.001$
-$v$ 0.01 0.08 -0.35 -0.32 $p<0.001$
-$\mu_{d}$ 0.1 0.3 -0.28 -0.25 $p<0.001$
-$\alpha$ 0.7 0.95 -0.25 -0.23 $p<0.001$
-$h_{s}$ 0.2 0.6 -0.21 -0.19 $p<0.001$
-$\delta_{s}$ 0.05 0.12 0.15 0.18 $p=0.002$
-5.2 Analytical Sensitivity Framework
-The normalized sensitivity index quantifying $\mathcal{R}_{0}$ dependence on parameter $p$ is defined as:
-$\Upsilon_{p}^{\mathcal{R}_{0}}=\frac{\partial\mathcal{R}_{0}}{\partial p}\times\frac{p}{\mathcal{R}_{0}}$ (18)
-Theorem 5.1 (Analytical Sensitivity Indices)
-The sensitivity indices for dominant parameters are given by:
-Proof.  Analytical differentiation of the reproduction number yields.
-$\text{Sensitivity indices for }\mathcal{R}_{0}:$ (19)
-$\Upsilon_{\beta}^{\mathcal{R}_{0}}=+1$
-$\Upsilon_{\sigma}^{\mathcal{R}_{0}}=1+\frac{\mathcal{R}_{0}^{(1)}}{\mathcal{R}_{0}}$
-$\Upsilon_{\eta_{a}}^{\mathcal{R}_{0}}=\frac{\mathcal{R}_{0}^{(I_{a})}}{\mathcal{R}_{0}}$
-$\Upsilon_{\eta_{d}}^{\mathcal{R}_{0}}=\frac{\mathcal{R}_{0}^{(D)}}{\mathcal{R}_{0}}$
-$\Upsilon_{\varepsilon}^{\mathcal{R}_{0}}=-\frac{V^{*}}{S^{*}+(1-\varepsilon)V^{*}}$
-$\Upsilon_{v}^{\mathcal{R}_{0}}=\frac{\beta\sigma}{\mathcal{R}_{0}(\Lambda/\mu)q_{2}}\cdot\frac{\partial}{\partial v}[S^{*}+(1-\varepsilon)V^{*}]\cdot K,$
-The following equations illustrate the different components that contribute to the basic reproduction number:
-$\mathcal{R}_{0}^{(I_{a})}$ $=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\cdot\frac{\eta_{a}(1-p)}{q_{4}}$ (20)
-$\mathcal{R}_{0}^{(D)}$ $=\frac{\beta\sigma[S^{*}+(1-\varepsilon)V^{*}]}{(\Lambda/\mu)q_{2}}\cdot\frac{p\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{3}q_{5}q_{6}}$ (21)
-$\frac{\partial}{\partial v}[S^{*}+(1-\varepsilon)V^{*}]$ $=-\frac{\Lambda\varepsilon(\omega+\mu)}{\mu(q_{1}+v)^{2}}$ (22)
-with the composite transmission factor given by
-$K=\frac{p}{q_{3}}\left(1+\frac{\eta_{d}(\delta_{s}q_{5}+\delta_{h}h_{s})}{q_{5}q_{6}}\right)+\frac{\eta_{a}(1-p)}{q_{4}}.$ (23)
-From equations (20) and (21) one may discern the individual contributions of asymptomatic infections and post-mortem transmission to the total basic reproduction number, $\mathcal{R}_{0}$. $\square$
-5.3 Sensitivity Analysis and Numerical Results
-Remark 5.1 (Parameter Influence Hierarchy)
-Analytical and PRCC sensitivity analyses establish the following parameter influence hierarchy on $\mathcal{R}_{0}$:
-$\text{High Positive: }\beta>\sigma>\eta_{d}$
-$\text{Moderate Positive: }p>\eta_{a}$
-$\text{High Negative: }\varepsilon$
-$\text{Moderate Negative: }\gamma_{s}>v$
-$\text{Low Negative: }\mu_{d}>\alpha>h_{s}$
-Classification follows PRCC thresholds: highly influential ($|\text{PRCC}|>0.7$), moderately influential ($0.5<|\text{PRCC}|\leq 0.7$), and less influential ($|\text{PRCC}|\leq 0.5$).
-Incorporating targeted integrated strategies aimed at reducing transmission ($\beta$), targeted rapid detection ($\sigma$), and safe burial ($\eta_{d}$) achieve an 85% reduction in outbreaks according to numerical simulations which is a significant increase from vaccination-only strategies that only attain a 45% reduction.
-(a) Threshold dynamics verification
-(b) Parameter sensitivity analysis
-(c) PRCC ranking results
-Figure 5: Numerical simulations validating theoretical findings and quantifying parameter impacts
-Table 4: Evidence-based intervention prioritization from sensitivity analysis
-Priority Interventions Impact
-Primary Contact reduction, Safe burial, Early detection 60-85% $\mathcal{R}_{0}$ reduction
-Secondary High-efficacy vaccination, Treatment access 25-45% $\mathcal{R}_{0}$ reduction
-Tertiary Mass vaccination, Hospital capacity 10-20% $\mathcal{R}_{0}$ reduction
-Figure 6: Contours of parameter sensitivity: (a) $\beta,\sigma$ (most sensitive), (b) $\eta_{d},p$ (intermediate), (c) $\varepsilon,\gamma_{s}$ (least sensitive), (d) $\eta_{a},v$ (mixed sensitivity)
-The results further affirm the developed theory by demonstrating very strong threshold behavior at $\mathcal{R}_{0}=1$.
-Such confirmation also supports the parameter sweep results which assert the dominance of the transmission parameters $\beta$, $\sigma$, and $\eta_{d}$, in the dynamics of the outbreaks.
-Contour plots illustrate the nonlinear interactions which add to the ever-growing complexity of the transmission dynamics of the Ebola virus.
-The primary community-based switch strategy employed to control the spread of the virus focuses on the interruption of transmission.
-Thus, the safe burial practices endorsed by the World Health Organization are critical in a community’s efforts to disturb transmission.
-Evidence of memory in the Ebola virus, captured by the sophisticated model elements, is indicated by the strong effect of the fractional order $\alpha$.
-Such results, in turn, offer the mathematical foundation for controlled community-centric Ebola response which is often lacking.
-This model control will inform the aggressive response necessary for targeted intervention strategies to control the Ebola virus.
+The fractional-order Ebola model (2) undergoes a forward bifurcation at $\mathcal{R}_{0}=1$.
+This bifurcation is characterized by the absence of backward bifurcation and the nonexistence of multiple endemic equilibria.
+4.10 Sensitivity Analysis Results
+Figure 3: Sensitivity analysis of parameters on basic reproduction number $\mathcal{R}_{0}$ for the Ebola model.
+Red bars indicate negative sensitivity (parameters that decrease $\mathcal{R}_{0}$), blue bars indicate positive sensitivity (parameters that increase $\mathcal{R}_{0}$).
+Table 2: Sensitivity analysis of basic reproduction number ($\mathcal{R}_{0}$) for the Ebola model
+Parameter $\Upsilon$ Effect Parameter $\Upsilon$ Effect
+$\beta$ +1.0000 High + $\varepsilon$ -0.4123 Medium –
+$\sigma$ +0.6250 High + $p$ +0.3752 Medium +
+$\eta_{d}$ +0.4678 High + $v$ -0.3561 Medium –
+$\eta_{a}$ +0.3117 Medium + $h_{s}$ -0.1987 Low –
+$\gamma_{s}$ -0.2541 Medium – $\gamma_{a}$ -0.1876 Low –
+$\delta_{s}$ +0.2345 Medium + $\alpha$ -0.1250 Low –
+$\omega$ +0.0942 Low + $\mu$ -0.0214 Very low –
+$\delta_{h}$ +0.0783 Very low + $\mu_{d}$ -0.0187 Very low –
+$\Lambda$ +0.0000 No effect
+The sensitivity analysis results, presented in Table 2, reveal the normalized sensitivity indices of model parameters with respect to $\mathcal{R}_{0}$.
+A positive index indicates that $\mathcal{R}_{0}$ increases with the parameter, while a negative index corresponds to a decrease in $\mathcal{R}_{0}$.
+Parameters with negative sensitivity indices ($\varepsilon$, $v$, $\gamma_{s}$, $h_{s}$, $\gamma_{a}$, $\alpha$, $\mu$, $\mu_{d}$) reduce $\mathcal{R}_{0}$ when increased, thereby suppressing disease transmission.
+Conversely, parameters with positive indices ($\beta$, $\sigma$, $\eta_{d}$, $p$, $\eta_{a}$, $\delta_{s}$, $\omega$, $\delta_{h}$) enhance $\mathcal{R}_{0}$ when increased, leading to increased transmission potential.
+The transmission rate $\beta$ exhibits the highest positive sensitivity index (+1.0000), indicating that $\mathcal{R}_{0}$ is directly proportional to $\beta$.
+The significance of contact reduction measures like social distancing and personal protection gear is highlighted by this.
+The crucial importance of secure burial practices provides for controlling outbreaks is demonstrated by the dead transmission rate $\eta_{d}$ (index +0.4678).
+Even at moderate coverage levels, vaccinations may reduce transmission, as indicated by the negative sensitivity (-0.4123) of vaccine efficacy $\varepsilon$.
+Interestingly, the hospitalization rate \$h_s\$ has a negative sensitivity index (-0.1987), indicating that through isolating infectious persons, expanding hospital capacity lowers transmission.
+The fractional order $\alpha$ displays low sensitivity (-0.1250), suggesting that decreased transmission potential in the fractional-order model is an outcome of memory effects.
+The recruitment rate $\Lambda$ has zero sensitivity, as expected from the structure of $\mathcal{R}_{0}$, indicating that population turnover does not directly affect the reproduction number in this model formulation.
+Figure 3 displays a tornado plot of the sensitivity indices, ranking parameters by their influence on $\mathcal{R}_{0}$.
+The dominance of $\beta$, $\sigma$, and $\eta_{d}$ as the top three positive-sensitivity parameters reinforces the importance of transmission reduction, rapid case detection, and safe burial practices in Ebola control strategies.
 5 Optimal Control Analysis
-6.1 Problem Formulation
-In this section, we delineate four time-dependent control variables for a complete management of the Ebola virus.
-Control variables: $u_{1}$ protection [0,0.8], $u_{2}$ vaccination [0,0.15], $u_{3}$ treatment [0,0.8], $u_{4}$ burial [0,0.5] (day-1).
-The controlled fractional-order system now becomes:
+In this section, we introduce four time-dependent control variables to manage Ebola transmission effectively.
+The controls are personal protection ($u_{1}$, range [0,0.8]), vaccination ($u_{2}$, range [0,0.15]), treatment ($u_{3}$, range [0,0.8]), and safe burial ($u_{4}$, range [0,0.5]) (all in units day-1).
+The controlled fractional-order system is given by
 ${}^{C}D_{t}^{\alpha}S(t)$ $=\Lambda-(1-u_{1})\lambda S+\omega V-(\mu+u_{2})S,$ (48)
@@
 ${}^{C}D_{t}^{\alpha}D(t)$ $=\delta_{s}I_{s}+\delta_{h}H-(u_{4}+\mu_{d})D,$
-${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R.$
-with modified force of infection $\lambda(t)=\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)/N$. The objective functional minimizes total cost over $[0,T]$:
-$J(\mathbf{u})=\int_{0}^{T}\left[\sum_{i=1}^{4}A_{i}X_{i}(t)+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}(t)\right]dt$ (25)
-Where $X=[I_{s},I_{a},H,D]$ indicates the infectious compartments, diseased burden costs are denoted by $A_{i}>0$, and the costs related to the implementation of controls are denoted as $B_{j}>0$ in quadratic form.
-We look to find the optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that fulfill the requirements of:
-$J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u})$ (26)
-for the controlled system (24), where the admissible set is:
-$\mathcal{U}=\left\{\mathbf{u}\in L^{\infty}(0,T):0\leq u_{i}(t)\leq u_{i}^{\max}\right\}$ (27)
-6.2 Necessary Conditions for Optimality
-Theorem 6.1 (Existence of Optimal Controls)
-Let $J(\mathbf{u})$ be the objective functional relative to the controlled fractional-order system (24) with controllable initial conditions $\mathbf{X}(0)\in\mathbb{R}^{8}_{\geq 0}$, and with $\mathbf{u}\in\mathcal{U}$ being the admissible controls.
-An optimal control pair $(\mathbf{X}^{*},\mathbf{u}^{*})$ such that $J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u})$ exists.
+${}^{C}D_{t}^{\alpha}R(t)$ $=\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R,$
+with the modified force of infection $\lambda(t)=\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)/N$. The objective functional to be minimized over the time horizon $[0,T]$ is
+$J(\mathbf{u})=\int_{0}^{T}\left[\sum_{i=1}^{4}A_{i}X_{i}(t)+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}(t)\right]dt,$ (49)
+where $X=[I_{s},I_{a},H,D]$ represents the infectious compartments, $A_{i}>0$ are the cost weights associated with the disease burden, and $B_{j}>0$ are the quadratic cost coefficients for implementing the controls.
+We seek the optimal controls $\mathbf{u}^{*}=(u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*})$ that satisfy
+$J(\mathbf{u}^{*})=\min_{\mathbf{u}\in\mathcal{U}}J(\mathbf{u}),$ (50)
+subject to the dynamics (48), where the admissible control set is
+$\mathcal{U}=\left\{\mathbf{u}\in L^{\infty}(0,T):0\leq u_{i}(t)\leq u_{i}^{\max}\right\}.$ (51)
+Theorem 5.1
+Let the fractional-order controlled system (48) with initial conditions $\mathbf{X}(0)\in\mathbb{R}^{8}_{\geq 0}$ and admissible control set $\mathcal{U}$ satisfy
+  1. (i)
+    The control set $\mathcal{U}$ is closed, convex, and bounded in $L^{\infty}(0,T)$.
+  2. (ii)
+    The right-hand side $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ of (48) is continuous in $(t,\mathbf{X},\mathbf{u})$, measurable in $t$, and satisfies as
+    • Linear growth $\|\mathbf{f}(t,\mathbf{X},\mathbf{u})\|\leq c_{1}+c_{2}\|\mathbf{X}\|+c_{3}\|\mathbf{u}\|$ for constants $c_{1},c_{2},c_{3}>0$.
+    • Lipschitz continuity $\|\mathbf{f}(t,\mathbf{X}_{1},\mathbf{u})-\mathbf{f}(t,\mathbf{X}_{2},\mathbf{u})\|\leq L\|\mathbf{X}_{1}-\mathbf{X}_{2}\|$ for some $L>0$.
+  3. (iii)
+    The integrand $\mathcal{L}(t,\mathbf{X},\mathbf{u})=\sum_{i=1}^{4}A_{i}X_{i}(t)+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}(t)$ of the objective functional (49) is convex in $\mathbf{u}$ and satisfies the coercivity condition $\mathcal{L}(t,\mathbf{X},\mathbf{u})\geq c_{4}\|\mathbf{u}\|^{2}-c_{5}$ for constants $c_{4}>0$, $c_{5}\geq 0$.
+Then there exists an optimal control pair $(\mathbf{X}^{*},\mathbf{u}^{*})$ that minimizes $J(\mathbf{u})$ subject to (48).
 Proof.
-Existence has been demonstrated by verifying the fundamental conditions in [56].
-The state space $\mathbb{R}^{8}_{\geq 0}$ is closed and the control set $\mathcal{U}$ is closed, convex and bounded (as per its definition).
-The system dynamics $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ has the required properties as follows:
-The system dynamics $\mathbf{f}(t,\mathbf{X},\mathbf{u})$ are continuous in $(t,\mathbf{X},\mathbf{u})$, is bounded such that $\|\mathbf{f}\|\leq C(1+\|\mathbf{X}\|)$, and is Lipschitz continuous $\|\mathbf{f}(\mathbf{X}_{1})-\mathbf{f}(\mathbf{X}_{2})\|\leq L\|\mathbf{X}_{1}-\mathbf{X}_{2}\|.$ The cost integrand $\mathcal{L}(t,\mathbf{X},\mathbf{u})$ exhibits convexity in $\mathbf{u}$ through quadratic terms and coercivity $\mathcal{L}\geq\frac{1}{2}\min B_{j}\|\mathbf{u}\|^{2}-K$.
-The inequality $^{C}D_{t}^{\alpha}N(t)\leq\Lambda-\mu N$ leads to population boundedness, ensuring that $N(t)\leq\max\{N(0),\Lambda/\mu\}$, and thus the state trajectories are bounded.
-Together, these conditions guarantee the existence of an optimal control pair.
+We verify the conditions of the Filippov-Cesari existence theorem for fractional optimal control problems [55, 56].
+By definition, $\mathcal{U}=\{\mathbf{u}\in L^{\infty}(0,T):0\leq u_{i}(t)\leq u_{i}^{\max}\}$ is closed, convex, and bounded in $L^{\infty}(0,T)$.
+The right-hand side $\mathbf{f}$ of (48) is polynomial in $\mathbf{X}$ and $\mathbf{u}$, hence continuous and measurable.
+For the linear growth condition, note that all terms are at most quadratic (products like $(1-u_{1})\lambda S$ where $\lambda$ is linear in $I_{s},I_{a},D$).
+Since $N(t)\geq\mu S^{*}>0$ (by Theorem 1) and the population is bounded by $\max\{N(0),\Lambda/\mu\}$, there exists $M>0$ such that $\|\mathbf{f}(t,\mathbf{X},\mathbf{u})\|\leq M(1+\|\mathbf{X}\|+\|\mathbf{u}\|)$.
+Lipschitz continuity follows from the boundedness of $\mathbf{u}$ and the fact that all partial derivatives $\partial\mathbf{f}/\partial\mathbf{X}$ are bounded on bounded sets.
+The integrand $\mathcal{L}$ is quadratic in $\mathbf{u}$, hence convex in $\mathbf{u}$.
+Moreover, with $B_{\min}=\min\{B_{1},B_{2},B_{3},B_{4}\}>0$, we have $\mathcal{L}(t,\mathbf{X},\mathbf{u})\geq\frac{1}{2}B_{\min}\|\mathbf{u}\|^{2}$, satisfying the coercivity condition.
+From the population dynamics $^{C}D_{t}^{\alpha}N\leq\Lambda-\mu N$, we have $N(t)\leq\max\{N(0),\Lambda/\mu\}$.
+Since all compartments are nonnegative and $N=S+V+E+I_{s}+I_{a}+H+D+R$, each compartment is bounded by this maximum.
+All conditions of Theorem 2.1 in [55] (existence theorem for fractional optimal control problems) are satisfied.
+Therefore, there exists an optimal control pair $(\mathbf{X}^{*},\mathbf{u}^{*})$ minimizing $J(\mathbf{u})$.
 $\square$
-6.3 Pontryagin’s Maximum Principle for Fractional Systems
 Theorem 5.2
-Assuming that $(\mathbf{X}^{*},\mathbf{u}^{*})$ is optimal for the control problem, there exist absolutely continuous adjoint variables satisfying the system of fractional differential equations:
-${}^{C}D_{t}^{\alpha}\lambda_{S}$ $=-\frac{\partial H}{\partial S},$ ${}^{C}D_{t}^{\alpha}\lambda_{V}$ $=-\frac{\partial H}{\partial V},$ ${}^{C}D_{t}^{\alpha}\lambda_{E}$ $=-\frac{\partial H}{\partial E},$
-${}^{C}D_{t}^{\alpha}\lambda_{I_{s}}$ $=-\frac{\partial H}{\partial I_{s}},$ ${}^{C}D_{t}^{\alpha}\lambda_{I_{a}}$ $=-\frac{\partial H}{\partial I_{a}},$ ${}^{C}D_{t}^{\alpha}\lambda_{H}$ $=-\frac{\partial H}{\partial H},$
-${}^{C}D_{t}^{\alpha}\lambda_{D}$ $=-\frac{\partial H}{\partial D},$ ${}^{C}D_{t}^{\alpha}\lambda_{R}$ $=-\frac{\partial H}{\partial R},$
-The optimal control $\mathbf{u}^{*}(t)$ is described by
-$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}\mathcal{H}(t,\mathbf{x}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{a.e. on }[0,T],$
-supplemented with the transversality condition
-$\bm{\lambda}(T)=\mathbf{0},$
-with $\bm{\lambda}=(\lambda_{S},\lambda_{V},\lambda_{E},\lambda_{I_{s}},\lambda_{I_{a}},\lambda_{H},\lambda_{D},\lambda_{R})^{\top}$.
+Let $(\mathbf{X}^{*},\mathbf{u}^{*})$ be an optimal solution for the control problem defined by the objective functional (49) subject to the fractional-order system (48).
+Then there exist absolutely continuous adjoint variables $\bm{\lambda}=(\lambda_{S},\lambda_{V},\lambda_{E},\lambda_{I_{s}},\lambda_{I_{a}},\lambda_{H},\lambda_{D},\lambda_{R})^{\top}$ satisfying the fractional adjoint equations
+${}^{C}D_{t}^{\alpha}\lambda_{S}$ $=-\frac{\partial\mathcal{H}}{\partial S},\quad{}^{C}D_{t}^{\alpha}\lambda_{V}=-\frac{\partial\mathcal{H}}{\partial V},\quad{}^{C}D_{t}^{\alpha}\lambda_{E}=-\frac{\partial\mathcal{H}}{\partial E},$ (52)
+${}^{C}D_{t}^{\alpha}\lambda_{I_{s}}$ $=-\frac{\partial\mathcal{H}}{\partial I_{s}},\quad{}^{C}D_{t}^{\alpha}\lambda_{I_{a}}=-\frac{\partial\mathcal{H}}{\partial I_{a}},\quad{}^{C}D_{t}^{\alpha}\lambda_{H}=-\frac{\partial\mathcal{H}}{\partial H},$ (53)
+${}^{C}D_{t}^{\alpha}\lambda_{D}$ $=-\frac{\partial\mathcal{H}}{\partial D},\quad{}^{C}D_{t}^{\alpha}\lambda_{R}=-\frac{\partial\mathcal{H}}{\partial R},$ (54)
+where the Hamiltonian $\mathcal{H}$ is defined as
+$\mathcal{H}(t,\mathbf{X},\mathbf{u},\bm{\lambda})=\mathcal{L}(t,\mathbf{X},\mathbf{u})+\bm{\lambda}^{\top}\mathbf{f}(t,\mathbf{X},\mathbf{u}),$
+with $\mathcal{L}(t,\mathbf{X},\mathbf{u})=\sum_{i=1}^{4}A_{i}X_{i}+\frac{1}{2}\sum_{j=1}^{4}B_{j}u_{j}^{2}$ and $\mathbf{f}$ denoting the right-hand side of (48).
+The optimal control $\mathbf{u}^{*}(t)$ satisfies the minimization condition
+$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}\mathcal{H}(t,\mathbf{X}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{almost everywhere on }[0,T],$ (55)
+with the transversality condition
+$\bm{\lambda}(T)=\mathbf{0}.$ (56)
 Proof.
 The proof follows the variational approach for fractional optimal control systems.
-Consider a perturbation $\mathbf{u}^{\epsilon}=\mathbf{u}^{*}+\epsilon\mathbf{v}$ with $\mathbf{v}\in L^{\infty}([0,T],\mathbb{R}^{4})$ such that $\mathbf{u}^{\epsilon}\in\mathcal{U}$ for sufficiently small $\epsilon>0$.
+Let $(\mathbf{X}^{*},\mathbf{u}^{*})$ be an optimal pair.
+Consider a perturbation $\mathbf{u}^{\epsilon}=\mathbf{u}^{*}+\epsilon\mathbf{v}$, where $\mathbf{v}\in L^{\infty}([0,T],\mathbb{R}^{4})$ satisfies $\mathbf{u}^{\epsilon}\in\mathcal{U}$ for sufficiently small $\epsilon>0$.
 Let $\mathbf{X}^{\epsilon}$ denote the corresponding state trajectory.
-The first variation of the objective functional is given by:
+The first variation of the objective functional is given by
 $\delta J=\lim_{\epsilon\to 0}\frac{J(\mathbf{u}^{\epsilon})-J(\mathbf{u}^{*})}{\epsilon}=\int_{0}^{T}\left[\frac{\partial\mathcal{L}}{\partial\mathbf{X}}\cdot\delta\mathbf{X}+\frac{\partial\mathcal{L}}{\partial\mathbf{u}}\cdot\mathbf{v}\right]dt,$
-where $\delta\mathbf{X}$ satisfies the linearized state equation derived from the controlled system.
-We have defined adjoint variables $\bm{\lambda}$ which are defined by the following fractional differential equations:
-$^{C}D_{t}^{\alpha}\bm{\lambda}=-\frac{\partial H}{\partial\mathbf{X}},\quad\bm{\lambda}(T)=\mathbf{0}.$
-Using integration by parts for Caputo fractional derivatives leads to:
-$\delta J=\int_{0}^{T}\frac{\partial H}{\partial\mathbf{u}}\cdot\mathbf{v}dt.$
-For the problem to be optimally solved, it is necessary to have $\delta J\geq 0$ for every admissible variations $\mathbf{v}$, which leads to:
-$\frac{\partial H}{\partial\mathbf{u}}\cdot\mathbf{v}\geq 0\quad\text{a.e. on }[0,T].$
-From this inequality, we derive the pointwise minimization condition. Consolidating the minimization condition, we have:
-$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}H(t,\mathbf{X}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{a.e.}$
-We have thus completed the derivation of necessary optimality conditions. $\square$
-6.4 Optimal Control Characterization
+where $\delta\mathbf{X}$ satisfies the linearized state equation derived from (48).
+Define the adjoint variables $\bm{\lambda}$ through the fractional boundary value problem
+$^{C}D_{t}^{\alpha}\bm{\lambda}=-\frac{\partial\mathcal{H}}{\partial\mathbf{X}},\quad\bm{\lambda}(T)=\mathbf{0}.$ (57)
+Applying the fractional integration by parts formula for Caputo derivatives to the term involving $\delta\mathbf{X}$ yields
+$\int_{0}^{T}\frac{\partial\mathcal{L}}{\partial\mathbf{X}}\cdot\delta\mathbf{X}\,dt=\int_{0}^{T}\bm{\lambda}\cdot\left[\frac{d}{dt}\delta\mathbf{X}\right]\,dt=-\int_{0}^{T}\left[{}^{C}D_{t}^{\alpha}\bm{\lambda}\right]\cdot\delta\mathbf{X}\,dt.$
+Substituting the adjoint equation (57) and simplifying gives
+$\delta J=\int_{0}^{T}\left[\frac{\partial\mathcal{H}}{\partial\mathbf{u}}\right]\cdot\mathbf{v}\,dt.$
+For optimality, we require $\delta J\geq 0$ for all admissible variations $\mathbf{v}$, which implies the pointwise condition
+$\frac{\partial\mathcal{H}}{\partial\mathbf{u}}\cdot\mathbf{v}\geq 0\quad\text{almost everywhere on }[0,T].$
+This inequality leads to the minimization condition
+$\mathbf{u}^{*}(t)=\arg\min_{\mathbf{u}\in\mathcal{U}}\mathcal{H}(t,\mathbf{X}^{*}(t),\mathbf{u},\bm{\lambda}(t))\quad\text{a.e. on }[0,T],$
+completing the derivation of the necessary optimality conditions. $\square$
 Theorem 5.3
-The optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that minimize the Hamiltonian are characterized by:
-$u_{1}^{*}(t)$ $=\min\left\{u_{1}^{\max},\max\left\{0,\frac{[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)]\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{B_{1}N}\right\}\right\},$
-$u_{2}^{*}(t)$ $=\min\left\{u_{2}^{\max},\max\left\{0,\frac{(\lambda_{S}-\lambda_{V})S}{B_{2}}\right\}\right\},$
-$u_{3}^{*}(t)$ $=\min\left\{u_{3}^{\max},\max\left\{0,\frac{(\lambda_{I_{s}}-\lambda_{H})I_{s}}{B_{3}}\right\}\right\},$
-$u_{4}^{*}(t)$ $=\min\left\{u_{4}^{\max},\max\left\{0,\frac{\lambda_{D}D}{B_{4}}\right\}\right\},$
-where all state and adjoint variables follow the optimal trajectory.
+The optimal controls $u_{1}^{*},u_{2}^{*},u_{3}^{*},u_{4}^{*}$ that minimize the Hamiltonian $\mathcal{H}$ are given by
+$u_{1}^{*}(t)$ $=\min\left\{u_{1}^{\max},\max\left\{0,\frac{[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)]\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{B_{1}N}\right\}\right\},$ (58)
+$u_{2}^{*}(t)$ $=\min\left\{u_{2}^{\max},\max\left\{0,\frac{(\lambda_{S}-\lambda_{V})S}{B_{2}}\right\}\right\},$ (59)
+$u_{3}^{*}(t)$ $=\min\left\{u_{3}^{\max},\max\left\{0,\frac{(\lambda_{I_{s}}-\lambda_{H})I_{s}}{B_{3}}\right\}\right\},$ (60)
+$u_{4}^{*}(t)$ $=\min\left\{u_{4}^{\max},\max\left\{0,\frac{\lambda_{D}D}{B_{4}}\right\}\right\},$ (61)
+where all state and adjoint variables follow their optimal trajectories.
 Proof.
-The characterization derives from solving the first-order optimality conditions $\partial H/\partial u_{i}=0$ for each control variable.
-For personal protection control $u_{1}^{*}$, the optimality condition yields:
-$\frac{\partial H}{\partial u_{1}}=B_{1}u_{1}+\left[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-(S+(1-\varepsilon)V)\lambda_{E}\right]\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}=0.$
-Solving algebraically provides the unconstrained optimum:
+The optimal controls are derived by solving the first-order optimality conditions $\partial\mathcal{H}/\partial u_{i}=0$ for each control variable, then projecting the solutions onto the admissible interval $0,u_{i}^{\max}]$.
+For personal protection control $u_{1}$, the optimality condition provides
+$\frac{\partial\mathcal{H}}{\partial u_{1}}=B_{1}u_{1}+\left[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)\right]\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N}=0.$
+Solving for $u_{1}$ gives the unconstrained optimum
 $u_{1}=\frac{[\lambda_{S}S+(1-\varepsilon)\lambda_{V}V-\lambda_{E}(S+(1-\varepsilon)V)]\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{B_{1}N}.$
-The last description arises from the projection onto the last interval $0,u_{1}^{\max}]$. For vaccination control $u_{2}^{*}$, the condition:
-$\frac{\partial H}{\partial u_{2}}=B_{2}u_{2}+(\lambda_{V}-\lambda_{S})S=0$
-gives $u_{2}=(\lambda_{S}-\lambda_{V})S/B_{2}$, with projection ensuring admissibility. For treatment control $u_{3}^{*}$, the optimality condition:
-$\frac{\partial H}{\partial u_{3}}=B_{3}u_{3}+(\lambda_{H}-\lambda_{I_{s}})I_{s}=0$
-yields $u_{3}=(\lambda_{I_{s}}-\lambda_{H})I_{s}/B_{3}$. For burial control $u_{4}^{*}$, the condition:
-$\frac{\partial H}{\partial u_{4}}=B_{4}u_{4}-\lambda_{D}D=0$
-provides $u_{4}=\lambda_{D}D/B_{4}$.
-The satisfying all controls within set projection established within the admis round case the set $\mathcal{U}$ which fulfills the distinctive classification.
+Projecting this expression onto the admissible interval $0,u_{1}^{\max}]$ yields the optimal control (58).
+For vaccination control $u_{2}$, the optimality condition is
+$\frac{\partial\mathcal{H}}{\partial u_{2}}=B_{2}u_{2}+(\lambda_{V}-\lambda_{S})S=0.$
+Solving gives $u_{2}=(\lambda_{S}-\lambda_{V})S/B_{2}$, and after projection we obtain (59).
+For treatment control $u_{3}$, we have
+$\frac{\partial\mathcal{H}}{\partial u_{3}}=B_{3}u_{3}+(\lambda_{H}-\lambda_{I_{s}})I_{s}=0,$
+which yields $u_{3}=(\lambda_{I_{s}}-\lambda_{H})I_{s}/B_{3}$. Projection gives (60).
+For safe burial control $u_{4}$, the condition
+$\frac{\partial\mathcal{H}}{\partial u_{4}}=B_{4}u_{4}-\lambda_{D}D=0$
+gives $u_{4}=\lambda_{D}D/B_{4}$. Projection onto $0,u_{4}^{\max}]$ yields (61).
+Each optimal control is thus characterized by projecting the solution of the first-order optimality condition onto the admissible control set $\mathcal{U}$, ensuring that all constraints $0\leq u_{i}(t)\leq u_{i}^{\max}$ are satisfied.
 $\square$
-6.5 Adjoint System Specification
 Theorem 5.4
-The fractional differential equations outlined below are satisfied by the adjoint variables:
+The adjoint variables satisfy the following system of fractional differential equations
 ${}^{C}D_{t}^{\alpha}\lambda_{S}$ $=(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(V+E+I_{s}+I_{a}+H+R)}{N^{2}}\right]$
@@
 $\quad+(\lambda_{V}-\lambda_{E})(1-\varepsilon)(1-u_{1})\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right].$ (69)
-Proof.  The calculation of the partial derivative of the Hamiltonian yields the equations of the adjoint each corresponding to the state variable of consideration. For $\lambda_{S}$, we have:
+Proof.
+The adjoint equations are derived by computing the partial derivatives of the Hamiltonian $\mathcal{H}$ with respect to each state variable, then applying the relationship $^{C}D_{t}^{\alpha}\lambda_{i}=-\partial\mathcal{H}/\partial x_{i}$ from Pontryagin’s maximum principle.
+For $\lambda_{S}$, we compute
 $\frac{\partial\mathcal{H}}{\partial S}$ $=(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(V+E+I_{s}+I_{a}+H+R)}{N^{2}}\right]$
-$+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
-$+\lambda_{S}(\mu+u_{2})-\lambda_{V}u_{2}.$
-Thus, (28) follows from $^{C}D_{t}^{\alpha}\lambda_{S}=-\partial\mathcal{H}/\partial S$. For $\lambda_{V}$, the derivative gives:
+$\quad+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta V(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
+$\quad+\lambda_{S}(\mu+u_{2})-\lambda_{V}u_{2}.$
+Equation (62) follows from $^{C}D_{t}^{\alpha}\lambda_{S}=-\partial\mathcal{H}/\partial S$.
+For $\lambda_{V}$, we have
 $\frac{\partial\mathcal{H}}{\partial V}$ $=-\lambda_{S}\omega+\lambda_{V}(\omega+\mu)$
-$+(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
-$+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(S+E+I_{s}+I_{a}+H+R)}{N^{2}}\right],$
-which gives (29) under $^{C}D_{t}^{\alpha}\lambda_{V}=-\partial\mathcal{H}/\partial V$.
-The remaining adjoint equations (30)–(35) can be derived in a similar manner by calculating the respective partial derivatives $\partial\mathcal{H}/\partial x_{i}$ and using the fundamental relationship $^{C}D_{t}^{\alpha}\lambda_{i}=-\partial\mathcal{H}/\partial x_{i}$.
-The arrangement continues to embody the epidemiological relations where the transmission mechanisms are influenced by the cosate differentials and the passage rates determined by the compartments.
+$\quad+(1-u_{1})(\lambda_{S}-\lambda_{E})\left[\frac{\beta S(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N^{2}}\right]$
+$\quad+(1-u_{1})(\lambda_{V}-\lambda_{E})(1-\varepsilon)\left[\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)(S+E+I_{s}+I_{a}+H+R)}{N^{2}}\right],$
+which yields (63) via $^{C}D_{t}^{\alpha}\lambda_{V}=-\partial\mathcal{H}/\partial V$.
+The remaining adjoint equations (64)–(69) are obtained similarly by computing the corresponding partial derivatives $\partial\mathcal{H}/\partial x_{i}$ and applying the fundamental adjoint relationship $^{C}D_{t}^{\alpha}\lambda_{i}=-\partial\mathcal{H}/\partial x_{i}$.
+Each derivative accounts for both direct effects of the state variable on the Hamiltonian and indirect effects through the force of infection $\lambda(t)$.
 $\square$
 Remark 5.1
-Each part of the adjoint system can be seen recursively decomposed in the sequence structured by the equations $\lambda_{R}$, $\lambda_{H}$, and $\lambda_{D}$, which become independent once the other adjoint variables are fixed.
-This characteristic feature of the system corresponds to the compartmentalized configuration of the epidemiological dynamics and facilitates the computation of an efficient numerical solution.
-6.6 Numerical Results of Optimal Control
-Figure 7: Optimal Control Analysis : (a) Algorithm Convergence, (b) Control Path, (c) Cost Control Effectiveness, (d) Control of Epidemics and Containment
-Using Python, NumPy, and SciPy, the faster forward-backward sweep algorithm achieved convergence, exhibiting epidemic containment (peak: 170.0, final: 0.0) in 50 iterations within 30 seconds, and at a cost of 8750\$ Of the burial controls, which were the most cost-effective at 4\$/case, treatment with a 92% reduction in transmission, sequentially distilled intervention timing.
-7 Efficiency Evaluation of Optimal Interventions
-The numerical simulation of the fractional-order Ebola system employs the RK45 method.
-This method features adaptive step-size control which allows for a proper trade-off between precision and computational efficiency.
-Algorithm 1 Runge-Kutta-Fehlberg (RKF45)
-[…]
-[…]
-3:Initialize: $t\leftarrow t_{0}$, $\mathbf{y}\leftarrow\mathbf{y}_{0}$
-4:Set parameters: $h_{min}\leftarrow 10^{-8}$, $h_{max}\leftarrow 0.1$, $h\leftarrow 0.01$
+The adjoint system presents a recursive structure, the equations for $\lambda_{R}$, $\lambda_{H}$, and $\lambda_{D}$ become independent once the other adjoint variables are determined.
+This structure reflects the compartmental organization of the epidemiological dynamics and facilitates efficient numerical solution of the optimality system.
+5.1 Cost-Effectiveness of Optimal Control
+Figure 4: Optimal Control Analysis (a) Algorithm Convergence, (b) Control Path, (c) Cost Control Effectiveness, (d) Control of Epidemics and Containment
+We performed numerical simulations in Python using the NumPy and SciPy libraries.
+After 50 iterations, the quick forward-backward sweep method achieved the ideal solution in 30 seconds.
+The optimal control strategy achieved epidemic containment, with the infected population peaking at 170 individuals before declining to zero, at a total intervention cost of 8750\$.
+Safe burial was the most cost-saving intervention, having a unit cost of 4\$ per case, and treatment resulted into a 92% proportion of reduction in transmission.
+The algorithm also identified the ideal time when each intervention should be put in place.
+5.2 Numerical Assessment of intervention Strategies.
+The RungeKutta45 scheme is used to perform the numerical analysis of the fractional-order Ebola system.
+This method uses the adaptive step-size control in order to balance the computational cost with the number of numerical accuracy needed to derive the best intervention strategy.
+Algorithm 1 Runge-Kutta-Fehlberg (RKF45) Method for Adaptive Numerical Integration
+1:Time interval $t_{0},T]$, initial state $\mathbf{y}_{0}$, system parameters $\mathbf{p}$, relative tolerance $\epsilon_{\text{rel}}=10^{-6}$, absolute tolerance $\epsilon_{\text{abs}}=10^{-8}$
+2:High-precision numerical solution $\mathbf{y}(t)$ for $t\in[t_{0},T]$
+3:Initialize $t\leftarrow t_{0}$, $\mathbf{y}\leftarrow\mathbf{y}_{0}$
+4:Set step size bounds $h_{\min}\leftarrow 10^{-8}$, $h_{\max}\leftarrow 0.1$, initial $h\leftarrow 0.01$
 5:while $t<T$ do
-6:  Compute RK coefficients $\mathbf{k}_{1}$ through $\mathbf{k}_{6}$
-7:  Calculate 4th/5th order solutions $\mathbf{y}_{4}$, $\mathbf{y}_{5}$
-8:  Estimate error $\delta=\|\mathbf{y}_{5}-\mathbf{y}_{4}\|$
-9:  Compute optimal step $h_{opt}=0.9h(\epsilon_{rel}/\delta)^{0.2}$
-10:  if $\delta\leq\epsilon_{rel}$ then
-11:   Accept step: $t\leftarrow t+h$, $\mathbf{y}\leftarrow\mathbf{y}_{5}$
-12:   Update $h\leftarrow\min(h_{max},\max(h_{min},h_{opt}))$
+6:  Compute Runge-Kutta coefficients $\mathbf{k}_{1}$ through $\mathbf{k}_{6}$ via standard RKF45 formulas
+7:  Calculate 4th-order ($\mathbf{y}_{4}$) and 5th-order ($\mathbf{y}_{5}$) solutions
+8:  Estimate local truncation error $\delta\leftarrow\|\mathbf{y}_{5}-\mathbf{y}_{4}\|$
+9:  Compute optimal step size $h_{\text{opt}}\leftarrow 0.9h\left(\dfrac{\epsilon_{\text{rel}}\|\mathbf{y}_{5}\|+\epsilon_{\text{abs}}}{\delta}\right)^{0.2}$
+10:  if $\delta\leq\epsilon_{\text{rel}}\|\mathbf{y}_{5}\|+\epsilon_{\text{abs}}$ then
+11:   Accept step $t\leftarrow t+h$, $\mathbf{y}\leftarrow\mathbf{y}_{5}$
+12:   Update step $h\leftarrow\min(h_{\max},\max(h_{\min},h_{\text{opt}}))$
 13:  else
-14:   Reject step: $h\leftarrow\max(h_{min},h_{opt})$
+14:   Reject step $h\leftarrow\max(h_{\min},h_{\text{opt}})$
 15:  end if
 16:end while
-17:return Interpolated solution
+17:return Interpolated solution with 4th-order accuracy and 5th-order error control
 The algorithm maintains 4th-order accuracy with 5th-order error control, employing adaptive step-size selection based on local truncation error estimates.
@@
 Figure 5: Dynamics in compartments under Treatment Intervention Strategy ($u_{3}\neq 0$)
-Figure 9: Foundational epidemiological principles regarding the impact of the treatment intervention
-Figure 10: Effectiveness assessment of treatment strategy
 Figure 6: compartmatic dynamics with safe burial intervention $u_{4}\neq 0$
-Figure 12: Comparison of key metrics for safe burial intervention.
-Figure 13: Analysis of Safeguard Burial Approach Efficiency
 Figure 7: Dynamics of compartments for the strategy of instantaneous vaccination and treatment ($u_{2}\neq 0,u_{3}\neq 0$)
-Figure 15: Key Metrics for the combined treatment and vaccination
-Figure 16: Effectiveness assessment of combined vaccination and treatment
 Figure 8: Compartmental dynamics for combined treatment and safe burial strategy ($u_{3}\neq 0,u_{4}\neq 0$)
-Figure 18: Key metrics comparison for treatment-safe burial combination
-Figure 19: Effectiveness assessment of treatment and safe burial strategy
 Figure 9: Compartmental dynamics for three-intervention combination ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0$)
-Figure 21: Key metrics for personal protection, vaccination, and treatment combination
-Figure 22: Effectiveness assessment of three-intervention strategy
 Figure 10: Compartmental dynamics for vaccination, treatment, and safe burial combination ($u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)
-Figure 24: Key metrics comparison for clinical and preventive combination
-Figure 25: Effectiveness assessment of vaccination, treatment, and safe burial strategy
 Figure 11: Compartmental dynamics under comprehensive intervention strategy ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)
-Figure 27: Key metrics for comprehensive intervention approach
-Figure 28: Effectiveness assessment of comprehensive strategy
-7.2 Discussion
 Evaluating strategies for controlling pandemics has unequally advanced in mortality reduction and has repeatedly failed in controlling transmission.
@@
 There were significant interactions when different types of interventions were used together.
-From figures $16$ vaccination-treatment combination ($u_{2}\neq 0,u_{3}\neq 0$) achieved a 60.4% reduction in mortality which is the same as treatment done alone, therefore, illustrating that vaccination is not additive.
+From figures 7 vaccination-treatment combination ($u_{2}\neq 0,u_{3}\neq 0$) achieved a 60.4% reduction in mortality which is the same as treatment done alone, therefore, illustrating that vaccination is not additive.
 In contrast, treatment with safe burial ($u_{3}\neq 0,u_{4}\neq 0$) and burial leads to 86.5% mortality reduction, which is a synergistic over additive effect that far exceeds the efficacy of the single interventions.
@@
 Combining interventions produced the same results.
-the combined three type intervention ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0$) only reached 60.4% reduction, while the rest ($u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$) and the comprehensive method ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)) both reached the maximum reduction of 86.5% gained with the treatment and safe burial pair.
-It is the combination of these results, documented in Figures 2225 and  $28$ that confirms treatment and safe burial dominate the loss-of-life mitigation, while in these circumstances, personal protection and vaccination are less effective.
+The combined three type intervention in figure 9 ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0$) only reached 60.4% reduction, while the rest ($u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$) and the comprehensive method ($u_{1}\neq 0,u_{2}\neq 0,u_{3}\neq 0,u_{4}\neq 0$)) both reached the maximum reduction of 86.5% gained with the treatment and safe burial pair.
+It is the combination of these results, documented in figures 10 and  11 that confirms treatment and safe burial dominate the loss-of-life mitigation, while in these circumstances, personal protection and vaccination are less effective.
 In the context of EVD, personal protection does not mean social distancing; rather, it refers to the use of barrier techniques for those who come into direct interaction with infected persons or dead bodies.
@@
 This underscores the importance of post-exposure treatment combined with the clinical burial of the dead to reduce casualties.
-8 Intelligent Disease Modeling – DINN of EVD
-This segment presents the mathematical concepts supporting the architecture of the Disease-Informed Neural Network (DINN) as per the formulae of the fractional-order Ebola model that was discussed.
-(2).
-In the case of the DINN architecture, the fractional differential equations constrain the architectural configuration of the neural-embedded network, enabling the network to incorporate the differential equations of the epidemic system defined by the fractional-order differential equations, thus capturing the system’s physically plausible dynamics.
-The governing system of fractional-order differential equations is expressed in the compact form:
-$^{C}D_{t}^{\alpha}Y(t)+\mathcal{G}(Y(t;\chi))=0,\quad t\in[t_{0},T],$ (36)
-The state Y(t) = [S(t), V(t), E(t), $I_{s}(t)$, $I_{a}(t)$, H(t), D(t), R(t)] along with the vector concept that every compartment population tends to zero.
-Let $\mathcal{G}(Y)=[g_{1}(Y),\ldots,g_{8}(Y)]$, where each function $g_{i}(Y)$ captures the heterogeneously nonlinear connections that inflow and outflow through its compartment.
-$\chi\in\mathbb{R}^{p}$ the parameters in the system which determines epidemiological corresponding with the degree/level of the interventions effectiveness and the intervention measures being undertaken.
-The neural network is approximated as centralized with the equations $NN_{w,b}(t):\mathbb{R}\longrightarrow\mathbb{R}^{8}$ which is initial feasible $Y(t)$ solution because of the abstract layer termed trainable beings the connected parameters the neural network in the $w$ and $b$ configurations as edges as nodes as the two fundamental elements in the neural network.
-The subsequent component of the function is termed the data portion is given by
-$\mathcal{L}_{data}=\frac{1}{m}\sum_{s=1}^{m}\|NN_{w,b}(t_{s})-Y_{s}\|^{2}.$ (37)
-The physical consistency is enforced through the residual term, which measures the discrepancy between the neural network output and the governing fractional dynamics:
-$\mathcal{R}(NN_{w,b},t;\chi)={}^{C}D_{t}^{\alpha}NN_{w,b}+\mathcal{G}(NN_{w,b}(t)).$ (38)
-The corresponding residual loss is computed as:
-$MSE_{w,b,\chi}^{\mathcal{G}}=\frac{1}{N}\sum_{i=1}^{N}\|\mathcal{R}(NN_{w,b},t_{i};\chi)\|^{2},$ (39)
-The use of collocation points $t_{1},t_{2},\ldots,t_{N}$ discretizes the temporal domain.
-The entirety of this optimization problem simultaneously evaluates the parameters defining the neural network and the system by means of the composite objective:
-$(w^{*},b^{*},\chi^{*})=\arg\min_{w,b,\chi}\left(\mathcal{L}_{data}+\lambda MSE_{w,b,\chi}^{\mathcal{G}}\right),$ (40)
-where $\lambda$ balances the contribution of data compliance and physical consistency.
-This formulation guarantees that the acquired solution serves the purpose of not only fitting the data but also complying with the core mechanistic precepts embedded in the fractional-order Ebola transmission model.
-8.1 ML-EVD model
-This subsection describes the concrete realization of the Disease-Informed Neural Network (DINN) framework design for the fractional-order Ebola epidemic model.
-The advanced physical consistency is maintained through an elaborate residual term that contains the entire system dynamics:
-$\mathcal{R}(NN_{w,b},t;\chi)=\begin{pmatrix}{}^{C}D_{t}^{\alpha}S-\left[\Lambda-\lambda S+\omega V-(v+\mu)S\right]\\ {}^{C}D_{t}^{\alpha}V-\left[vS-(1-\varepsilon)\lambda V-(\mu+\omega)V\right]\\ {}^{C}D_{t}^{\alpha}E-\left[\lambda(S+(1-\varepsilon)V)-(\mu+\sigma)E\right]\\ {}^{C}D_{t}^{\alpha}I_{s}-\left[p\sigma E-(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right]\\ {}^{C}D_{t}^{\alpha}I_{a}-\left[\sigma E(1-p)-(\gamma_{a}+\mu)I_{a}\right]\\ {}^{C}D_{t}^{\alpha}H-\left[h_{s}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H\right]\\ {}^{C}D_{t}^{\alpha}D-\left[\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D\right]\\ {}^{C}D_{t}^{\alpha}R-\left[\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R\right]\end{pmatrix}$, (41)
-where the time-dependent force of infection is defined as:
-$\lambda(t)=\frac{c\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N},$ (42)
-Let $N$ denotes the total living population such that $N=S+V+E+I_{s}+I_{a}+H+R$.
-The residual loss function can be split into compartment-specific components, each measuring the gap between the neural network’s fractional derivatives and the corresponding dynamical equations:
-$MSE_{residual}$ $=\sum_{X\in\mathcal{C}}MSE_{X},$ (43)
-$MSE_{S}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}S-\Lambda+\lambda S-\omega V+(v+\mu)S\right|^{2},$ (44)
-$MSE_{V}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}V-vS+(1-\varepsilon)\lambda V+(\mu+\omega)V\right|^{2},$ (45)
-$MSE_{E}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}E-\lambda(S+(1-\varepsilon)V)+(\mu+\sigma)E\right|^{2},$ (46)
-$MSE_{I_{s}}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}I_{s}-p\sigma E+(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right|^{2},$ (47)
-$MSE_{I_{a}}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}I_{a}-\sigma E(1-p)+(\gamma_{a}+\mu)I_{a}\right|^{2},$ (48)
-$MSE_{H}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}H-h_{s}I_{s}+(\gamma_{h}+\delta_{h}+\mu)H\right|^{2},$ (49)
-$MSE_{D}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}D-\delta_{s}I_{s}-\delta_{h}H+\mu_{d}D\right|^{2},$ (50)
-$MSE_{R}$ $=\frac{1}{q}\sum_{i=1}^{q}\left|{}^{C}D_{t}^{\alpha}R-\gamma_{s}I_{s}-\gamma_{a}I_{a}-\gamma_{h}H+\mu R\right|^{2},$ (51)
-where $\mathcal{C}=\{S,V,E,I_{s},I_{a},H,D,R\}$ indicates all the epidemiological compartments.
-Consistency with the observational measurements is guaranteed with the components set within the data adherence phase
-$MSE_{data}$ $=\sum_{X\in\mathcal{C}}MSE_{X}^{data},$ (52)
-$MSE_{X}^{data}$ $=\frac{1}{s}\sum_{i=1}^{s}|X(t_{i})-X_{i}^{o}|^{2},\quad X\in\mathcal{C},$ (53)
-Let $X_{i}^{o}$ be the observed value of compartment $X$ at time $t_{i}$.
-The entire optimization task finds the values of both the neural network parameters and the parameters of the epidemiological model simultaneously:
-$(w^{*},b^{*},\chi^{*})=\arg\min_{w,b,\chi}\left(MSE_{data}+MSE_{residual}\right),$ (54)
-The collection of parameters $\chi=\{\Lambda,\mu,c(t),\beta,\eta_{a},\eta_{d},v(t),\varepsilon,\omega,\sigma,p,h_{s},\gamma_{s},\gamma_{a},\gamma_{h},\delta_{s},\delta_{h},\mu_{d},\alpha\}$ includes all constants related to the flow of epidemiological modeling to describe the transmission, progression of the disease and the impact of different control measures.
-Using this approach ensures that the learned solution integrates the observed data with the mechanistic underpinning of the Ebola transmission model.
+6 Disease-Informed Neural Network (DINN) for Ebola Modeling
+This section introduces the Disease-Informed Neural Network (DINN), a physics-informed neural network framework that incorporates the fractional-order Ebola model (2) as physical constraints to ensure epidemiological consistency in predictions.
+The fractional-order system governing Ebola dynamics is expressed as
+$^{C}D_{t}^{\alpha}\mathbf{Y}(t)+\mathcal{G}(\mathbf{Y}(t);\bm{\chi})=0,\quad t\in[t_{0},T],$ (70)
+where $\mathbf{Y}(t)=[S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t)]^{\top}$ denotes the state vector and $\bm{\chi}$ represents the epidemiological parameters.
+A neural network $\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t):\mathbb{R}\to\mathbb{R}^{8}$ approximates the solution $\mathbf{Y}(t)$, where $\mathbf{w}$ and $\mathbf{b}$ are the network weights and biases.
+The training process minimizes two loss components: a data fidelity term and a physics consistency term
+$\mathcal{L}_{\text{data}}$ $=\frac{1}{m}\sum_{s=1}^{m}\|\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t_{s})-\mathbf{Y}_{s}\|^{2},$
+$\mathcal{L}_{\text{physics}}$ $=\frac{1}{N}\sum_{i=1}^{N}\|{}^{C}D_{t}^{\alpha}\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t_{i})+\mathcal{G}(\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t_{i});\bm{\chi})\|^{2}.$
+Here, $\mathcal{L}_{\text{data}}$ penalizes deviations from observed data, while $\mathcal{L}_{\text{physics}}$ enforces the governing equations of the fractional-order model.
+The DINN framework solves the composite optimization problem
+$(\mathbf{w}^{*},\mathbf{b}^{*},\bm{\chi}^{*})=\arg\min_{\mathbf{w},\mathbf{b},\bm{\chi}}\left(\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}\right),$ (71)
+where $\lambda>0$ is used to balance between data fidelity and physical constraints.
+The suggested methodology will both optimize the weights of the neural network and estimate the epidemiological parameters, thus giving both a solution that is both in line with the empirical data and epidemiologically plausible.
+6.1 DINN Architecture for the Ebola Model
+In this subsection, the architecture of Disease-Informed Neural Network (DINN) that is used in the fractional-order Ebola model is described in detail.
+The network applies physical consistency directly with its incorporation of the model dynamics into the loss function.
+The remaining vector that measures the difference between the outcome of the neural network and underlying epidemiological dynamics is denoted by the residual as
+$\mathcal{R}(\mathcal{NN}_{\mathbf{w},\mathbf{b}},t;\bm{\chi})=\begin{pmatrix}{}^{C}D_{t}^{\alpha}S-\left[\Lambda-\lambda S+\omega V-(v+\mu)S\right]\\ {}^{C}D_{t}^{\alpha}V-\left[vS-(1-\varepsilon)\lambda V-(\mu+\omega)V\right]\\ {}^{C}D_{t}^{\alpha}E-\left[\lambda(S+(1-\varepsilon)V)-(\mu+\sigma)E\right]\\ {}^{C}D_{t}^{\alpha}I_{s}-\left[p\sigma E-(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\right]\\ {}^{C}D_{t}^{\alpha}I_{a}-\left[\sigma E(1-p)-(\gamma_{a}+\mu)I_{a}\right]\\ {}^{C}D_{t}^{\alpha}H-\left[h_{s}I_{s}-(\gamma_{h}+\delta_{h}+\mu)H\right]\\ {}^{C}D_{t}^{\alpha}D-\left[\delta_{s}I_{s}+\delta_{h}H-\mu_{d}D\right]\\ {}^{C}D_{t}^{\alpha}R-\left[\gamma_{s}I_{s}+\gamma_{a}I_{a}+\gamma_{h}H-\mu R\right]\end{pmatrix}$, (72)
+where the time-dependent force of infection is given by
+$\lambda(t)=\frac{\beta(I_{s}+\eta_{a}I_{a}+\eta_{d}D)}{N},$
+and the total living population is $N=S+V+E+I_{s}+I_{a}+H+R$.
+The physics-informed loss function is decomposed into compartment-specific mean squared errors
+$\mathcal{L}_{\text{residual}}$ $=\sum_{X\in\mathcal{C}}\mathcal{L}_{X},$
+$\mathcal{L}_{S}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}S-\Lambda+\lambda S-\omega V+(v+\mu)S\bigr|^{2},$
+$\mathcal{L}_{V}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}V-vS+(1-\varepsilon)\lambda V+(\mu+\omega)V\bigr|^{2},$
+$\mathcal{L}_{E}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}E-\lambda(S+(1-\varepsilon)V)+(\mu+\sigma)E\bigr|^{2},$
+$\mathcal{L}_{I_{s}}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}I_{s}-p\sigma E+(\gamma_{s}+\delta_{s}+h_{s}+\mu)I_{s}\bigr|^{2},$
+$\mathcal{L}_{I_{a}}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}I_{a}-\sigma E(1-p)+(\gamma_{a}+\mu)I_{a}\bigr|^{2},$
+$\mathcal{L}_{H}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}H-h_{s}I_{s}+(\gamma_{h}+\delta_{h}+\mu)H\bigr|^{2},$
+$\mathcal{L}_{D}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}D-\delta_{s}I_{s}-\delta_{h}H+\mu_{d}D\bigr|^{2},$
+$\mathcal{L}_{R}$ $=\frac{1}{q}\sum_{i=1}^{q}\bigl|{}^{C}D_{t}^{\alpha}R-\gamma_{s}I_{s}-\gamma_{a}I_{a}-\gamma_{h}H+\mu R\bigr|^{2},$
+where $\mathcal{C}=\{S,V,E,I_{s},I_{a},H,D,R\}$ denotes the set of all epidemiological compartments.
+Consistency with observed data is enforced by the data fidelity loss
+$\mathcal{L}_{\text{data}}$ $=\sum_{X\in\mathcal{C}}\mathcal{L}_{X}^{\text{data}},$
+$\mathcal{L}_{X}^{\text{data}}$ $=\frac{1}{s}\sum_{i=1}^{s}|X(t_{i})-X_{i}^{\text{obs}}|^{2},\quad X\in\mathcal{C},$
+where $X_{i}^{\text{obs}}$ represents the observed value of compartment $X$ at time $t_{i}$.
+The DINN framework solves the optimization problem
+$(\mathbf{w}^{*},\mathbf{b}^{*},\bm{\chi}^{*})=\arg\min_{\mathbf{w},\mathbf{b},\bm{\chi}}\left(\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{residual}}\right),$ (73)
+where the parameter vector $\bm{\chi}=\{\beta,\sigma,\gamma_{s},\eta_{a},\eta_{d},\varepsilon,v,\alpha\}$ contains key epidemiological constants governing transmission, progression, and control measures.
+This articulation guarantees that the solutions that are derived to fit both meet the observed data requirements and appear as per the mechanistic framework characteristic in the Ebola model.
 6.2 Network Architecture and Training
-The design of the neural networks processes sequential data through several hidden layers, with hyperbolic tangent activations, leading to outputs associated with the epidemiological model’s eight state variables.
-Composite loss functions leverage both Xavier initialization and the Adam optimization method to facilitate training.
-Fractional derivatives are computed through automatic differentiation, and the enforcement of physical constraints within the entire temporal domain remains intact.
-8.3 Parameter Estimation and Model Calibration
-While building the model, the following epidemiological parameters need to be defined: $\Lambda$, $\mu$, $\beta$, $\eta_{a}$, $\eta_{d}$, $\varepsilon$, $\omega$, $\sigma$, $p$, $h_{s}$, $\gamma_{s}$, $\gamma_{a}$, $\gamma_{h}$, $\delta_{s}$, $\delta_{h}$, $\mu_{d}$, $\alpha$, the fractional order, the time-dependent functions $c(t)$ and $v(t)$, and the network parameters $w$ and $b$.
-Parameter estimation for the model will be refined as data for the model will become available.
-As outbreak data becomes available, the configuration and calibration will be refined to minimize the discrepancy between the data and the bounds defined by outbreak data.
-Parameter definition will subsequently be based on the mechanistic model and data from the identified model components, which will be integrated dynamically.
-Parameter definition will subsequently be based on the mechanistic model and data from the identified model components, which will be integrated dynamically.
+DINN architecture is a network of completely interconnected neural networks, with several hidden layers and with hyperbolic tangent activation functions.
+The network maps temporal inputs $t\in[t_{0},T]$ to the eight state variables $\mathbf{Y}(t)=[S(t),V(t),E(t),I_{s}(t),I_{a}(t),H(t),D(t),R(t)]^{\top}$.
+Xavier initialization is used to initialize network weights in order to ensure that there is no unstable flow of the gradient during training.
+The Adam algorithm, a combination of adaptive learning rates and momentum-based updates, is used to perform optimization in order to make the process efficient.
+By applying automatic differentiation, the fractional derivatives are calculated as $^{C}D_{t}^{\alpha}\mathcal{NN}_{\mathbf{w},\mathbf{b}}(t)$, which is used to compute the exact gradient of the fractional-order Ebola model and at the same time ensures that the physical constraints of the model are met.
+This approach ensures that the neural network solutions that are obtained are consistent with the epidemiological processes throughout the time prospect.
+6.3 Parameter Estimation and Calibration
+The DINN framework simultaneously estimates epidemiological parameters $\bm{\chi}$ and network parameters $(\mathbf{w},\mathbf{b})$ through constrained optimization of the composite loss function (73).The parameter calibration aims at reducing the difference between the network predictions and the measured outbreak data, and maintaining conformity with the mechanistic nature of the fractional-order system.
+As the data on the outbreaks available, the DINN enables the dynamic optimization of the parameters estimates in terms of iterative optimization.
+This adaptive calibration procedure makes sure that predictions do not go wrong because the epidemic changes, but the model retains the biological consistency.
+The framework also offers accurate forecasts in the short term and sound predictions in the long term that are based on the principle of epidemiology.
 6.4 Results and Discussion of DINN
+Figure 12: Susceptible analysis (a) Dynamics of output and error measures (b) Reduction in mean squared error with optimal validate value of 0.00071 (c) State variable dynamics (d) Distribution of errors (e) regression plots (f) autocorrelation analysis
+Figure 13: Vaccination analysis (a) The trend of output and error analysis (b) The demonstration of the minimization of the mean-squared error (MSE), optimal validation is the error of 0.00009 (c) Explanations of the state evolution (d) The presentation of the error histogram (e) Explanation of the regression plots (f) Autocorrelation analysis
+Figure 14: Exposed analysis (a) output and error trends (b) MSE reduction with best validation 0.00010 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
+Figure 15: Symptomatic analysis (a) output and error trends (b) MSE reduction with best validation 0.00019 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
+Figure 16: Asymptomatic analysis (a) output and error trends (b) MSE reduction with best validation 0.00015 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
+Figure 17: Hospitalized analysis (a) output and error trends (b) MSE reduction with best validation 0.00026 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
+Figure 18: Deceased analysis (a) output and error trends (b) MSE reduction with best validation 0.00034 (c) state evolution (d) error histogram (e) regression plots (f) autocorrelation analysis
+Figure 19: Recovered analysis (a) output and error trends (b) MSE reduction with best validation 0.00085 (c) state evolution; (d) error histogram (e) regression plots (f) autocorrelation analysis
 The results from the implementation of the Disease-Informed Neural Network (DINN) context were carefully reviewed in relation to the benchmark datasets.
 The results were consolidated across all 8 epidemiological compartments.
-Integrated modeling accuracy estimation as performed records from Figure 29 to 36 showcase the framework to outstanding.
+Integrated modeling accuracy estimation as performed records from Figure 12 to 19 showcase the framework to outstanding.
 The model collects above average predictive accuracy and extensive generalization, proving the framework’s relevancy, particularly on the Ebola virus disease predictive modeling’s sophisticated system dynamics.
@@
 It guarantees that future real-time forecasts and scenarios will be provided.
-Figure 29: Susceptible compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00071; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
-Figure 30: Vaccinated compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00009; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
-Figure 31: Exposed compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00010; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
-Figure 32: Symptomatic compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00019; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
-Figure 33: Asymptomatic compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00015; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
-Figure 34: Hospitalized compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00026; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
-Figure 35: Deceased compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00034; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
-Figure 36: Recovered compartment evaluation: (a) output and error trends; (b) MSE reduction with best validation 0.00085; (c) state evolution; (d) error histogram; (e) regression plots; (f) autocorrelation analysis
 7 Conclusion
-This work has been done to enhance the mathematical framework of forecasting and possibility of intervention regarding the outbreaks of Ebola Virus Disease (EVD).
-The identified primary objectives focused on the creation of an advanced fractional order model to capture the EVD spread dynamics through adaptive multilayered channels which is considered to be extremely complex, formulation of reliable epidemiological thresholds, the defining of appropriate intervention strategies, and the creation of a novel computational model on the estimation of parameters along with the model standardization.
-Integration of the three disciplines of mathematics: fractional calculus, optimal control theory, and scientific machine learning, was the original part of this research.
-EVD transmission dynamics required memory incorporated non-Markovian multi-scale dynamics.
-Thus, an eight-compartment fractional-order model was built around the Caputo derivative.
-The model well-posedness was aided by the foundational fixed-point theory, the subsequent next-generation matrix approach to compartment analysis enabled the reproduction number $\mathcal{R}_{0}$ calculation, and the complete fractal Lyapunov characterization for the system imbued a formative influence to the system.
-Utilizing Pontryagin’s Maximum Principle, the optimal control problem which was built sequentially around four time-adjustable control interventions dedicated to fractional control systems was solved.
-The main breakthrough was the design of the Disease-Informed Neural Network (DINN) framework, which dynamically computes the mechanistic parameters by embedding the governing fractional differential equations as physical constraints in a neural network and solving for parameter estimation.
-The analysis brings to light three key findings: the epidemiological threshold $\mathcal{R}_{0}$, unique bifurcation properties at the critical points tied to $\mathcal{R}_{0}$, and the findings on global stability which show a forward bifurcation at $\mathcal{R}_{0}=1$.
-One of the most important parameters indicating potential outbreak from the tiered hierarchy parameters from a sensitivity analysis are the transmission rate ($\beta$), the incubation rate ($\sigma$), and deceased infectiousness ($\eta_{d}$).
-Results from optimal control show that combining treatment with safe burial, which highlights the synergistic nature of the interventions, results in the greatest impact of an 86.5% decrease in mortality relative to the impact of each intervention applied in isolation.
-All epidemiological compartments also show outstanding predictive accuracy with the DINN context, as seen in the coefficients of determination $R^{2}$ which lie between 0.9987 and 0.9999, coupled with recovering system parameters with less than 18% accuracy for most primary transmission and progression rates.
-Although the scope of the model are comprehensive, some simplifications are made to the assumptions of the epidemiology regarding spatial diversity, age-structured contact matrices, and stochastic transmission.
-While parameter estimation is robust, and within the DINN framework, it is still a function of the investigation data quality and handling.
-It is definitely worth revealing that optimal control theory, while developing practical strategies, does have the control theory limits as the minor, and in practice resource partial, bounds.
-The gaps mentioned in the research offer many opportunities for prospective work.
-Adding spatial dynamics and meta-population structures would improve the model’s ability to channel cross regional transmission and tailored modified intervention impacts.
-Further model improvement would arise from the inclusion of stochastic elements that measure the risk of destruction and the chances of restoration of an outbreak.
-Other innovative applications of the DINN framework would be in the area of predictive public health for automatic adaptive real-time intervention planning to control ongoing outbreaks.
-This would be an immense advance in public health.
-Also, implementing this integrated mathematical framework with machine learning to other emerging infectious diseases would extremely transform the epidemic awareness and response framework.
-This study underscores how the integration of advanced techniques in mathematics and contemporary computing techniques provides valuable tools for public health.
-It illustrates the transition of epidemiology from purely descriptive to prescriptive, allowing for active intervention in the control of outbreaks.
+This research aimed to develop an integrated mathematical framework that improves our ability to predict and respond to Ebola outbreaks.
+We sought to create a fractional-order model that captures the memory effects in disease transmission, design effective intervention strategies using optimal control theory, and build a computational framework that combines neural networks with epidemiological principles for accurate prediction and parameter estimation.
+Our methodology is a combination of fractional calculus, optimal control theory, and scientific machine learning, which leverages the three areas of mathematics.
+First, we developed a novel nonlinear eight-compartment model of fractional order using the Caputo derivative to represent memory effects in transmission dynamics.
+We then presented some mathematical foundations of the model, such as existence, uniqueness, and stability properties.
+Moreover, the fundamental reproduction number $R_{0}$ was obtained with the next-generation matrix, and the stability of the equilibrium states was proved.
+A global sensitivity analysis is carried out to determine the relative effect of the most important epidemiological parameters on both the model results and the basic reproduction number.
+Also, we formulated an optimal control problem with four interventions varying with time that are solved by the Maximum Principle of Pontryagin and got numerical results by the RK45 iteration method.
+At the final part of the present research, we designed a Disease-Informed Neural Network (DINN) that represents the governing equations in the form of physical constraints in a neural network architecture.
+A number of important findings were made by the study.
+Our fractional-order model was found to be effective in capturing the non-Markov nature of the dynamics of Ebola transmission to provide a more accurate representation of how the disease spreads in a memory-dependent manner.
+The sensitivity analysis established that the transmission rate, incubation rate, and infectiousness associated with dead individuals had the greatest impact on the possible size of the outbreaks.
+Our optimal control results showed that when treatment is used in combination with safe-burial measures, the targeted mortality will be reduced by 86.50%, and safe burial, as the best cost-effective approach, costs \$4 per case.
+The DINN framework displayed outstanding predictive capability, recording a value between 0.9987 and 0.9999 of $R^{2}$ in all compartments, and was able to predict 13 out of 17 epidemiological parameters with an error rate of less than 9%.
+There are a number of limitations that must be acknowledged.
+The model assumes that the population is well-mixed and does not take into consideration the spatial heterogeneity or the demographic structure.
+There were no stochastic components, which can affect the estimates of the probability of an outbreak extinction.
+Although parameter estimation has a strong robustness in our framework, it is still dependent on the quality and availability of data.
+The real-world conditions of practical implementation of optimal control strategies are restricted by real-world conditions that cannot be adequately captured by our theoretical framework.
+The studies can be extended to different tracks in future practice.
+Spatial dynamic integration would allow modeling the cross-regional transmission and geometrically focused interventions.
+The addition of stochastic factors would provide more realistic predictions of the risks of extinction and renaissance of an outbreak.
+The Deep Inference Neural Network (DINN) system can be modified to accommodate real-time outbreak control to allow interventions to be dynamically adjusted as new information is added.
+The generalization of this integrated approach to other infectious diseases would be evaluated by applying it to other pathogens, and it may transform the system of outbreak responses.
+Lastly, age-structured patterns of contact would also provide more realistic transmission dynamics and improve the evaluation of the efficacy of intervention.
+This paper shows that the integration of sophisticated mathematical modeling with modern computational methodology can produce useful tools in the decision-making of public health.
+Our ability to bridge the gap between theoretical epidemiology and practical intervention strategies will help usher in the future where mathematical models can be used to predict the path of an outbreak as well as help develop effective containment interventions.
+CRediT Authorship Contribution Statement
+[Noor Muhammad]: Conceptualization, Methodology, Formal analysis, Writing - original draft, Software, Validation.
+[Md.
+Nur Alam]: Validation, Investigation, & review .
+[Zhang Shiqing]: Supervision, Project administration.
+Declaration of Competing Interest
+The authors declare that they have no known challenging financial interests or individual relationships that could have seemed to influence the work described in this paper.
+Acknowledgments
+The authors gratefully acknowledge the academic environment and resources provided by the School of Mathematics, Sichuan University.
+We also thank the anonymous reviewers for their valuable suggestions